### v1

# Goldfish: Monolingual Language Models for 350 Languages

Abstract
For many low-resource languages, the only available language models are large multilingual models trained on many languages simultaneously. However, using FLORES perplexity as a metric, we find that these models perform worse than bigrams for many languages (e.g. 24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B). To facilitate research that focuses on low-resource languages, we pre-train and release Goldfish, a suite of monolingual autoregressive Transformer language models up to 125M parameters for 350 languages. The Goldfish reach lower FLORES perplexities than BLOOM, XGLM, and MaLA-500 on 98 of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. However, the Goldfish significantly underperform larger multilingual models on reasoning benchmarks, suggesting that for low-resource languages, multilinguality primarily improves general reasoning abilities rather than basic text generation. We release models trained on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) of text data where available. The Goldfish models are available as baselines, fine-tuning sources, or augmentations to existing models in low-resource NLP research, and they are further useful for crosslinguistic studies requiring maximally comparable models across languages.
Goldfish: Monolingual Language Models for 350 Languages
[…]
1 Introduction
Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023). For many low-resource languages, a dedicated model optimized for that language does not exist. This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022). These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020).
To address this lack of available models, we introduce Goldfish, a suite of over 1000 monolingual language models for 350 diverse languages. [^1] The name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024). The models reach lower perplexities than XGLM (Lin et al., 2022), BLOOM 7.1B (Scao et al., 2022), and MaLA-500 (Lin et al., 2024) on 98 out of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller. The Goldfish also outperform simple bigram models, which are surprisingly competitive with larger models for low-resource languages (e.g. lower perplexities than BLOOM 7.1B on 43% of its languages; §4). However, despite better perplexities, the Goldfish underperform larger multilingual models on reasoning benchmarks, suggesting that multilingual pre-training may benefit abstract reasoning capabilities over more basic grammatical text generation (§5).
Data size Model output 5MB ‘‘Goldfish are a few years of the most of the most of the most...’’ 10MB ‘‘Goldfish are a great way to the best way to the best way...’’ 100MB ‘‘Goldfish are a great way to get your fish in the wild.’’ 1GB ‘‘Goldfish are a species of fish that are found in the sea.’’
Figure 1: Left: Map of the 350 languages for which Goldfish models are available, using coordinates from Glottolog (Hammarström et al., 2023). Right: Sample model outputs completing the prompt ‘‘Goldfish are’’ for the eng_latn (English) model for each dataset size, using sampling temperature zero. Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4).
Finally, to enable comparisons across languages, we release monolingual models trained on comparable dataset sizes for all languages: 5MB, 10MB, 100MB, and 1GB when available, after accounting for the fact that languages require different numbers of UTF-8 bytes to encode comparable content (Arnett et al., 2024). These Goldfish serve as baselines, allowing results in diverse languages to be situated relative to comparable models. They can also be used as source models for fine-tuning or to enhance larger multilingual models in areas where those models fall short (§4). Models and code are available at https://huggingface.co/goldfish-models.
2 Related Work
Low resource language modeling often leverages multilingual pre-training, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020). Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2023). However, monolingual models have still been shown to achieve better performance than multilingual models for many languages (e.g. Martin et al., 2020; Pyysalo et al., 2021; Gutiérrez-Fandiño et al., 2021; Luukkonen et al., 2023). Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020; Chang et al., 2023).
Notably, the training datasets for massively multilingual models are often heavily skewed towards high-resource languages. For example, XGLM 4.5B is trained on over 7000$\times$ more Norwegian (71GB; 5.4M native speakers) than Quechua (0.01GB; 7.3M native speakers; Lin et al., 2022; Ethnologue, 2024). In a more extreme case, BLOOM is trained on only 0.07MB of Akan (8.1M native speakers) out of 1.61TB total (4e-6% of the pre-training dataset; Scao et al., 2022). These extremely small quantities of low-resource language data often do not leverage recent efforts to compile text data in low-resource languages (Costa-jussà et al., 2022; Imani et al., 2023; Kudugunta et al., 2023), and the data imbalances are likely to severely hinder performance in low-resource languages. Indeed, we find that these models have worse perplexities than simple bigram models for many languages (§4). Unfortunately, comparable monolingual language models across many diverse languages have yet to be studied or released.
3 Models and Datasets
We introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pre-trained for 350 languages. The largest model for each language is 125M parameters. We train models on 5MB, 10MB, 100MB, and 1GB of text when available after byte premium scaling (Arnett et al., 2024). Figure 1 shows a geographic map of the 350 languages, with coordinates from Glottolog (Hammarström et al., 2023), along with sample outputs from the English model for each dataset size.
Goldfish data size # Langs Goldfish Bigrams XGLM 4.5B MaLA-500 10B
1000MB 73 76.9 112.3 78.6 84.7
100MB 22 102.7 132.6 143.9 121.7
10MB, 5MB 5 130.5 148.3 183.1 135.0
Table 1: Mean FLORES log-perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size. The Goldfish languages are a strict superset of these languages.
3.1 Training Datasets
We merge the massively multilingual text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language. To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022). We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data. Full dataset details are in §A.1.
To sample pre-training datasets of the desired sizes in a language $L$, we first use the Byte Premium Tool (Arnett et al., 2024) to estimate the byte premium for $L$, the number of UTF-8 bytes required to encode comparable text in $L$ relative to eng_latn (English). For example, khm_khmr (Khmer) has byte premium 3.91, meaning that it uses approximately 3.91$\times$ as many UTF-8 bytes as English to encode content-matched text. We divide each dataset size by the estimated byte premium for the corresponding language, thus measuring all datasets in units of “equivalent” English text bytes. We sample datasets to train monolingual language models on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) when available after byte premium scaling. [^2] The languages with 5MB-dataset models are a subset of the languages with 10MB-dataset models, and similarly for the 100MB and 1GB dataset sizes. These are equivalent to roughly 1M, 2M, 20M, and 200M tokens of English text respectively; including 10 epochs of repetition, the 1GB-dataset models are trained on the equivalent of roughly 2B English tokens. When a 1GB dataset is not available for a language after byte premium scaling, we include a full model (267 languages) trained on the entire dataset in that language, for use cases that seek to maximize performance in a specific low-resource language.
3.2 Architectures and Pre-Training
For each language and each dataset size, we pre-train an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019). For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019). Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes. Full hyperparameters are reported in §A.2.
We tokenize each dataset using using a monolingual SentencePiece tokenizer (Kudo and Richardson, 2018) trained on that dataset size, limiting tokenizer training text to 100MB after byte premium scaling. Following Liu et al. (2019), we use vocabulary size 50K and a maximum sequence length of 512 tokens for all models. We train each language model on 10 epochs of its corresponding dataset. [^3] Multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but we find that more than 10 epochs of training leads to overfitting for extremely small datasets (e.g. 5MB). Pre-training details, compute costs, and all available models are reported in §A.2.
Bigrams XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
Bigrams 24 / 102 0 / 30 20 / 46 11 / 175
Goldfish (ours) 202 / 202 60 / 102 2 / 30 32 / 46 111 / 175
Table 2: FLORES perplexity win rates for each row vs. column model. For example, Goldfish reach lower log-perplexities than MaLA-500 for 111/175 (63%) of FLORES languages in both Goldfish and MaLA-500.
# Langs Chance Goldfish XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
Belebele 121 25.0 28.2 30.1 30.6 30.2 30.6
XCOPA 11 50.0 54.9 57.9 60.6 56.9 55.6
XStoryCloze 10 50.0 52.5 57.1 59.9 58.2 55.7
Table 3: Reasoning benchmark accuracies averaged over non-English languages. Despite better perplexities, the Goldfish perform significantly worse than larger multilingual models on reasoning.
4 FLORES Log-Perplexity Evaluations
We first evaluate our models on FLORES-200 log-perplexity (Costa-jussà et al., 2022) (equivalently, negative log-likelihood; Lin et al., 2024). To avoid tokenization confounds from computing log-perplexity per token, we compute log-perplexity per FLORES sequence. Regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES. In most cases, $s$ is a single sentence. For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$. We then compute the mean over sequences:
$\textrm{LogPPL}_{\mathcal{M}}=\textrm{mean}_{s}\Big{(}-\textrm{log}(P_{\mathcal{M}}(s_{1}|s_{0}))\Big{)}$ (1)
A lower log-perplexity indicates better performance, where $\mathcal{M}$ assigns higher probabilities to ground truth text (FLORES sequences). While imperfect, perplexity does not require annotated text data, it is predictive of performance on a variety of downstream tasks (Xia et al., 2023), and it has been used to measure language model quality in previous work (Kaplan et al., 2020; Hoffmann et al., 2022; Lin et al., 2024).
We compare the Goldfish models with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages). We also compare to simple bigram models trained on the Goldfish datasets. [^4] Bigram and perplexity implementation details in §A.3. In all cases, we use the Goldfish model trained on the maximum amount of data in each language (maximum 1GB).
FLORES log-perplexity results.
The Goldfish reach lower log-perplexities than all four comparison models on 98 of the 204 FLORES languages. Average log-perplexities for the 100 FLORES languages included in both XGLM 4.5B and MaLA-500 are reported in Table 1 (excluding XGLM 7.5B and BLOOM 7.1B there because they are trained on far fewer languages). On average, the Goldfish reach 13% lower log-perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B.
To ensure that these results are not driven by a small subset of specific languages, in Table 2 we report the pairwise “win” rates for Goldfish and bigrams vs. all four comparison models, for the set of FLORES languages shared between each pair. The Goldfish models have a perplexity win rate above 50% against all comparison models except XGLM 7.5B, which considers only 30 fairly high-resource languages (Lin et al., 2022). Notably, the bigram models also reach lower perplexities than large multilingual models for a nontrivial number of languages: 24% of languages in XGLM 4.5B and 43% of languages in BLOOM 7.1B. Still, the bigrams have worse perplexities than Goldfish for all languages. Log-perplexities for individual languages and models are reported in Table LABEL:tab:ppl-table.
5 Multilingual Reasoning Benchmarks
Because FLORES perplexities are not necessarily reflective of complex capabilities in language models, we also evaluate Goldfish, XGLM, BLOOM, and MaLA-500 (as in §4) on non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense; Lin et al., 2022). All models are evaluated zero shot with no fine-tuning. Evaluation task details are in §A.4.
Results for all three reasoning tasks are reported in Table 3. Although all models perform quite poorly (close to chance accuracy), the Goldfish perform substantially worse than the multilingual models. [^5] It is unlikely that this effect is due to model size alone; the Goldfish models (125M parameters) have easily enough capacity for their maximum of 1GB of text data. This indicates that the combination of larger datasets and model sizes in multilingual pre-training can allow language models to develop reasoning capabilities in specific languages, even when perplexities in those languages remain high. For example, XGLM 7.5B has worse perplexities than Goldfish for 82 Belebele languages (in fact, worse than bigrams for 77 languages), but it outperforms Goldfish on Belebele (reading comprehension) for 56 of those languages. This is in stark contrast with monolingual language models, which generally must reach low perplexities and acquire basic grammatical capabilities before developing reasoning abilities (Liu et al., 2021; Choshen et al., 2022; Xia et al., 2023; Chang et al., 2024). Intuitively, it may be that abstract reasoning patterns are often more language-agnostic than grammatical text generation, and thus multilingual pre-training primarily benefits the former.
6 Conclusion
We pre-train and release Goldfish, a suite of over 1000 monolingual language models for 350 languages. The Goldfish achieve perplexities that are competitive with, and on average lower than, state-of-the-art multilingual language models across languages. However, they underperform large multilingual models on reasoning tasks; in low-resource languages, it appears that multilingual pre-training facilitates nontrivial reasoning capabilities despite extremely poor perplexities. We publicly release all Goldfish models to be used as comparable baselines, fine-tuning sources, or augmentations to larger models (e.g. cross-lingual experts; Blevins et al., 2024) in future low-resource NLP research.
Limitations
Comparability and availability.
In order to include as many low-resource languages as possible, the Goldfish models are trained on corpora compiled from a wide variety sources (§A.1). Still, 5MB of text (roughly 1M tokens) is not publicly available for many of the world’s languages. Even where text is available, corpora for different languages vary significantly both in cleanliness and domain coverage (e.g. news vs. social media vs. books). Thus, while we release models trained on comparable quantities of text in different languages (including accounting for byte premiums; Arnett et al., 2024; §3.1), the models are not perfectly comparable across languages. In fact, it is likely that such perfect comparability is impossible given the diversity of the world’s languages, cultures, and language use. Even directly translated datasets are not perfectly comparable across languages (Jill Levine and Lateef-Jan, 2018). Thus, the Goldfish models aim to maximize model and dataset comparability across languages while still covering a wide variety of languages.
Monolinguality.
By design, all of the Goldfish models are monolingual. For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020; Chang et al., 2023). However, adding multilingual data introduces concerns such as the choice of added languages (some languages have more closely related languages in our dataset than others), quantities of added data, and model capacity limitations. To maximize comparability across languages and to allow the models to serve as clearly-defined baselines, we train all Goldfish models monolingually. Of course, language-annotated text datasets inevitably contain mislabeled text, particularly for similar languages (Caswell et al., 2020; Blevins and Zettlemoyer, 2022; Kreutzer et al., 2022). Thus, we cannot guarantee that our models are entirely free from cross-language contamination, although they are monolingual to the best ability of current language identification models.
Model and dataset sizes.
Because the Goldfish are focused on low-resource languages, we restrict all models to 1GB of training text (after byte premium scaling; Arnett et al., 2024). For the majority of the world’s languages, 1GB is sufficient to include all publicly available text data in the language. At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2023). Thus, the largest Goldfish model that we train for each language has 125M parameters and is trained on a maximum of 1GB of text. This is the same model size as GPT-1 (Radford et al., 2018) or BERT (Devlin et al., 2019), and the 1GB dataset size is approximately 20% of the dataset size of GPT-1 (Radford et al., 2018).
Downstream tasks.
We evaluate the Goldfish models on FLORES log-perplexity (§4) and three reasoning benchmarks (§5). These are some of the only evaluations that can be used for autoregressive language models in many languages, but they have significant limitations. Perplexity is not necessarily predictive of grammatical text generation (Hu et al., 2020) or complex reasoning capabilities (Levy et al., 2024), but it still provides reasonable signal for model performance (Xia et al., 2023) and it is often used to roughly quantify language model quality (Kaplan et al., 2020; Hoffmann et al., 2022). On the other hand, reasoning benchmarks require annotated datasets and thus often cover fewer languages. One notable exception is Belebele (121 non-English languages; Bandarkar et al., 2024), but even large state-of-the-art models perform quite poorly on Belebele without tuning or few-shot prompting (§5). Thus, our evaluations of model reasoning are not entirely conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”). We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future.
Risks and dataset licensing.
Trained on a maximum of 1GB of text each, the Goldfish models have very limited capabilities relative to modern language models in high-resource languages. The Goldfish are trained on publicly-released corpora used in previous NLP research (§A.1), but we cannot guarantee that the data is free from offensive content or personally identifying information. We do not redistribute the data itself. Furthermore, our models are small, which reduces the likelihood that they will regurgitate memorized text (Carlini et al., 2023). As far as we are aware, we do not include any datasets that prohibit use for language model training. We report all included datasets in §A.1. We will remove models for affected languages if contacted by dataset owners.
Acknowledgments
We would like to thank the UCSD Language and Cognition Lab for valuable discussion. Some models were trained on hardware provided by the NVIDIA Corporation as part of an NVIDIA Academic Hardware Grant. Some models were also trained on the UCSD Social Sciences Research and Development Environment (SSRDE). Zhuowen Tu is supported by NSF IIS-2127544. Tyler Chang is partially supported by the UCSD HDSI graduate fellowship.
Appendix A Appendix
A.1 Training Dataset Details
Data sources.
As described in §3.1, we merge the text datasets compiled in Chang et al. (2023), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023). These datasets include popular multilingual corpora such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), and others. Together, these datasets take advantage of both automatically crawled datasets with automated language identification and targeted datasets manually annotated for specific low-resource languages. All included datasets are publicly available; see Limitations for licensing concerns. Comprehensively, the Goldfish dataset includes:
  • Chang et al. (2023):
    OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), Leipzig Corpora Collection (Goldhahn et al., 2012), eBible translations (eBible, 2023), Tatoeba (Tiedemann, 2012, 2020), AfriBERTa (Ogueji et al., 2021), NusaX (Winata et al., 2023), AmericasNLP (Mager et al., 2021), Nunavut Hansard Inuktitut–English Parallel Corpus (Joanis et al., 2020), Cherokee-English ChrEn dataset (Zhang et al., 2020), Cherokee Corpus (Cherokee Corpus, 2023), Cree Corpus (Teodorescu et al., 2022), Languages of Russia (Zaydelman et al., 2016), Evenki Life newspaper (Zueva et al., 2020), transcribed Fula Speech Corpora (Cawoylel, 2023), IsiXhosa (Podile and Eiselen, 2016), Ewe Language Corpus (Gbedevi Akouyo et al., 2021), Makerere Luganda Corpora (Mukiibi et al., 2022), CMU Haitian Creole dataset (CMU, 2010), Tigrinya Language Modeling Dataset (Gaim et al., 2021), and Ulukau (Ulukau, 2023).
  • Glot500 (Imani et al., 2023):
    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008), and XLSum (Hasan et al., 2021).
  • MADLAD-400 (Kudugunta et al., 2023):
    CommonCrawl (Common Crawl, 2022).
We start with the corpus from Chang et al. (2023). We then merge the dataset per language with Glot500 for languages that have not yet reached our 1GB maximum (after byte premium scaling). Then, we merge the dataset with MADLAD-400 for languages that have still not reached our 1GB maximum. We also add MADLAD-400 for languages with short average line lengths (less than 25.0 tokens), to make use of MADLAD-400’s longer contiguous sequences. To allow comparisons on popular low-resource language evaluations, we exclude FLORES-200 (Costa-jussà et al., 2022) and AmericasNLI (Ebrahimi et al., 2022) from all dataset merging. For each dataset, we exclude languages that contain only Bible data. Because there is likely significant overlap between different dataset sources, we deduplicate repeated sequences of 100 UTF-8 bytes for each language (Lee et al., 2022).
Language codes.
To enable dataset merging per language, several datasets must be converted to ISO 639-3 language codes and ISO 15924 script codes. In some cases, this introduces ambiguity because datasets can be labeled as individual language codes (e.g. quy_latn for Ayacucho Quechua and quz_latn for Cusco Quechua) or as macrolanguage codes (e.g. que_latn for Quechua). In these cases, we compile both a macrolanguage dataset and individual language datasets. Datasets labeled with individual codes contribute both to their individual dataset and their umbrella macrolanguage dataset; datasets labeled with macrolanguage codes contribute only to the macrolanguage dataset. For example, we have individual quy_latn and quz_latn datasets, both of which contribute to a larger que_latn dataset, which also contains datasets labeled only with que_latn. These ambiguities primarily appear for lower-resource languages.
Additionally, we drop several redundant language codes:
  • We drop ory_orya (Odia) in favor of the macrocode ori_orya because ory_orya is the only individual language within ori_orya for which we have any data.
  • For the same reason, we drop npi_deva (Nepali) in favor of the macrocode nep_deva.
  • For the same reason, we drop swh_latn (Swahili) in favor of the macrocode swa_latn.
  • We drop cmn_hans (Mandarin) in favor of the macrocode zho_hans (Chinese) because the zho_hans data is almost entirely in Mandarin. While less specific, zho_hans is commonly used by other datasets. For other Chinese languages, see their individual codes (e.g. yue_hant for Cantonese). We note that the similar code zho_hant (traditional characters) is not primarily Mandarin.
  • We drop hbs_cyrl and hbs_latn (Serbo-Croatian) because we have the individual languages Serbian (srp_cyrl and srp_latn), Croatian (hrv_latn), and Bosnian (bos_cyrl and bos_latn).
  • We drop the deprecated code ajp_arab (Levantine Arabic) in favor of apc_arab.
  • We drop ber_latn (Berber) because it is a collective code for distinct (and often not mutually intelligible) languages. We keep the constituent individual languages.
  • We drop nah_latn (Nahuatl) because it is a collective code for distinct languages. We keep the constituent individual languages.
After merging, we have a dataset of 547GB of text covering 523 language-script combinations (486 unique language codes, 32 unique script codes).
Byte premiums.
As described in §3.1, we then scale our dataset sizes by estimated byte premiums (Arnett et al., 2024). A byte premium $b$ for a language $L$ indicates that content-matched (i.e. parallel) text in $L$ takes $b\times$ as many UTF-8 bytes to encode as English. We use the Byte Premium Tool (Arnett et al., 2024) to compute or estimate the byte premium for all of our languages. Byte premiums are pre-computed in the tool for high-resource languages. For each novel low-resource language $L$, we use the tool (which uses a linear regression) to predict the byte premium for $L$ based on the character entropy for text in $L$ and the script type for $L$ (alphabet, abjad, abugida, or logography), as recommended for low-resource languages in Arnett et al. (2024). Then, we have an estimated byte premium for every language in our dataset. We clip each byte premium to a minimum of 0.70 and a maximum of 5.00; clipping occurs for only three languages (lzh_hant, wuu_hani $\to$ 0.70, mya_mymr $\to$ 5.00). As described in §3.1, all of our training datasets (both for tokenizers and for the models themselves) are sampled based on size in bytes after byte premium scaling. We drop languages with less than 5MB of text after byte premium scaling.
Dataset statistics.
The resulting 350 Goldfish languages cover five continents, 28 top-level language families (Hammarström et al., 2023), and 32 scripts (writing systems). All languages for which Goldfish models are available are listed in Table LABEL:tab:lang-list. We include the language name, ISO 639-3 language code, ISO 15924 script code, estimated byte premium, dataset size after byte premium scaling, dataset size in tokens, and proportion of the dataset from each of our four largest sources. Raw dataset sizes before byte premium scaling can be obtained by multiplying the dataset size after byte premium scaling by the estimated byte premium. Source dataset proportions are reported before deduplication. The reported dataset sizes reflect the dataset for the Goldfish model trained on the maximum amount of data for that language (the 1GB-dataset Goldfish when available, otherwise the full-dataset Goldfish). Reported token counts use the tokenizer for the largest Goldfish model for that language. All dataset statistics can be downloaded at https://github.com/tylerachang/goldfish.
A.2 Pre-Training Details
As described in §3.2, we train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full. The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language. In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total). Full hyperparameters are reported in Table 4.
Hyperparameter 5MB,10MB 100MB,1GB,full
Total parameters 39M 125M
Layers 4 12
Embedding size 512 768
Hidden size 512 768
Intermediate hidden size 2048 3072
Attention heads 8 12
Attention head size 64 64
Learning rate 1e-4
Batch size 5MB: 4, 10MB: 8,
100MB: 32, 1GB: 64
Epochs 10
Activation function GELU
Max sequence length 512
Position embedding Absolute
Learning rate decay Linear
Warmup steps 10% of pre-training
Adam $\epsilon$ 1e-6
Adam $\beta_{1}$ 0.9
Adam $\beta_{2}$ 0.999
Dropout 0.1
Attention dropout 0.1
Table 4: Pre-training hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
Tokenizers.
All tokenizers are trained with vocabulary size 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling). We use SentencePiece tokenizers (Kudo and Richardson, 2018) with training text randomly sampled from the dataset for the desired language. To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling. After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens. We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form. Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling. [^6] When de-tokenized, the tokenized datasets result in slightly smaller datasets than the original text datasets, because the tokenizer truncates lines to create 512-token sequences. All reported dataset sizes account for this truncation.
Architectures.
All of our models use the GPT-2 architecture (Radford et al., 2019), changing only the number of layers, attention heads, and embedding sizes as in Turc et al. (2019). For the 100MB-, 1GB-, and full-dataset models, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018) (similar to BERT-base and RoBERTa; Devlin et al., 2019; Liu et al., 2019). Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al. (2019) for the 10MB and 5MB dataset sizes.
Training hyperparameters.
Language models are pre-trained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022). We refrain from extensive hyperparameter tuning to avoid biasing our hyperparameters towards English (or any other selected tuning language). Instead, we adopt hyperparameters from previous work with minimal modifications. To match the setup of our models and to prevent overfitting, we select hyperparameters based on models with fairly small training datasets relative to modern standards. Specifically, following BERT (Devlin et al., 2019), we use learning rate 1e-4 for the 125M-parameter models (the same as RoBERTa for small batch sizes; Liu et al., 2019; GPT-1 uses learning rate 2.5e-4; Radford et al., 2018). Based on initial results using randomly-sampled languages, we find that learning rate 1e-4 also works well for the 39M-parameter models; this is in line with Chang et al. (2023), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
We train each model for 10 epochs of the training data; multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pre-training on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios. For batch sizes, following GPT-1 (most similar to our models; Radford et al., 2018), we use batch size 64 (64$\times$512 = 32K tokens) for the 1GB-dataset models. We find that these larger batch sizes lead to overfitting for small datasets, so we use batch sizes 4, 8, and 32 for 5MB-, 10MB-, and 100MB-dataset models respectively (determined based on initial experiments with randomly-sampled languages). These correspond to batches of 2K, 4K, or 16K tokens. For full-dataset models, we use the batch size that would be used if rounding the dataset size down to 5MB, 10MB, or 100MB (recall that we do not train a full-dataset model when the 1GB dataset is available for a language).
Compute costs.
All language model pre-training runs together take a total of $1.65\times 10^{20}$ FLOPs. This is less than 1/1900×1/1900$\times$1 / 1900 × the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020; $3.14\times 10^{23}$ FLOPs). Models are each trained on one NVIDIA GeForce GTX TITAN X, GeForce RTX 2080 Ti, TITAN Xp, Quadro P6000, RTX A4500, RTX A5000, or RTX A6000 GPU. In total, Goldfish pre-training takes the equivalent of approximately 15600 A6000 GPU hours. Inference for FLORES perplexities and reasoning benchmarks takes approximately 250 A6000 GPU hours (primarily due to the large multilingual models used for comparison). Dataset merging, deduplication, and tokenization takes approximately 1600 CPU core hours.
A.3 FLORES Evaluation Details
[…]
FLORES log-perplexities for all models and languages are reported in Table LABEL:tab:ppl-table. For Goldfish models, we report the log-perplexity for the model trained on the largest dataset for the language (i.e. the 1GB-dataset model when available, otherwise the full-dataset model). Log-perplexities of the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.
Bigram model details.
For each FLORES language, we train a bigram model on the entire Goldfish dataset for that language, up to 1GB after byte premium scaling §3.1. The bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset. The tokenizer is the same as the Goldfish tokenizer for that dataset (i.e. the 1GB-dataset model when available, or the full-dataset model). When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e. “stupid backoff”; Brants et al., 2007). We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.
Ambiguous or missing languages.
Several of the FLORES and Belebele languages are either missing from Goldfish or have multiple possible Goldfish available (e.g. either the macrolanguage que_latn or individual language quy_latn for FLORES language quy_latn). We make the following substitutions:
  • taq_tfng $\to$ None,
    tzm_tfng $\to$ None.
    None of the language models evaluated are trained on these languages, and no Goldfish are trained with the Tifinagh (tfng) script.
  • awa_deva $\to$ hin_deva,
    kam_latn $\to$ kik_latn.
    kas_arab $\to$ urd_arab,
    mni_beng $\to$ ben_beng,
    nus_latn $\to$ din_latn,
    taq_latn $\to$ kab_latn,
    Here, we use the closest relative in Goldfish that uses the same script.
  • ace_arab $\to$ urd_arab,
    arb_latn $\to$ mlt_latn,
    ben_latn $\to$ hin_latn,
    bjn_arab $\to$ urd_arab,
    min_arab $\to$ urd_arab,
    npi_latn $\to$ hin_latn,
    sin_latn $\to$ hin_latn,
    urd_latn $\to$ hin_latn,
    These are languages that are missing from Goldfish and that are written in a nonstandard script for the language (e.g. Arabic in Latin script). We use the closest relative in Goldfish that uses that script.
  • acm_arab $\to$ arb_arab,
    acq_arab $\to$ arb_arab,
    aeb_arab $\to$ arb_arab,
    ajp_arab $\to$ arb_arab,
    als_latn $\to$ sqi_latn,
    ars_arab $\to$ arb_arab,
    ary_arab $\to$ arb_arab,
    ayr_latn $\to$ aym_latn,
    azb_arab $\to$ aze_arab,
    azj_latn $\to$ aze_latn,
    dik_latn $\to$ din_latn,
    gaz_latn $\to$ orm_latn,
    khk_cyrl $\to$ mon_cyrl,
    kmr_latn $\to$ kur_latn,
    lvs_latn $\to$ lav_latn,
    npi_deva $\to$ nep_deva,
    ory_orya $\to$ ori_orya,
    pbt_arab $\to$ pus_arab,
    plt_latn $\to$ mlg_latn,
    quy_latn $\to$ que_latn,
    swh_latn $\to$ swa_latn,
    uzn_latn $\to$ uzb_latn,
    ydd_hebr $\to$ yid_hebr,
    yue_hant $\to$ zho_hant,
    zsm_latn $\to$ msa_latn,
    These languages map to multiple different Goldfish languages or are individual languages within a macrolanguage code included in Goldfish. When the option is available, we use the Goldfish language with more data.
A.4 Reasoning Task Details
In §5, we evaluate the Goldfish models, XGLM 4.5B, XGLM 7.5B, BLOOM 7.1B, and MaLA-500 10B on:
  • Non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024). For languages that are ambiguous or missing from Goldfish, we use the same language code mapping as in §A.3. Each Belebele example consists of a passage, a question, and four candidate answers. We evaluate model accuracy in selecting the correct answer by computing text probabilities for each “[passage] [question] [answer_option]”. No model exceeds 41% accuracy for any language (random chance 25%).
  • XCOPA (11 languages, commonsense reasoning; Ponti et al., 2020). Each example consists of a premise sentence and two possible causes or effects (i.e. answer options). We use the task format and evaluation implementation in Gao et al. (2023). This selects answers based on a model’s computed text probabilities for each “[premise] [connecting_word] [cause/effect_option]”, where the connecting word is the translation of “because” (for causes) or “therefore” (for effects).
  • Non-English XStoryCloze (10 languages, story commonsense; Lin et al., 2022). Each example consists of a context story and two possible story completions. We use the task format and evaluation implementation in Gao et al. (2023). This selects answers based on a model’s computed text probabilities for each “[story_context] [completion_option]”.
All models are evaluated zero shot with no fine-tuning. Results per language are available at https://github.com/tylerachang/goldfish.
Table 5: FLORES log-perplexity score ($\downarrow$) for each model and FLORES language. Parentheses indicate that the model is not trained specifically on that language.
Language   FLORES log-perplexity
Goldfish Bigram XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
ace_arab (287.55) (365.59) (260.50) (263.32) (232.76) (251.92)
ace_latn 144.62 169.33 (202.67) (208.64) (198.50) 133.10
acm_arab (96.60) (124.48) (93.25) (88.91) (85.28) 102.65
acq_arab (95.54) (124.94) (92.14) (88.09) (82.88) (105.38)
aeb_arab (116.01) (140.30) (113.84) (108.98) (102.76) (120.88)
afr_latn 79.88 115.57 79.54 (162.73) (153.05) 85.61
ajp_arab (98.92) (125.28) (96.25) (91.49) (85.73) 103.56
aka_latn 132.48 162.51 (234.93) (239.68) 187.66 128.37
als_latn 77.28 119.91 76.37 (220.78) (178.58) 89.49
amh_ethi 83.36 111.87 110.54 (266.96) (195.14) 108.99
apc_arab 173.18 179.37 (99.97) (94.47) (87.38) 106.58
arb_arab 82.43 117.61 79.33 75.07 69.24 96.84
arb_latn (245.97) (346.80) 211.13 (226.14) (221.60) (229.07)
ars_arab (83.68) (118.50) (80.72) (76.51) (70.75) (98.25)
ary_arab (128.66) (155.55) (131.80) (125.88) (114.30) 123.30
arz_arab 116.98 146.50 (98.72) (92.96) (87.42) 112.84
asm_beng 93.78 118.79 135.60 (227.40) 113.74 108.20
ast_latn 86.86 118.36 (113.30) (112.31) (99.39) 82.54
awa_deva (128.70) (169.05) (148.81) (135.84) (123.28) (141.50)
ayr_latn 123.33 148.81 (239.30) (231.82) (228.59) 146.85
azb_arab 154.24 185.83 163.32 (218.12) (225.63) 165.85
azj_latn 74.28 106.31 75.33 (183.10) (185.84) 84.13
bak_cyrl 79.24 108.34 (259.93) (272.19) (209.63) 92.72
bam_latn 158.88 175.25 188.31 (212.45) 203.82 143.14
ban_latn 121.16 137.57 (154.09) (183.02) (188.39) 114.41
bel_cyrl 79.50 124.53 81.13 (226.99) (211.85) 91.22
bem_latn 150.39 174.20 (218.51) (188.14) (237.98) 158.61
ben_beng 80.71 109.08 90.55 79.78 76.54 94.62
bho_deva 121.88 138.45 (168.84) (149.88) (120.97) (130.48)
bjn_arab (297.57) (383.82) (260.05) (261.94) (238.42) (257.15)
bjn_latn 111.90 140.40 (151.76) (150.82) (153.76) 104.37
bod_tibt 118.58 142.65 (134.47) (137.26) (206.06) 120.94
bos_latn 73.84 113.42 63.67 (155.27) (135.77) 67.75
bug_latn 172.63 181.80 (206.57) (214.07) (218.78) (211.21)
bul_cyrl 71.36 109.94 64.51 59.03 (148.72) 70.69
cat_latn 76.30 115.98 65.93 60.65 60.55 66.57
ceb_latn 94.04 125.05 (111.22) (192.82) (171.75) 103.37
ces_latn 75.07 115.11 63.68 (154.38) (140.67) 71.35
cjk_latn 219.90 239.42 (212.05) (211.78) (220.85) 189.97
ckb_arab 89.11 121.96 (112.62) (290.64) (199.22) 107.81
crh_latn 105.56 127.10 (175.84) (185.40) (180.14) 119.20
cym_latn 77.50 114.42 97.27 (226.16) (204.15) 100.22
dan_latn 74.09 111.50 60.26 (122.87) (129.58) 67.08
deu_latn 73.91 118.06 57.93 57.08 (98.55) 62.85
dik_latn 152.14 169.32 (197.55) (202.82) (216.99) (225.78)
dyu_latn 183.05 210.69 (209.72) (216.63) (209.24) 189.79
dzo_tibt 125.44 144.63 (213.72) (217.28) (238.66) 110.58
ell_grek 78.38 119.29 68.55 65.99 (157.23) 93.41
eng_latn 68.73 103.16 51.10 50.39 50.56 48.43
epo_latn 75.57 110.48 118.68 (175.58) (149.97) 88.92
est_latn 73.18 109.48 71.72 60.94 (173.50) 96.32
eus_latn 70.76 103.94 112.97 70.24 70.78 101.32
ewe_latn 128.71 144.86 (181.68) (202.81) (228.49) 137.60
fao_latn 89.22 120.77 (182.64) (216.89) (180.74) 100.02
fij_latn 107.24 126.01 (186.08) (136.05) (215.93) (121.46)
fin_latn 75.34 114.20 63.80 55.50 (164.25) 82.10
fon_latn 190.75 205.52 (209.87) (255.39) 239.02 190.66
fra_latn 70.55 116.07 56.67 55.76 53.43 59.64
fur_latn 114.09 131.26 (203.28) (198.54) (185.86) 107.32
fuv_latn 165.36 188.23 (188.00) (201.33) (192.94) (184.41)
gaz_latn 120.21 202.78 (163.96) (260.02) (262.77) (144.71)
gla_latn 97.03 133.97 172.79 (245.08) (209.10) 120.94
gle_latn 79.48 118.25 150.60 (228.43) (204.19) 106.38
glg_latn 76.34 114.53 86.26 (107.84) (87.42) 74.13
grn_latn 121.55 141.64 193.46 (230.52) (225.65) 119.96
guj_gujr 84.11 110.95 100.35 (276.27) 96.00 97.04
hat_latn 89.82 114.98 114.15 85.51 (163.36) 102.63
hau_latn 86.28 116.50 115.81 (215.68) (215.07) 107.27
heb_hebr 77.17 108.77 73.71 (175.48) (147.78) 111.24
hin_deva 76.64 110.74 79.74 71.13 70.06 88.97
hne_deva 141.43 149.99 (164.54) (153.61) (144.13) 110.53
hrv_latn 71.16 110.07 61.93 (153.35) (135.73) 66.55
hun_latn 75.05 113.40 64.13 (176.50) (164.44) 78.54
hye_armn 78.37 115.07 85.23 (259.42) (224.51) 94.83
ibo_latn 110.84 135.25 147.33 (248.34) 149.46 123.89
ilo_latn 111.06 129.57 (133.69) (227.74) (209.68) 115.75
ind_latn 72.11 101.11 62.46 60.13 59.44 65.23
isl_latn 75.36 113.10 82.56 (200.00) (179.47) 93.17
ita_latn 75.27 116.41 60.06 59.00 (86.91) 62.98
jav_latn 90.19 112.98 102.58 (161.56) (154.48) 102.21
jpn_jpan 68.51 100.99 63.11 61.53 (93.28) 68.24
kab_latn 134.05 154.27 (214.03) (256.69) (215.19) 159.99
kac_latn 128.22 137.82 (187.08) (248.17) (239.55) 146.59
kam_latn (240.45) (264.47) (186.89) (202.02) (200.47) 172.19
kan_knda 76.34 103.10 92.36 (268.95) 93.76 97.03
kas_arab (252.81) (304.08) (276.83) (267.41) (245.94) (260.94)
kas_deva 221.04 240.29 (235.98) (228.86) (231.57) (246.51)
kat_geor 72.44 108.39 82.24 (261.79) (236.64) 96.53
kaz_cyrl 73.52 102.69 76.14 (192.47) (186.61) 89.71
kbp_latn 145.05 159.84 (217.85) (264.27) (286.31) 143.65
kea_latn 145.29 157.19 (180.88) (187.08) (182.18) 135.49
khk_cyrl 77.40 103.60 87.49 (238.21) (192.81) 97.14
khm_khmr 98.82 139.14 114.88 (323.85) (240.85) 117.47
kik_latn 165.85 177.82 (222.18) (244.09) 227.12 148.48
kin_latn 87.00 118.79 (227.82) (225.85) 135.94 113.52
kir_cyrl 70.91 99.21 114.55 (224.69) (202.97) 94.96
kmb_latn 179.18 199.36 (205.24) (196.76) (216.45) 168.66
kmr_latn 99.93 130.74 155.62 (234.52) (213.35) 120.91
knc_arab 181.38 274.16 (223.08) (222.78) (214.43) (228.52)
knc_latn 170.17 206.34 (229.82) (227.51) (239.24) (242.18)
kon_latn 132.91 143.02 182.94 (190.80) (189.74) (126.76)
kor_hang 72.23 102.86 69.83 63.54 (122.61) 73.28
lao_laoo 91.00 120.61 110.10 (268.95) (231.49) 107.59
lij_latn 140.38 168.14 (198.93) (195.96) (193.23) 122.99
lim_latn 123.42 154.99 (180.85) (201.88) (192.90) (116.94)
lin_latn 106.45 122.73 144.57 (183.46) 167.57 116.17
lit_latn 71.55 110.06 67.23 (188.51) (163.54) 92.94
lmo_latn 162.18 201.95 (203.26) (199.20) (198.36) 152.10
ltg_latn 121.88 138.23 (217.38) (241.02) (229.77) (227.25)
ltz_latn 85.00 123.40 (154.55) (149.83) (188.39) 103.80
lua_latn 152.56 162.72 (175.04) (167.40) (201.83) 147.08
lug_latn 118.73 139.82 168.52 (213.70) 165.48 141.39
luo_latn 139.03 155.32 (210.18) (218.56) (222.85) 163.70
lus_latn 95.13 126.88 (157.04) (217.70) (211.72) 131.48
lvs_latn 70.94 110.34 70.45 (187.39) (179.57) 86.42
mag_deva 126.54 139.42 (156.88) (143.63) (128.48) (133.86)
mai_deva 123.50 142.56 (174.69) (165.19) (126.39) 102.51
mal_mlym 80.56 111.47 92.78 (221.87) 88.15 99.60
mar_deva 82.32 110.97 94.21 (200.05) 92.70 100.33
min_arab (308.31) (399.00) (269.74) (273.93) (254.70) (275.42)
min_latn 108.38 128.37 (167.37) (164.01) (160.32) 125.86
mkd_cyrl 73.08 109.43 73.22 (134.89) (162.92) 76.94
mlt_latn 83.60 125.90 (280.75) (279.00) (237.59) 90.70
mni_beng (176.58) (274.13) (279.36) (271.00) (188.18) (275.78)
mos_latn 187.64 198.01 (228.35) (236.65) (241.18) 188.11
mri_latn 97.39 130.22 (191.98) (187.99) (180.89) 109.67
mya_mymr 86.45 125.64 119.52 90.49 (224.44) 121.00
nld_latn 71.72 112.22 60.19 (111.54) (115.43) 65.13
nno_latn 80.80 114.77 (73.23) (153.12) (151.17) 76.16
nob_latn 76.13 109.96 64.70 (122.09) (131.71) 68.04
npi_deva 82.42 111.11 90.72 (193.40) 86.21 86.53
nso_latn 119.03 139.17 (166.19) (234.76) 181.13 123.81
nus_latn (217.98) (276.12) (259.51) (266.18) (293.48) (319.71)
nya_latn 97.55 121.06 (213.77) (202.58) 180.39 118.77
oci_latn 97.75 135.49 (152.88) (126.52) (129.14) 93.79
ory_orya 87.64 114.38 (113.19) (387.80) (103.65) 105.48
pag_latn 122.99 139.15 (175.08) (184.33) (184.39) 120.19
pan_guru 85.22 116.57 107.69 (298.41) 99.43 101.39
pap_latn 94.63 126.00 (166.14) (177.79) (183.93) 118.24
pbt_arab 104.87 136.61 120.56 (263.66) (210.05) 130.62
pes_arab 75.11 112.64 70.67 (152.72) (145.67) 84.23
plt_latn 87.89 121.62 116.37 (236.48) (176.35) 102.04
pol_latn 72.87 114.38 61.18 (146.68) (130.92) 69.19
por_latn 72.38 110.34 59.80 57.79 55.73 63.06
prs_arab 96.31 120.44 (80.74) (150.59) (141.22) 83.70
quy_latn 121.48 144.34 185.50 125.42 (196.23) 132.70
ron_latn 75.68 118.45 62.78 (151.24) (135.97) 70.39
run_latn 112.80 135.64 (229.50) (226.86) 152.96 127.91
rus_cyrl 73.59 117.13 58.22 57.38 (110.95) 65.18
sag_latn 162.70 167.88 (182.49) (171.77) (196.64) 150.49
san_deva 134.36 156.91 167.12 (188.71) (167.33) 140.17
sat_olck 148.03 156.26 (217.91) (217.35) (298.99) 124.11
scn_latn 124.37 157.13 (172.60) (187.74) (175.53) 108.90
shn_mymr 162.52 182.33 (253.11) (475.53) (373.38) (341.88)
sin_sinh 83.19 114.20 97.91 (304.98) (235.47) 106.82
slk_latn 72.22 113.10 63.60 (181.84) (157.25) 78.82
slv_latn 71.40 111.05 64.99 (176.07) (147.81) 75.62
smo_latn 103.20 142.22 (197.35) (213.38) (214.55) 120.15
sna_latn 93.53 123.08 (228.56) (225.74) 180.82 125.19
snd_arab 91.04 116.69 150.37 (257.35) (212.74) 117.36
som_latn 104.72 138.92 125.40 (245.50) (229.18) 135.24
sot_latn 99.85 142.34 (171.86) (235.46) 192.76 122.30
spa_latn 77.45 116.81 63.10 61.91 59.82 66.71
srd_latn 116.75 135.26 (202.61) (200.61) (191.63) 106.39
srp_cyrl 76.66 116.32 72.33 (175.02) (149.07) 71.86
ssw_latn 124.51 143.31 186.61 (232.90) (220.26) 133.78
sun_latn 91.07 116.80 109.40 (167.76) (162.37) 98.90
swe_latn 73.76 109.95 62.27 (106.34) (126.73) 65.90
swh_latn 79.45 109.31 89.60 76.81 98.22 95.63
szl_latn 131.79 157.70 (184.61) (229.41) (208.14) 122.68
tam_taml 79.87 107.64 88.40 78.39 79.09 95.70
taq_latn (241.82) (267.82) (216.97) (231.56) (222.07) (225.98)
taq_tfng None None (301.75) (289.98) (261.59) (396.24)
tat_cyrl 76.67 107.20 (222.35) (227.31) (192.83) 88.82
tel_telu 80.76 106.79 89.99 77.70 88.52 95.52
tgk_cyrl 79.61 113.99 (267.56) (281.19) (213.45) 100.89
tgl_latn 86.11 123.04 91.13 (168.86) (162.76) 89.23
tha_thai 73.84 102.82 71.03 64.31 (177.26) 87.04
tir_ethi 107.10 134.77 142.22 (300.08) (258.62) 133.08
tpi_latn 141.91 164.61 (218.27) (212.39) (197.37) 109.12
tsn_latn 111.75 143.35 184.89 (240.07) 186.95 127.47
tso_latn 111.93 133.30 (236.13) (239.86) 205.36 130.50
tuk_latn 81.23 107.98 (239.28) (256.07) (198.90) 114.93
tum_latn 132.63 154.17 (233.00) (187.16) 237.42 141.81
tur_latn 69.75 100.52 64.44 61.13 (130.11) 86.26
twi_latn 131.51 148.42 (211.42) (216.89) 174.63 128.30
tzm_tfng None None (206.77) (206.42) (243.59) (332.15)
uig_arab 75.68 105.62 (317.09) (367.46) (206.46) 105.21
ukr_cyrl 76.60 116.07 64.77 (129.73) (145.72) 72.19
umb_latn 182.34 211.34 (199.72) (209.91) (221.59) 174.09
urd_arab 83.96 117.74 90.05 79.04 85.58 98.80
uzn_latn 71.09 107.26 112.09 (243.88) (217.18) 96.67
vec_latn 114.88 147.99 (161.71) (155.44) (160.51) 108.87
vie_latn 77.66 120.08 68.06 64.89 61.00 74.36
war_latn 118.02 153.85 (161.42) (203.73) (174.18) 132.17
wol_latn 141.12 158.76 202.72 (225.23) 167.95 161.96
xho_latn 93.68 121.39 144.06 (216.55) 155.76 122.42
ydd_hebr 109.90 144.63 (260.53) (286.49) (210.28) 128.80
yor_latn 123.24 167.30 174.33 (246.05) 154.45 148.67
yue_hant 90.03 121.49 (70.31) (86.41) 61.81 69.76
zho_hans 78.92 121.82 66.34 65.42 59.08 70.09
zho_hant 93.56 125.41 (72.53) (89.36) 63.06 (75.40)
zsm_latn 73.25 100.45 67.03 (83.84) (76.22) 72.90
zul_latn 87.90 118.40 135.47 (218.33) 186.47 115.47
Table 6: Goldfish languages with corresponding dataset sizes.
Language   Language Script Byte Scaled Tokens Dataset Proportions
(ISO 639-3) (ISO 15924) Premium MB
  OSCAR
  NLLB
  MADLAD-400
  Glot500
  Other
Afrikaans afr latn 1.04 1000.00 239682048
Amharic amh ethi 1.72 1000.00 211767808
Standard Arabic arb arab 1.47 1000.00 196197376
Azerbaijani aze latn 1.30 1000.00 233091584
Belarusian bel cyrl 2.01 1000.00 254138368
Bengali ben beng 2.43 1000.00 194737152
Bosnian bos cyrl 1.15 1000.00 232501760
Bosnian bos latn 0.97 1000.00 228266496
Bulgarian bul cyrl 1.81 1000.00 224346112
Catalan cat latn 1.09 1000.00 238915072
Czech ces latn 1.04 1000.00 206113280
Welsh cym latn 1.03 1000.00 236230144
Danish dan latn 1.02 1000.00 208085504
German deu latn 1.05 1000.00 210817024
Modern Greek ell grek 1.97 1000.00 238704128
English eng latn 1.00 1000.00 213977088
Esperanto epo latn 1.00 1000.00 231384576
Estonian est latn 0.97 1000.00 189518336
Basque eus latn 1.06 1000.00 209921536
Persian fas arab 1.59 1000.00 244359680
Filipino fil latn 1.33 1000.00 274955776
Finnish fin latn 1.06 1000.00 186050560
French fra latn 1.17 1000.00 251415552
Galician glg latn 1.06 1000.00 222080000
Gujarati guj gujr 2.16 1000.00 193794560
Hausa hau latn 1.18 1000.00 277416448
Hebrew heb hebr 1.36 1000.00 192904704
Hindi hin deva 2.37 1000.00 228020736
Croatian hrv latn 0.99 1000.00 219422208
Hungarian hun latn 1.02 1000.00 191089664
Armenian hye armn 1.72 1000.00 203630592
Indonesian ind latn 1.18 1000.00 210432000
Icelandic isl latn 1.15 1000.00 236872704
Italian ita latn 1.07 1000.00 216099840
Japanese jpn jpan 1.32 1000.00 219063296
Kara-Kalpak kaa cyrl 1.92 1000.00 212100608
Kannada kan knda 2.64 1000.00 212683264
Georgian kat geor 4.34 1000.00 354762752
Kazakh kaz cyrl 1.76 1000.00 199970304
Kirghiz kir cyrl 1.96 1000.00 223066112
Korean kor hang 1.29 1000.00 227021824
Latin lat latn 0.88 1000.00 188774912
Latvian lav latn 1.29 1000.00 243401728
Lithuanian lit latn 1.03 1000.00 201228800
Malayalam mal mlym 2.88 1000.00 244708864
Marathi mar deva 2.48 1000.00 206630400
Macedonian mkd cyrl 1.83 1000.00 221346304
Maltese mlt latn 1.09 1000.00 283158528
Mongolian mon cyrl 1.78 1000.00 205737472
Malay msa latn 1.29 1000.00 236371456
Nepali nep deva 2.63 1000.00 215368192
Dutch nld latn 1.05 1000.00 216978432
Norwegian Bokmål nob latn 1.00 1000.00 205949952
Norwegian nor latn 1.13 1000.00 255482880
Panjabi pan guru 2.22 1000.00 215775232
Iranian Persian pes arab 1.60 1000.00 215946240
Polish pol latn 1.08 1000.00 216235008
Portuguese por latn 1.10 1000.00 225242112
Pushto pus arab 1.59 1000.00 237871616
Romanian ron latn 1.12 1000.00 230580224
Russian rus cyrl 1.82 1000.00 220467712
Sinhala sin sinh 2.45 1000.00 233098752
Slovak slk latn 1.04 1000.00 211206144
Slovenian slv latn 0.97 1000.00 198052864
Somali som latn 1.42 1000.00 302652928
Spanish spa latn 1.08 1000.00 221790720
Albanian sqi latn 1.34 1000.00 274664448
Serbian srp cyrl 1.42 1000.00 184423424
Serbian srp latn 0.83 1000.00 207482368
Swahili swa latn 1.26 1000.00 260033024
Swedish swe latn 1.02 1000.00 206359552
Tamil tam taml 2.73 1000.00 200523264
Tatar tat cyrl 1.85 1000.00 232933888
Telugu tel telu 2.62 1000.00 209365504
Tajik tgk cyrl 1.75 1000.00 216990208
Tagalog tgl latn 1.12 1000.00 245370880
Thai tha thai 2.74 1000.00 205872640
Turkish tur latn 1.04 1000.00 186848768
Ukrainian ukr cyrl 1.75 1000.00 215392768
Urdu urd arab 1.71 1000.00 247899648
Uzbek uzb latn 1.23 1000.00 261058560
Vietnamese vie latn 1.35 1000.00 262306304
Chinese zho hans 0.94 1000.00 206204416
Irish gle latn 1.98 976.70 404823040
Kurdish kur arab 1.57 902.39 196483584
Standard Malay zsm latn 1.14 859.52 185929728
Central Kurdish ckb arab 1.65 838.87 190565888
Kinyarwanda kin latn 1.13 810.96 193561088
Haitian hat latn 0.97 775.80 185333248
Odia ori orya 2.60 774.55 165528576
Zulu zul latn 1.16 764.14 199965696
Burmese mya mymr 5.00 762.14 315374592
Central Khmer khm khmr 3.90 742.37 235559424
Malagasy mlg latn 1.27 720.80 210497024
Kurdish kur latn 1.29 685.53 189872128
Dhivehi div thaa 2.00 634.02 114510336
Shona sna latn 1.12 608.11 151712256
Luxembourgish ltz latn 1.23 579.07 160200192
Sundanese sun latn 1.10 577.96 142266368
Scottish Gaelic gla latn 0.99 558.84 123736064
Cebuano ceb latn 1.11 540.21 140301312
Lao lao laoo 2.71 532.98 124077056
Uzbek uzb cyrl 1.98 525.51 110868992
Yoruba yor latn 1.37 502.55 155829248
Norwegian Nynorsk nno latn 1.03 498.93 116016128
Xhosa xho latn 1.20 477.36 127885824
Western Frisian fry latn 1.23 472.81 133072384
Javanese jav latn 1.15 465.58 115332096
Sindhi snd arab 1.59 459.14 114626048
Maori mri latn 1.18 450.17 136011776
Yiddish yid hebr 1.55 446.04 85695488
Nyanja nya latn 1.21 444.13 112440832
Corsican cos latn 1.18 414.00 126150656
Faroese fao latn 1.16 400.34 96587776
Bashkir bak cyrl 2.27 398.36 118369280
Uighur uig arab 2.31 397.21 104039936
Igbo ibo latn 1.35 388.31 119706112
Modern Greek ell latn 1.24 376.42 92225536
Occitan oci latn 1.01 375.38 99783680
Plateau Malagasy plt latn 1.15 370.58 97517568
Assamese asm beng 2.53 348.88 77216256
Hmong hmn latn 1.19 345.97 100051968
Tosk Albanian als latn 1.17 336.30 87609344
Southern Sotho sot latn 1.17 332.91 94144000
Samoan smo latn 1.18 314.93 101910016
Azerbaijani aze arab 1.20 267.26 56526848
Hawaiian haw latn 1.11 260.95 86747136
Chuvash chv cyrl 1.80 256.36 84293120
Papiamento pap latn 1.00 255.51 60037632
Tigrinya tir ethi 1.76 252.98 56515072
Asturian ast latn 1.75 225.68 93333504
Southern Pashto pbt arab 1.74 225.11 60608000
Central Kanuri knc arab 2.50 221.65 237422592
Lushai lus latn 1.17 213.03 62735360
Northern Uzbek uzn cyrl 2.01 208.92 44960768
Yakut sah cyrl 1.88 206.06 47289344
Ancient Greek grc grek 1.77 205.45 47620608
Turkmen tuk latn 1.79 186.44 57201664
Chinese zho hant 0.99 177.32 42692096
Waray war latn 1.09 175.25 48998912
Kara-Kalpak kaa latn 1.23 165.22 38767104
Breton bre latn 1.01 163.11 43437056
Dari prs arab 1.66 162.70 37549568
Venetian vec latn 1.00 150.70 40523776
North Azerbaijani azj latn 1.08 149.82 27041792
Northern Uzbek uzn latn 1.65 145.59 52049408
Limburgan lim latn 1.00 142.31 39700480
Kalaallisut kal latn 1.34 140.44 30082048
Quechua que latn 1.21 139.38 40595968
Oromo orm latn 1.26 137.90 39742976
Ganda lug latn 1.22 132.42 37459968
Tibetan bod tibt 2.62 131.94 23463424
Hindi hin latn 1.26 131.86 37683712
Swiss German gsw latn 1.14 128.81 38605824
Ayacucho Quechua quy latn 1.16 123.58 34850816
Lombard lmo latn 0.94 123.24 35603456
Egyptian Arabic arz arab 1.55 122.38 30322176
Western Panjabi pnb arab 1.41 121.58 30110208
Eastern Yiddish ydd hebr 1.81 120.20 28306432
Sanskrit san deva 2.54 119.34 31856128
Sicilian scn latn 1.04 113.80 32010752
Halh Mongolian khk cyrl 1.80 108.25 23605760
South Azerbaijani azb arab 1.49 107.56 26922496
Walloon wln latn 1.22 102.32 29091328
Tswana tsn latn 1.17 101.85 31488512
Gujarati guj latn 1.19 101.60 24635392
Gilaki glk arab 1.68 98.73 25519104
Iloko ilo latn 1.08 97.44 25450496
Tetum tet latn 1.40 96.03 28032512
Banjar bjn latn 1.17 93.17 25012224
Rundi run latn 1.12 90.59 23721984
Romansh roh latn 1.27 86.73 23623680
Chechen che cyrl 1.83 86.11 23590400
West Central Oromo gaz latn 1.33 79.04 25565184
Yue Chinese yue hant 0.86 78.42 16084992
Low German nds latn 1.14 75.35 20312064
Minangkabau min latn 0.95 75.07 17732608
Inuktitut iku cans 2.16 74.41 13798400
Tsonga tso latn 1.21 71.85 21684224
Achinese ace latn 1.24 71.09 21666816
Tuvinian tyv cyrl 1.86 68.39 15576576
Northern Sami sme latn 1.27 66.64 15802880
Ewe ewe latn 1.08 63.27 18470400
Twi twi latn 1.03 62.79 18900480
Standard Estonian ekk latn 0.99 61.41 12375552
Guarani grn latn 0.99 60.38 15366656
Pedi nso latn 1.12 59.40 17516544
Northern Kurdish kmr latn 1.03 53.71 12299264
Udmurt udm cyrl 1.74 51.77 10932736
Akan aka latn 1.57 49.51 22551040
Mari (Russia) chm cyrl 1.76 49.43 11290624
Mongolian mon latn 1.18 49.21 12692480
Lingala lin latn 1.14 47.33 13213184
Crimean Tatar crh latn 1.31 47.20 12994560
Zaza zza latn 1.20 46.78 14813184
Kabyle kab latn 1.03 45.19 14035456
Min Nan Chinese nan latn 1.15 44.38 16624128
Scots sco latn 1.19 42.97 12578304
Aragonese arg latn 1.19 42.82 12469760
Maithili mai deva 2.39 41.73 11159040
Fon fon latn 1.54 40.84 13993984
Buriat bua cyrl 1.70 39.10 8951808
Ossetian oss cyrl 1.85 38.60 14059008
Pampanga pam latn 1.19 38.14 11270656
Dimli diq latn 0.96 37.98 9935872
Wolof wol latn 1.08 37.32 12005888
Tedim Chin ctd latn 1.30 37.10 11405824
Tumbuka tum latn 1.21 36.69 9842688
Pangasinan pag latn 1.04 36.43 10441728
Fijian fij latn 1.21 35.48 10642944
Standard Latvian lvs latn 1.21 35.42 8333312
Bemba bem latn 1.16 35.35 10177024
Kabardian kbd cyrl 1.78 34.89 9802752
Luo luo latn 1.04 34.50 9859072
Hakha Chin cnh latn 1.32 33.20 10364928
Hiligaynon hil latn 1.35 32.12 9034752
Balinese ban latn 1.27 31.84 9161216
Aymara aym latn 1.21 30.74 9201152
Avaric ava cyrl 1.94 30.73 8009728
Central Aymara ayr latn 1.10 28.37 7641088
Fiji Hindi hif latn 1.28 28.00 8768000
Ligurian lij latn 1.14 27.89 8498176
Eastern Mari mhr cyrl 1.81 27.86 6580224
Bavarian bar latn 1.13 27.68 7961600
Silesian szl latn 1.07 27.04 7593472
Russian rus latn 1.18 26.62 7373824
Ido ido latn 1.18 26.18 7369216
Russia Buriat bxr cyrl 1.59 25.38 6060544
Abkhazian abk cyrl 2.01 25.24 6408192
Sardinian srd latn 1.11 24.71 6834176
Nigerian Pidgin pcm latn 0.95 24.62 5281280
Wu Chinese wuu hani 0.70 24.53 4112384
Fulah ful latn 1.26 24.03 7806464
Bhojpuri bho deva 2.52 23.74 6156800
Betawi bew cyrl 1.74 23.52 5288960
Volapük vol latn 1.13 21.39 6030336
Nigerian Fulfulde fuv latn 1.11 21.23 6159872
Karachay-Balkar krc cyrl 1.87 21.02 4627456
Swati ssw latn 1.14 20.97 5566976
Luba-Lulua lua latn 1.19 20.82 6322688
Friulian fur latn 1.07 20.72 5487616
Khasi kha latn 1.30 20.56 6209536
Telugu tel latn 1.28 20.02 5266432
Iban iba latn 1.30 19.98 5278208
Bikol bik latn 1.27 19.26 5440512
Interlingua ina latn 1.24 19.15 5581824
Latgalian ltg latn 1.00 18.70 4046848
Komi kom cyrl 1.61 18.20 4716032
Querétaro Otomi otq latn 1.25 17.48 5702656
Tonga (Tonga Islands) ton latn 1.27 17.46 6237184
Azerbaijani aze cyrl 1.82 17.12 3627008
Dargwa dar cyrl 2.02 16.99 4506624
Erzya myv cyrl 1.77 16.81 3851776
Piemontese pms latn 1.23 16.75 5307904
Tok Pisin tpi latn 1.18 16.61 5102592
Umbundu umb latn 1.17 16.12 4743168
Sango sag latn 1.16 15.87 4929024
Kabuverdianu kea latn 0.78 15.74 3247616
Adyghe ady cyrl 1.81 15.22 4124160
Literary Chinese lzh hant 0.70 15.20 2767872
Gulf Arabic afb arab 1.37 14.25 3247616
Falam Chin cfm latn 1.32 14.09 4315648
Kabiyè kbp latn 1.44 13.93 4698624
Bambara bam latn 1.26 12.84 4511744
Kachin kac latn 1.35 12.74 4453888
Newari new deva 2.56 12.44 2927616
Syriac syr syrc 1.41 12.17 2641408
Chokwe cjk latn 1.17 12.10 3622400
Dyula dyu latn 1.15 11.94 3849216
Betawi bew latn 1.30 11.84 3186176
Venda ven latn 1.30 11.82 3268608
Dinka din latn 1.24 11.69 4125696
Shan shn mymr 2.82 11.66 2238976
Southern Altai alt cyrl 1.86 11.65 2694144
Southwestern Dinka dik latn 1.12 11.61 3753984
Goan Konkani gom deva 1.74 11.50 2219520
Sranan Tongo srn latn 1.06 11.47 3098112
Yucateco yua latn 1.24 11.41 3645440
Kongo kon latn 1.23 11.32 3549184
Kimbundu kmb latn 1.13 11.09 3359744
Kumyk kum cyrl 1.96 11.04 2208768
Buginese bug latn 1.23 10.72 3269632
Goan Konkani gom latn 1.21 10.38 2806784
Mossi mos latn 1.14 10.37 3537920
Upper Sorbian hsb latn 1.12 10.31 2503680
Lak lbe cyrl 2.01 10.24 2470912
North Ndebele nde latn 0.97 10.17 1766912
Central Kanuri knc latn 1.18 10.07 3433472
Ingush inh cyrl 1.70 9.59 2764800
Zapotec zap latn 1.08 9.58 2395136
Central Bikol bcl latn 1.22 9.49 2638336
Lezghian lez cyrl 1.83 9.38 2358784
Kituba mkw cyrl 1.81 9.37 2266112
Cusco Quechua quz latn 1.30 9.32 2070528
Bishnupriya bpy beng 2.33 9.29 2019328
Mam mam latn 1.34 9.27 3580416
Magahi mag deva 2.56 9.08 2488832
Tzotzil tzo latn 1.49 9.02 3463680
Tamil tam latn 1.27 9.00 2260992
Western Mari mrj cyrl 1.51 8.74 1812992
Brunei Bisaya bsb latn 1.31 8.69 2460672
Chhattisgarhi hne deva 2.17 8.61 2106880
Luba-Katanga lub latn 1.30 8.61 2269184
Kaqchikel cak latn 1.82 8.51 4157952
Santali sat olck 2.80 8.49 2224128
Vlaams vls latn 1.21 8.49 2484736
Kikuyu kik latn 1.29 8.36 2418176
Mirandese mwl latn 1.24 8.12 2293760
Isoko iso latn 1.48 8.11 2638336
Uighur uig latn 1.19 7.88 1662976
Dzongkha dzo tibt 3.26 7.70 2019328
Bashkir bak latn 1.19 7.53 1793024
Dombe dov latn 0.99 7.43 1389056
Madurese mad latn 1.29 7.29 2044416
Levantine Arabic apc arab 1.47 7.06 1687040
Pohnpeian pon latn 0.90 7.02 1412608
Kashmiri kas deva 2.53 6.96 1990656
Paite Chin pck latn 1.32 6.94 2163712
Veps vep latn 1.17 6.89 1751552
Boko (Benin) bqc latn 0.98 6.80 1806336
Neapolitan nap latn 1.23 6.73 2123776
Manx glv latn 1.22 6.63 1939968
Nande nnb latn 1.31 6.49 1764352
Batak Toba bbc latn 1.33 6.48 1846784
Malayalam mal latn 1.27 6.38 1556480
Tiv tiv latn 1.31 6.32 2119168
Cornish cor latn 1.22 6.31 1936896
Khakas kjh cyrl 1.93 6.17 1271808
Moksha mdf cyrl 1.71 6.17 1302016
Kalmyk xal cyrl 1.72 6.05 1474048
Guerrero Nahuatl ngu latn 1.44 5.99 1508864
Klingon tlh latn 1.14 5.91 1741312
Crimean Tatar crh cyrl 1.89 5.86 1265664
Makhuwa-Meetto mgh latn 1.11 5.77 1251328
Sanskrit san latn 0.97 5.72 1164800
Northern Frisian frr latn 1.17 5.68 1594368
Eastern Balochi bgp latn 1.29 5.64 1735680
Carpathian Romani rmc latn 1.02 5.61 1241600
Georgian kat latn 1.20 5.57 1422336
Old English ang latn 1.29 5.47 1671168
Kedah Malay meo latn 1.28 5.44 1670656
Mingrelian xmf geor 2.51 5.44 1367040
Tulu tcy knda 2.67 5.29 1210368
Tandroy-Mahafaly Malagasy tdx latn 1.00 5.23 1303552
Komi-Zyrian kpv cyrl 1.67 5.19 1355776
Lingua Franca Nova lfn latn 1.30 5.12 1593344
Ditammari tbz latn 1.33 5.12 1868800
Nzima nzi latn 1.42 5.07 1514496
Rusyn rue cyrl 1.56 5.03 1160704
Eastern Huasteca Nahuatl nhe latn 1.49 5.02 1268224

### v2

@@
 For many low-resource languages, the only available language models are large multilingual models trained on many languages simultaneously.
-However, using FLORES perplexity as a metric, we find that these models perform worse than bigrams for many languages (e.g.
-24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B).
-To facilitate research that focuses on low-resource languages, we pre-train and release Goldfish, a suite of monolingual autoregressive Transformer language models up to 125M parameters for 350 languages.
-The Goldfish reach lower FLORES perplexities than BLOOM, XGLM, and MaLA-500 on 98 of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller.
-However, the Goldfish significantly underperform larger multilingual models on reasoning benchmarks, suggesting that for low-resource languages, multilinguality primarily improves general reasoning abilities rather than basic text generation.
-We release models trained on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) of text data where available.
-The Goldfish models are available as baselines, fine-tuning sources, or augmentations to existing models in low-resource NLP research, and they are further useful for crosslinguistic studies requiring maximally comparable models across languages.
+Despite state-of-the-art performance on reasoning tasks, we find that these models still struggle with basic grammatical text generation in many languages.
+First, large multilingual models perform worse than bigrams for many languages (e.g.
+24% of languages in XGLM 4.5B; 43% in BLOOM 7.1B) using FLORES perplexity as an evaluation metric.
+Second, when we train small monolingual models with only 125M parameters on 1GB or less data for 350 languages, these small models outperform large multilingual models both in perplexity and on a massively multilingual grammaticality benchmark.
+To facilitate future work on low-resource language modeling, we release Goldfish, a suite of over 1,000 small monolingual language models trained comparably for 350 languages.
+These models represent the first publicly-available monolingual language models for 215 of the languages included.
+Keywords: multilingual NLP, low-resource languages
+[Uncaptioned image]  Models and training data       Training and evaluation code
+[…]
 Goldfish: Monolingual Language Models for 350 Languages
+Tyler A. Chang1, Catherine Arnett2, Zhuowen Tu1,3, Benjamin K. Bergen1
+1Department of Cognitive Science, 2Department of Linguistics,
+3Department of Computer Science,
+University of California San Diego
+Abstract content
+1.  Introduction
+Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020b; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023).
+For many low-resource languages, a dedicated model optimized for that language does not exist.
+For example, in an analysis of publicly available models on Hugging Face, we find that 215 of the 350 languages in this paper did not have a monolingual text-generation model prior to the Goldfish models, and 47 of the languages did not have any text-generation models at all (Arnett and Chang, 2025).
+This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022).
+These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020a).
+In particular, we find in this paper that large multilingual language models still struggle with basic next token prediction—a prerequisite for text generation—in low-resource languages.
+To establish a baseline for basic text generation performance in low-resource languages, we introduce Goldfish [^1] This name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024).
+If an acronym is desired, Goldfish can stand for Generative autOregressive Low-resource moDels For lImited-compute System Hardware., a suite of over 1000 monolingual language models for 350 diverse languages.
 […]
-1 Introduction
-Language modeling research in low-resource languages often relies on large multilingual models trained on many languages simultaneously (Conneau et al., 2020; Adelani et al., 2021b; Ebrahimi et al., 2022; Lin et al., 2022; Hangya et al., 2022; Imani et al., 2023).
-For many low-resource languages, a dedicated model optimized for that language does not exist.
-This lack of dedicated models hinders comparability of results across models and languages (Bandarkar et al., 2024), and it contributes to model under-performance in low-resource languages (Wu and Dredze, 2020; Blasi et al., 2022).
-These barriers to research in low-resource languages are likely to exacerbate existing inequities across language communities in NLP research (Bender, 2011; Joshi et al., 2020).
-To address this lack of available models, we introduce Goldfish, a suite of over 1000 monolingual language models for 350 diverse languages.
-[^1] The name refers to shared qualities between our models and goldfish (Carassius auratus); they are small, there are many of them, and they are known for their poor memories (perhaps inaccurately; Carey, 2024).
-The models reach lower perplexities than XGLM (Lin et al., 2022), BLOOM 7.1B (Scao et al., 2022), and MaLA-500 (Lin et al., 2024) on 98 out of 204 FLORES languages, despite each Goldfish model being over 10$\times$ smaller.
-The Goldfish also outperform simple bigram models, which are surprisingly competitive with larger models for low-resource languages (e.g.
-lower perplexities than BLOOM 7.1B on 43% of its languages; §4).
-However, despite better perplexities, the Goldfish underperform larger multilingual models on reasoning benchmarks, suggesting that multilingual pre-training may benefit abstract reasoning capabilities over more basic grammatical text generation (§5).
-Data size Model output 5MB ‘‘Goldfish are a few years of the most of the most of the most...’’ 10MB ‘‘Goldfish are a great way to the best way to the best way...’’ 100MB ‘‘Goldfish are a great way to get your fish in the wild.’’ 1GB ‘‘Goldfish are a species of fish that are found in the sea.’’
+Thus, the primary contributions of this short paper are (1) to show that large multilingual models still suffer from poor next token prediction performance in low-resource languages, (2) to demonstrate that small monolingual models often exhibit lower perplexities and more grammatical next token predictions for such languages, and (3) to address the lack of available models in low-resource languages by releasing comparable monolingual models for 350 languages.
+Models, training data, and code are available at: https://huggingface.co/goldfish-models.
+Data size Model output
+5MB Goldfish are a few years of the most of the most of the most...
+10MB Goldfish are a great way to the best way to the best way...
+100MB Goldfish are a great way to get your fish in the wild.
+1GB Goldfish are a species of fish that are found in the sea.
 Figure 1: Left: Map of the 350 languages for which Goldfish models are available, using coordinates from Glottolog (Hammarström et al., 2023).
+Color indicates the largest dataset size for that language.
 Right: Sample model outputs completing the prompt ‘‘Goldfish are’’ for the eng_latn (English) model for each dataset size, using sampling temperature zero.
-Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4).
-Finally, to enable comparisons across languages, we release monolingual models trained on comparable dataset sizes for all languages: 5MB, 10MB, 100MB, and 1GB when available, after accounting for the fact that languages require different numbers of UTF-8 bytes to encode comparable content (Arnett et al., 2024).
-These Goldfish serve as baselines, allowing results in diverse languages to be situated relative to comparable models.
-They can also be used as source models for fine-tuning or to enhance larger multilingual models in areas where those models fall short (§4).
-Models and code are available at https://huggingface.co/goldfish-models.
-2 Related Work
-Low resource language modeling often leverages multilingual pre-training, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020).
-Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2023).
+Grammatical text generation begins to emerge in the 100MB-dataset model (available for 166 languages), but the lower-resource models still achieve better perplexities than previous models for many low-resource languages (§4.1).
+2.  Related Work
+Low resource language modeling often leverages multilingual pretraining, where a model is trained on multiple languages simultaneously (Pires et al., 2019; Conneau et al., 2020b).
+Indeed, this can improve low-resource performance, particularly when models have sufficient capacity and the multilingual data is from related or typologically similar languages (Kakwani et al., 2020; Ogueji et al., 2021; Chang et al., 2024a).
 However, monolingual models have still been shown to achieve better performance than multilingual models for many languages (e.g.
 Martin et al., 2020; Pyysalo et al., 2021; Gutiérrez-Fandiño et al., 2021; Luukkonen et al., 2023).
-Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020; Chang et al., 2023).
-Notably, the training datasets for massively multilingual models are often heavily skewed towards high-resource languages.
-For example, XGLM 4.5B is trained on over 7000$\times$ more Norwegian (71GB; 5.4M native speakers) than Quechua (0.01GB; 7.3M native speakers; Lin et al., 2022; Ethnologue, 2024).
-In a more extreme case, BLOOM is trained on only 0.07MB of Akan (8.1M native speakers) out of 1.61TB total (4e-6% of the pre-training dataset; Scao et al., 2022).
-These extremely small quantities of low-resource language data often do not leverage recent efforts to compile text data in low-resource languages (Costa-jussà et al., 2022; Imani et al., 2023; Kudugunta et al., 2023), and the data imbalances are likely to severely hinder performance in low-resource languages.
-Indeed, we find that these models have worse perplexities than simple bigram models for many languages (§4).
-Unfortunately, comparable monolingual language models across many diverse languages have yet to be studied or released.
-3 Models and Datasets
-We introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pre-trained for 350 languages.
+Thus, it appears that existing multilingual language models are still limited by model capacity or limited data in low-resource languages (Conneau et al., 2020b; Chang et al., 2024a).
+[…]
+3.  Models and Datasets
+To compare to larger multilingual models, we introduce the Goldfish models, a suite of 1154 monolingual Transformer language models pretrained for 350 languages.
 The largest model for each language is 125M parameters.
@@
 Figure 1 shows a geographic map of the 350 languages, with coordinates from Glottolog (Hammarström et al., 2023), along with sample outputs from the English model for each dataset size.
+3.1.  Training Datasets
+We merge the massively multilingual text datasets compiled in Chang et al.
+(2024a), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language.
+We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data.
+Full dataset details are in §A.1.
+To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022).
+We conduct a contamination analysis to determine whether FLORES is contaminated in our datasets and find that for 98% of languages, less than 10 out of 2000 FLORES sequences appear in the training dataset at all (§A.2).
+[…]
+As we cap our dataset sizes at 1GB per language, our models are trained on much less data than is otherwise available for high-resource languages.
+Still, when we compare our dataset sizes to FineWeb2 (Penedo et al., 2025) for the languages covered by both Goldfish and FineWeb2, 106 of those languages have more data in our dataset than in FineWeb2.
+We attribute this partially to our approach to data collection, which involves seeking out language-specific resources rather than relying solely on language identification models, which can be unreliable for web data in low-resource languages (Ortiz Suarez et al., 2026).
+We release our training datasets at: https://huggingface.co/datasets/goldfish-models/fish-food.
 Goldfish data size # Langs Goldfish Bigrams XGLM 4.5B MaLA-500 10B
@@
 10MB, 5MB 5 130.5 148.3 183.1 135.0
-Table 1: Mean FLORES log-perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size.
+Table 1: Mean FLORES perplexity ($\downarrow$) for the 100 languages in XGLM 4.5B, MaLA-500, and FLORES, separated by maximum Goldfish dataset size.
 The Goldfish languages are a strict superset of these languages.
-3.1 Training Datasets
-We merge the massively multilingual text datasets compiled in Chang et al.
-(2023), Glot500 (Imani et al., 2023), and MADLAD-400 (Kudugunta et al., 2023) per language.
-To facilitate fair evaluations, we hold out FLORES-200 and AmericasNLI from all datasets (Costa-jussà et al., 2022; Ebrahimi et al., 2022).
-We deduplicate repeated sequences of 100 UTF-8 bytes and drop languages with only Bible data.
-Full dataset details are in §A.1.
-To sample pre-training datasets of the desired sizes in a language $L$, we first use the Byte Premium Tool (Arnett et al., 2024) to estimate the byte premium for $L$, the number of UTF-8 bytes required to encode comparable text in $L$ relative to eng_latn (English).
-For example, khm_khmr (Khmer) has byte premium 3.91, meaning that it uses approximately 3.91$\times$ as many UTF-8 bytes as English to encode content-matched text.
-We divide each dataset size by the estimated byte premium for the corresponding language, thus measuring all datasets in units of “equivalent” English text bytes.
-We sample datasets to train monolingual language models on 5MB (350 languages), 10MB (288 languages), 100MB (166 languages), and 1GB (83 languages) when available after byte premium scaling.
-[^2] The languages with 5MB-dataset models are a subset of the languages with 10MB-dataset models, and similarly for the 100MB and 1GB dataset sizes.
-These are equivalent to roughly 1M, 2M, 20M, and 200M tokens of English text respectively; including 10 epochs of repetition, the 1GB-dataset models are trained on the equivalent of roughly 2B English tokens.
-When a 1GB dataset is not available for a language after byte premium scaling, we include a full model (267 languages) trained on the entire dataset in that language, for use cases that seek to maximize performance in a specific low-resource language.
-3.2 Architectures and Pre-Training
-For each language and each dataset size, we pre-train an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019).
-For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019).
-Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al.
-(2019) for the 10MB and 5MB dataset sizes.
-Full hyperparameters are reported in §A.2.
-We tokenize each dataset using using a monolingual SentencePiece tokenizer (Kudo and Richardson, 2018) trained on that dataset size, limiting tokenizer training text to 100MB after byte premium scaling.
-Following Liu et al.
-(2019), we use vocabulary size 50K and a maximum sequence length of 512 tokens for all models.
-We train each language model on 10 epochs of its corresponding dataset.
-[^3] Multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but we find that more than 10 epochs of training leads to overfitting for extremely small datasets (e.g.
-5MB).
-Pre-training details, compute costs, and all available models are reported in §A.2.
 Bigrams XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
@@
 For example, Goldfish reach lower log-perplexities than MaLA-500 for 111/175 (63%) of FLORES languages in both Goldfish and MaLA-500.
-# Langs Chance Goldfish XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
-Belebele 121 25.0 28.2 30.1 30.6 30.2 30.6
-XCOPA 11 50.0 54.9 57.9 60.6 56.9 55.6
-XStoryCloze 10 50.0 52.5 57.1 59.9 58.2 55.7
-Table 3: Reasoning benchmark accuracies averaged over non-English languages. Despite better perplexities, the Goldfish perform significantly worse than larger multilingual models on reasoning.
-4 FLORES Log-Perplexity Evaluations
+3.2.  Architectures and Pretraining
+We train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full.
+The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language.
+In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total).
+For each language and each dataset size, we pretrain an autoregressive GPT-2 Transformer language model from scratch (Radford et al., 2019).
+For the 1GB, 100MB, and full dataset sizes, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018), which has a similar parameter count to BERT-base and RoBERTa (Devlin et al., 2019; Liu et al., 2019).
+Because larger models do not appear to outperform smaller models for very small datasets (Chang et al., 2024a), we use the small model size (39M parameters) from Turc et al.
+(2019) for the 10MB and 5MB dataset sizes.
+Full hyperparameters are reported in §A.3.
+Each model has a custom monolingual tokenizer, which is trained with a vocabulary size of 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling).
+We use Unigram tokenizers trained with the SentencePiece implementation (Kudo and Richardson, 2018).
+Training text is randomly sampled from the dataset for the desired language.
+[^3] To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling.
+After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens.
+We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form.
+Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling.
+[^4] When de-tokenized, the tokenized datasets are slightly smaller than the original text datasets, because the tokenizer truncates lines to create 512-token sequences.
+Reported dataset sizes account for truncation.
+We train each model for 10 epochs of the training data; multiple epochs of pretraining is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pretraining on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios.
+All language model pretraining runs together take a total of $1.65\times 10^{20}$ FLOPs.
+This is less than $1/1900\times$ the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020a; $3.14\times 10^{23}$ FLOPs).
+Further pretraining details are reported in §A.3.
+$3.$  Training Bigram Baselines
+Because there do not exist other models for many of these languages, we also train bigram baseline models on the same training data and with the same tokenizer for each language.
+Each bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset.
+When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e.
+“stupid backoff”; Brants et al., 2007).
+We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.
+4.  Evaluations
+4.1.  FLORES Log-Perplexity
 We first evaluate our models on FLORES-200 log-perplexity (Costa-jussà et al., 2022) (equivalently, negative log-likelihood; Lin et al., 2024).
-To avoid tokenization confounds from computing log-perplexity per token, we compute log-perplexity per FLORES sequence.
-Regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES.
+To avoid confounds from different tokenizers across models, we compute log-perplexities at the sequence level.
+Specifically, regardless of its tokenization, a language model $\mathcal{M}$ assigns some probability $P_{\mathcal{M}}(s)$ to each sequence $s$ in FLORES.
 In most cases, $s$ is a single sentence.
@@
 While imperfect, perplexity does not require annotated text data, it is predictive of performance on a variety of downstream tasks (Xia et al., 2023), and it has been used to measure language model quality in previous work (Kaplan et al., 2020; Hoffmann et al., 2022; Lin et al., 2024).
-We compare the Goldfish models with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages).
-We also compare to simple bigram models trained on the Goldfish datasets.
-[^4] Bigram and perplexity implementation details in §A.3.
-In all cases, we use the Goldfish model trained on the maximum amount of data in each language (maximum 1GB).
-FLORES log-perplexity results.
+We compare the Goldfish with XGLM 4.5B (Lin et al., 2022; 134 languages), XGLM 7.5B (30 languages), BLOOM 7.1B (Scao et al., 2022; 46 languages), and MaLA-500 10B (Lin et al., 2024; 534 languages).
+We also compare to our bigram baselines, which were trained on the same datasets as the Goldfish models.
+In all cases, we use the Goldfish model trained on the maximum amount of data in a language (maximum 1GB).
 The Goldfish reach lower log-perplexities than all four comparison models on 98 of the 204 FLORES languages.
-Average log-perplexities for the 100 FLORES languages included in both XGLM 4.5B and MaLA-500 are reported in Table 1 (excluding XGLM 7.5B and BLOOM 7.1B there because they are trained on far fewer languages).
-On average, the Goldfish reach 13% lower log-perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B.
-To ensure that these results are not driven by a small subset of specific languages, in Table 2 we report the pairwise “win” rates for Goldfish and bigrams vs.
+On average, the Goldfish reach 13% lower perplexities than XGLM 4.5B, and 11% lower than MaLA-500 10B (Table 1).
+To ensure that these results are not driven by a small subset of specific languages, in Table 2 we also report the pairwise “win” rates for Goldfish and bigrams vs.
 all four comparison models, for the set of FLORES languages shared between each pair.
@@
 Still, the bigrams have worse perplexities than Goldfish for all languages.
-Log-perplexities for individual languages and models are reported in Table LABEL:tab:ppl-table.
-5 Multilingual Reasoning Benchmarks
-Because FLORES perplexities are not necessarily reflective of complex capabilities in language models, we also evaluate Goldfish, XGLM, BLOOM, and MaLA-500 (as in §4) on non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense; Lin et al., 2022).
-All models are evaluated zero shot with no fine-tuning.
-Evaluation task details are in §A.4.
-Results for all three reasoning tasks are reported in Table 3.
-Although all models perform quite poorly (close to chance accuracy), the Goldfish perform substantially worse than the multilingual models.
-[^5] It is unlikely that this effect is due to model size alone; the Goldfish models (125M parameters) have easily enough capacity for their maximum of 1GB of text data.
-This indicates that the combination of larger datasets and model sizes in multilingual pre-training can allow language models to develop reasoning capabilities in specific languages, even when perplexities in those languages remain high.
-For example, XGLM 7.5B has worse perplexities than Goldfish for 82 Belebele languages (in fact, worse than bigrams for 77 languages), but it outperforms Goldfish on Belebele (reading comprehension) for 56 of those languages.
-This is in stark contrast with monolingual language models, which generally must reach low perplexities and acquire basic grammatical capabilities before developing reasoning abilities (Liu et al., 2021; Choshen et al., 2022; Xia et al., 2023; Chang et al., 2024).
-Intuitively, it may be that abstract reasoning patterns are often more language-agnostic than grammatical text generation, and thus multilingual pre-training primarily benefits the former.
-6 Conclusion
-We pre-train and release Goldfish, a suite of over 1000 monolingual language models for 350 languages.
+Perplexities for individual languages and models are available in our GitHub.
+Dataset / Model Chance Goldfish Gemma 3 BLOOM XGLM Gemma 3 LLaMA 3 XGLM
+124M 270M 560M 564M 1B 1B 1.7B
+MultiBLiMP (avg) 50.0 78.8 72.3 64.2 66.6 78.0 77.4 62.7
+Belebele (avg) 25.0 28.2 22.7 28.9 29.0 22.7 28.9 28.2
+XCOPA (avg) 50.0 55.1 54.6 54.4 55.2 59.7 55.9 57.6
+XStoryCloze (avg) 50.0 52.3 52.7 52.6 53.1 58.5 55.1 56.2
+Table 3: Average performance on downstream benchmarks for all models tested.
+4.2.  Downstream Tasks
+Next, we evaluate linguistic knowledge (grammaticality) with MultiBLiMP (Jumelet et al., 2025) as implemented in the LM Evaluation Harness (Gao et al., 2023); this covers 74 of the Goldfish languages.
+We compare against popular small multilingual models: BLOOM 560M (Scao et al., 2022), XGLM (564M and 1.7B; Lin et al., 2021), Gemma 3 (270M and 1B base models; Gemma Team, 2025), and Llama 3.2 (1B base model; Meta AI, 2024).
+Results are in Table 3.
+The Goldfish models have higher average accuracy than any of the multilingual models, and they have the highest accuracy of all models tested for 25 of the 74 languages.
+This result highlights the benefits of small monolingual models, especially for languages which only account for a small portion of the training data in multilingual models.
+We also evaluate the same models on three popular multilingual reasoning benchmarks: Belebele (121 languages, reading comprehension; Bandarkar et al., 2024), XCOPA (11 languages, commonsense reasoning; Ponti et al., 2020), and XStoryCloze (10 languages, story commonsense reasoning; Lin et al., 2022).
+All models are evaluated zero shot with log-probability solution ranking, with no fine-tuning or instruction tuning.
+Unfortunately, all models perform quite poorly (close to chance accuracy; Table 3), suggesting that available multilingual reasoning evaluations are not appropriate for pretrained-only models of this scale.
+5.  Discussion
+Our results demonstrate that large multilingual language models still struggle with basic grammatical text generation for many languages (e.g.
+often worse perplexities than bigrams).
+The Goldfish (small monolingual models) exhibit lower perplexities and more grammatical next token predictions for many low-resource languages.
+These results make the Goldfish suitable baselines for basic grammatical text generation in diverse languages, motivating future work developing low-resource language models.
+Furthermore, the Goldfish are accessible to labs with limited compute budgets, they are trained on a roughly human-like amount of data (Warstadt et al., 2023), and they are trained to be maximally comparable across languages.
+Future work may investigate precisely when larger-scale multilingual pretraining provides benefits to lower-resource languages; for example, it may be that abstract reasoning patterns and heuristics are often more language-agnostic than grammatical text generation, and thus larger-scale multilingual pretraining primarily benefits the former.
+6.  Conclusion
+In this paper, we pretrain and release Goldfish, a suite of over 1000 monolingual language models for 350 languages.
+For the majority of these languages, the Goldfish represent the first monolingual model dedicated for that language.
 The Goldfish achieve perplexities that are competitive with, and on average lower than, state-of-the-art multilingual language models across languages.
-However, they underperform large multilingual models on reasoning tasks; in low-resource languages, it appears that multilingual pre-training facilitates nontrivial reasoning capabilities despite extremely poor perplexities.
-We publicly release all Goldfish models to be used as comparable baselines, fine-tuning sources, or augmentations to larger models (e.g.
-cross-lingual experts; Blevins et al., 2024) in future low-resource NLP research.
+However, similar to multilingual models of the same scale, the Goldfish still struggle with reasoning tasks.
+This suggests that smaller monolingual models may better represent linguistic knowledge of the target language, but they do not perform well on more complex tasks at this scale.
+We publicly release the Goldfish models to promote future research pushing the limits of language model capabilities in low-resource languages.
 Limitations
@@
 By design, all of the Goldfish models are monolingual.
-For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020; Chang et al., 2023).
+For low-resource languages, training on closely related languages would likely improve performance (Conneau et al., 2020b; Chang et al., 2024a).
 However, adding multilingual data introduces concerns such as the choice of added languages (some languages have more closely related languages in our dataset than others), quantities of added data, and model capacity limitations.
@@
 For the majority of the world’s languages, 1GB is sufficient to include all publicly available text data in the language.
-At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2023).
+At these small dataset sizes, larger models do not appear to provide significant benefit over smaller models (Kaplan et al., 2020; Hoffmann et al., 2022; Chang et al., 2024a).
 Thus, the largest Goldfish model that we train for each language has 125M parameters and is trained on a maximum of 1GB of text.
@@
 Downstream tasks.
-We evaluate the Goldfish models on FLORES log-perplexity (§4) and three reasoning benchmarks (§5).
-These are some of the only evaluations that can be used for autoregressive language models in many languages, but they have significant limitations.
-Perplexity is not necessarily predictive of grammatical text generation (Hu et al., 2020) or complex reasoning capabilities (Levy et al., 2024), but it still provides reasonable signal for model performance (Xia et al., 2023) and it is often used to roughly quantify language model quality (Kaplan et al., 2020; Hoffmann et al., 2022).
+We evaluate the Goldfish models on FLORES log-perplexity (§4.1) and four multilingual benchmarks, including three reasoning benchmarks (§4.2).
+Perplexity is the only evaluation available for autoregressive language models in many languages before instruction-tuning and RLHF, but it has significant limitations.
+Perplexity is not necessarily predictive of more specific capabilities (Hu et al., 2020a; Levy et al., 2024), although it still provides reasonable signal for model performance (Xia et al., 2023).
 On the other hand, reasoning benchmarks require annotated datasets and thus often cover fewer languages.
-One notable exception is Belebele (121 non-English languages; Bandarkar et al., 2024), but even large state-of-the-art models perform quite poorly on Belebele without tuning or few-shot prompting (§5).
-Thus, our evaluations of model reasoning are not entirely conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”).
-We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future.
+Belebele (121 non-English languages; Bandarkar et al., 2024) is an exception, but even state-of-the-art models perform quite poorly on Belebele without instruction-tuning or few-shot prompting (§4.2).
+Thus, our evaluations of model reasoning are not conclusive; we may primarily be measuring heuristics that allow the models to perform only somewhat above chance (arguably, this might still be considered a basic form of “reasoning”).
+Outside of reasoning benchmarks, we also evaluate on MultiBLiMP (Jumelet et al., 2025), which evaluates linguistic knowledge.
+MultiBLiMP significantly expands the language coverage of multilingual grammaticality benchmarks; however, MultiBLiMP evaluates a very narrow aspect of grammatical knowledge (two types of subject-verb agreement).
+We hope that tractable evaluation datasets with broad language coverage will become increasingly available in the future to enable more informative evaluation of models in a broader range of languages.
 Risks and dataset licensing.
@@
 The Goldfish are trained on publicly-released corpora used in previous NLP research (§A.1), but we cannot guarantee that the data is free from offensive content or personally identifying information.
-We do not redistribute the data itself.
-Furthermore, our models are small, which reduces the likelihood that they will regurgitate memorized text (Carlini et al., 2023).
+Our models are small, which reduces the likelihood that they will regurgitate memorized text (Carlini et al., 2023).
 As far as we are aware, we do not include any datasets that prohibit use for language model training.
@@
 We will remove models for affected languages if contacted by dataset owners.
-Acknowledgments
-We would like to thank the UCSD Language and Cognition Lab for valuable discussion.
-Some models were trained on hardware provided by the NVIDIA Corporation as part of an NVIDIA Academic Hardware Grant.
-Some models were also trained on the UCSD Social Sciences Research and Development Environment (SSRDE).
-Zhuowen Tu is supported by NSF IIS-2127544.
-Tyler Chang is partially supported by the UCSD HDSI graduate fellowship.
+7.  Bibliographical References
+[…]
+  • Abadji et al.
+(2021) Julien Abadji, Pedro Javier Ortiz Suárez, Laurent Romary, and Benoît Sagot.
+2021.
+Ungoliant: An optimized pipeline for the generation of a very large-scale multilingual web corpus.
+In Proceedings of the Workshop on Challenges in the Management of Large Corpora (CMLC-9) 2021.
+Limerick, 12 July 2021 (Online-Event), pages 1 – 9.
+  • Abdelali et al.
+(2021) Ahmed Abdelali, Hamdy Mubarak, Younes Samih, Sabit Hassan, and Kareem Darwish.
+2021.
+QADI: Arabic dialect identification in the wild.
+In Proceedings of the Sixth Arabic Natural Language Processing Workshop, pages 1–10, Kyiv, Ukraine (Virtual).
+Association for Computational Linguistics.
+  • Abu Kwaik et al.
+(2018) Kathrein Abu Kwaik, Motaz Saad, Stergios Chatzikyriakidis, and Simon Dobnik.
+2018.
+Shami: A corpus of Levantine Arabic dialects.
+In Proceedings of the Eleventh International Conference on Language Resources and Evaluation (LREC 2018), Miyazaki, Japan.
+European Language Resources Association (ELRA).
+  • Adebara et al.
+(2022) Ife Adebara, AbdelRahim Elmadany, Muhammad Abdul-Mageed, and Alcides Alcoba Inciarte.
+2022.
+SERENGETI: Massively multilingual language models for Africa.
+arXiv preprint arXiv:2212.10785.
+  • Adelani et al.
+(2022) David Adelani, Jesujoba Alabi, Angela Fan, Julia Kreutzer, Xiaoyu Shen, Machel Reid, Dana Ruiter, Dietrich Klakow, Peter Nabende, Ernie Chang, Tajuddeen Gwadabe, Freshia Sackey, Bonaventure F.
+P.
+Dossou, Chris Emezue, Colin Leong, Michael Beukman, Shamsuddeen Muhammad, Guyo Jarso, Oreen Yousuf, Andre Niyongabo Rubungo, Gilles Hacheme, Eric Peter Wairagala, Muhammad Umair Nasir, Benjamin Ajibade, Tunde Ajayi, Yvonne Gitau, Jade Abbott, Mohamed Ahmed, Millicent Ochieng, Anuoluwapo Aremu, Perez Ogayo, Jonathan Mukiibi, Fatoumata Ouoba Kabore, Godson Kalipe, Derguene Mbaye, Allahsera Auguste Tapo, Victoire Memdjokam Koagne, Edwin Munkoh-Buabeng, Valencia Wagner, Idris Abdulmumin, Ayodele Awokoya, Happy Buzaaba, Blessing Sibanda, Andiswa Bukula, and Sam Manthalu.
+2022.
+A few thousand translations go a long way!
+leveraging pre-trained models for African news translation.
+In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 3053–3070, Seattle, United States.
+Association for Computational Linguistics.
+  • Adelani et al.
+(2021a) David Adelani, Dana Ruiter, Jesujoba Alabi, Damilola Adebonojo, Adesina Ayeni, Mofe Adeyemi, Ayodele Esther Awokoya, and Cristina España-Bonet.
+2021a.
+The effect of domain and diacritics in Yoruba–English neural machine translation.
+In Proceedings of Machine Translation Summit XVIII: Research Track, pages 61–75, Virtual.
+Association for Machine Translation in the Americas.
+  • Adelani et al.
+(2021b) David Ifeoluwa Adelani, Jade Abbott, Graham Neubig, Daniel D’souza, Julia Kreutzer, Constantine Lignos, Chester Palen-Michel, Happy Buzaaba, Shruti Rijhwani, Sebastian Ruder, Stephen Mayhew, Israel Abebe Azime, Shamsuddeen H.
+Muhammad, Chris Chinenye Emezue, Joyce Nakatumba-Nabende, Perez Ogayo, Aremu Anuoluwapo, Catherine Gitau, Derguene Mbaye, Jesujoba Alabi, Seid Muhie Yimam, Tajuddeen Rabiu Gwadabe, Ignatius Ezeani, Rubungo Andre Niyongabo, Jonathan Mukiibi, Verrah Otiende, Iroro Orife, Davis David, Samba Ngom, Tosin Adewumi, Paul Rayson, Mofetoluwa Adeyemi, Gerald Muriuki, Emmanuel Anebi, Chiamaka Chukwuneke, Nkiruka Odu, Eric Peter Wairagala, Samuel Oyerinde, Clemencia Siro, Tobius Saul Bateesa, Temilola Oloyede, Yvonne Wambui, Victor Akinode, Deborah Nabagereka, Maurice Katusiime, Ayodele Awokoya, Mouhamadane MBOUP, Dibora Gebreyohannes, Henok Tilaye, Kelechi Nwaike, Degaga Wolde, Abdoulaye Faye, Blessing Sibanda, Orevaoghene Ahia, Bonaventure F.
+P.
+Dossou, Kelechi Ogueji, Thierno Ibrahima DIOP, Abdoulaye Diallo, Adewale Akinfaderin, Tendai Marengereke, and Salomey Osei.
+2021b.
+MasakhaNER: Named entity recognition for African languages.
+Transactions of the Association for Computational Linguistics, 9:1116–1131.
+  • Agerri et al.
+(2018) Rodrigo Agerri, Xavier Gómez Guinovart, German Rigau, and Miguel Anxo Solla Portela.
+2018.
+Developing new linguistic resources and tools for the Galician language.
+In Proceedings of the Eleventh International Conference on Language Resources and Evaluation (LREC 2018), Miyazaki, Japan.
+European Language Resources Association (ELRA).
+  • Aho and Ullman (1972) Alfred V. Aho and Jeffrey D. Ullman. 1972. The Theory of Parsing, Translation and Compiling, volume 1. Prentice-Hall, Englewood Cliffs, NJ.
+  • Ahuja et al.
+(2022) Kabir Ahuja, Sunayana Sitaram, Sandipan Dandapat, and Monojit Choudhury.
+2022.
+On the calibration of massively multilingual language models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 4310–4323.
+Association for Computational Linguistics.
+  • AI FOR THAI (2023) AI FOR THAI. 2023. Ai for thai lotuscorpus. Dataset.
+  • AI4Bharat (2023) AI4Bharat. 2023. AI4Bharat. Dataset.
+  • Alabi et al.
+(2022) Jesujoba O.
+Alabi, David Ifeoluwa Adelani, Marius Mosbach, and Dietrich Klakow.
+2022.
+Adapting pre-trained language models to African languages via multilingual adaptive fine-tuning.
+In Proceedings of the 29th International Conference on Computational Linguistics, pages 4336–4349.
+International Committee on Computational Linguistics.
+  • Alsarsour et al.
+(2018) Israa Alsarsour, Esraa Mohamed, Reem Suwaileh, and Tamer Elsayed.
+2018.
+DART: A large dataset of dialectal Arabic tweets.
+In Proceedings of the Eleventh International Conference on Language Resources and Evaluation (LREC 2018), Miyazaki, Japan.
+European Language Resources Association (ELRA).
+  • American Psychological Association (1983) American Psychological Association. 1983. Publications Manual. American Psychological Association, Washington, DC.
+  • Anastasopoulos et al.
+(2020) Antonios Anastasopoulos, Alessandro Cattelan, Zi-Yi Dou, Marcello Federico, Christian Federmann, Dmitriy Genzel, Franscisco Guzmán, Junjie Hu, Macduff Hughes, Philipp Koehn, Rosie Lazar, Will Lewis, Graham Neubig, Mengmeng Niu, Alp Öktem, Eric Paquin, Grace Tang, and Sylwia Tur.
+2020.
+TICO-19: the translation initiative for COvid-19.
+In Proceedings of the 1st Workshop on NLP for COVID-19 (Part 2) at EMNLP 2020, Online.
+Association for Computational Linguistics.
+  • Ando and Zhang (2005) Rie Kubota Ando and Tong Zhang.
+2005.
+A framework for learning predictive structures from multiple tasks and unlabeled data.
+Journal of Machine Learning Research, 6:1817–1853.
+  • Andrew and Gao (2007) Galen Andrew and Jianfeng Gao.
+2007.
+Scalable training of $L_{1}$-regularized log-linear models.
+In Proceedings of the 24th International Conference on Machine Learning, pages 33–40.
+  • Ansell et al.
+(2022) Alan Ansell, Edoardo Ponti, Anna Korhonen, and Ivan Vulić.
+2022.
+Composable sparse fine-tuning for cross-lingual transfer.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1778–1796.
+  • Anuvaad (2023) Anuvaad. 2023. Anuvaad project. Dataset.
+  • Armengol-Estapé et al.
+(2022) Jordi Armengol-Estapé, Ona de Gibert Bonet, and Maite Melero.
+2022.
+On the multilingual capabilities of very large-scale english language models.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 3056–3068.
+  • Arnett and Chang (2025) Catherine Arnett and Tyler Chang. 2025. An Analysis of Multilingual Models on Hugging Face. Hugging Face blog post.
+  • Arnett et al.
+(2024) Catherine Arnett, Tyler A.
+Chang, and Benjamin Bergen.
+2024.
+A bit of a problem: Measurement disparities in dataset sizes across languages.
+In Proceedings of the 3rd Annual Meeting of the Special Interest Group on Under-resourced Languages @ LREC-COLING 2024, pages 1–9, Torino, Italia.
+ELRA and ICCL.
+  • Artetxe et al.
+(2020) Mikel Artetxe, Sebastian Ruder, and Dani Yogatama.
+2020.
+On the cross-lingual transferability of monolingual representations.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4623–4637.
+  • Autshumato (2023) Autshumato. 2023. Autshumato. Dataset.
+  • Bafna (2022) Niyati Bafna. 2022. Empirical models for an indic language continuum.
+  • Bandarkar et al.
+(2024) Lucas Bandarkar, Davis Liang, Benjamin Muller, Mikel Artetxe, Satya Narayan Shukla, Donald Husa, Naman Goyal, Abhinandan Krishnan, Luke Zettlemoyer, and Madian Khabsa.
+2024.
+The Belebele benchmark: A parallel reading comprehension dataset in 122 language variants.
+In Annual Meeting of the Association for Computational Linguistics.
+  • Bañón et al.
+(2020) Marta Bañón, Pinzhen Chen, Barry Haddow, Kenneth Heafield, Hieu Hoang, Miquel Esplà-Gomis, Mikel L.
+Forcada, Amir Kamran, Faheem Kirefu, Philipp Koehn, Sergio Ortiz Rojas, Leopoldo Pla Sempere, Gema Ramírez-Sánchez, Elsa Sarrías, Marek Strelec, Brian Thompson, William Waites, Dion Wiggins, and Jaume Zaragoza.
+2020.
+ParaCrawl: Web-scale acquisition of parallel corpora.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 4555–4567, Online.
+Association for Computational Linguistics.
+  • Bañón et al.
+(2022) Marta Bañón, Miquel Esplà-Gomis, Mikel L.
+Forcada, Cristian García-Romero, Taja Kuzman, Nikola Ljubesic, Rik van Noord, Leopoldo Pla Sempere, Gema Ramírez-Sánchez, Peter Rupnik, Vít Suchomel, Antonio Toral, Tobias van der Werff, and Jaume Zaragoza.
+2022.
+Macocu: Massive collection and curation of monolingual and bilingual data: focus on under-resourced languages.
+In Proceedings of the 23rd Annual Conference of the European Association for Machine Translation, EAMT 2022, Ghent, Belgium, June 1-3, 2022, pages 301–302.
+European Association for Machine Translation.
+  • Bapna et al.
+(2022) Ankur Bapna, Isaac Caswell, Julia Kreutzer, Orhan Firat, Daan van Esch, Aditya Siddhant, Mengmeng Niu, Pallavi Baljekar, Xavier Garcia, Wolfgang Macherey, et al.
+2022.
+Building machine translation systems for the next thousand languages.
+arXiv preprint arXiv:2205.03983.
+  • Bender (2019) Emily Bender. 2019. The# benderrule: On naming the languages we study and why it matters. The Gradient, 14(1).
+  • Bender (2009) Emily M Bender.
+2009.
+Linguistically naïve != language independent: Why NLP needs linguistic typology.
+In Proceedings of the EACL 2009 Workshop on the Interaction between Linguistics and Computational Linguistics: Virtuous, Vicious or Vacuous?, pages 26–32.
+  • Bender (2011) Emily M Bender. 2011. On achieving and evaluating language-independence in NLP. Linguistic Issues in Language Technology, 6.
+  • BigScience et al.
+(2022) Workshop BigScience, Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, Jonathan Tow, Alexander M.
+Rush, Stella Biderman, Albert Webson, Pawan Sasanka Ammanamanchi, Thomas Wang, Benoît Sagot, Niklas Muennighoff, Albert Villanova del Moral, Olatunji Ruwase, Rachel Bawden, Stas Bekman, Angelina McMillan-Major, Iz Beltagy, Huu Nguyen, Lucile Saulnier, Samson Tan, Pedro Ortiz Suarez, Victor Sanh, Hugo Laurençon, Yacine Jernite, Julien Launay, Margaret Mitchell, Colin Raffel, Aaron Gokaslan, Adi Simhi, Aitor Soroa, Alham Fikri Aji, Amit Alfassy, Anna Rogers, Ariel Kreisberg Nitzav, Canwen Xu, Chenghao Mou, Chris Emezue, Christopher Klamm, Colin Leong, Daniel van Strien, David Ifeoluwa Adelani, Dragomir Radev, Eduardo González Ponferrada, Efrat Levkovizh, Ethan Kim, Eyal Bar Natan, Francesco De Toni, Gérard Dupont, Germán Kruszewski, Giada Pistilli, Hady Elsahar, Hamza Benyamina, Hieu Tran, Ian Yu, Idris Abdulmumin, Isaac Johnson, Itziar Gonzalez-Dios, Javier de la Rosa, Jenny Chim, Jesse Dodge, Jian Zhu, Jonathan Chang, Jörg Frohberg, Joseph Tobing, Joydeep Bhattacharjee, Khalid Almubarak, Kimbo Chen, Kyle Lo, Leandro Von Werra, Leon Weber, Long Phan, Loubna Ben allal, Ludovic Tanguy, Manan Dey, Manuel Romero Muñoz, Maraim Masoud, María Grandury, Mario Šaško, Max Huang, Maximin Coavoux, Mayank Singh, Mike Tian-Jian Jiang, Minh Chien Vu, Mohammad A.
+Jauhar, Mustafa Ghaleb, Nishant Subramani, Nora Kassner, Nurulaqilla Khamis, Olivier Nguyen, Omar Espejel, Ona de Gibert, Paulo Villegas, Peter Henderson, Pierre Colombo, Priscilla Amuok, Quentin Lhoest, Rheza Harliman, Rishi Bommasani, Roberto Luis López, Rui Ribeiro, Salomey Osei, Sampo Pyysalo, Sebastian Nagel, Shamik Bose, Shamsuddeen Hassan Muhammad, Shanya Sharma, Shayne Longpre, Somaieh Nikpoor, Stanislav Silberberg, Suhas Pai, Sydney Zink, Tiago Timponi Torrent, Timo Schick, Tristan Thrush, Valentin Danchev, Vassilina Nikoulina, Veronika Laippala, Violette Lepercq, Vrinda Prabhu, Zaid Alyafeai, Zeerak Talat, Arun Raja, Benjamin Heinzerling, Chenglei Si, Davut Emre Taşar, Elizabeth Salesky, Sabrina J.
+Mielke, Wilson Y.
+Lee, Abheesht Sharma, Andrea Santilli, Antoine Chaffin, Arnaud Stiegler, Debajyoti Datta, Eliza Szczechla, Gunjan Chhablani, Han Wang, Harshit Pandey, Hendrik Strobelt, Jason Alan Fries, Jos Rozen, Leo Gao, Lintang Sutawika, M Saiful Bari, Maged S.
+Al-shaibani, Matteo Manica, Nihal Nayak, Ryan Teehan, Samuel Albanie, Sheng Shen, Srulik Ben-David, Stephen H.
+Bach, Taewoon Kim, Tali Bers, Thibault Fevry, Trishala Neeraj, Urmish Thakker, Vikas Raunak, Xiangru Tang, Zheng-Xin Yong, Zhiqing Sun, Shaked Brody, Yallow Uri, Hadar Tojarieh, Adam Roberts, Hyung Won Chung, Jaesung Tae, Jason Phang, Ofir Press, Conglong Li, Deepak Narayanan, Hatim Bourfoune, Jared Casper, Jeff Rasley, Max Ryabinin, Mayank Mishra, Minjia Zhang, Mohammad Shoeybi, Myriam Peyrounette, Nicolas Patry, Nouamane Tazi, Omar Sanseviero, Patrick von Platen, Pierre Cornette, Pierre François Lavallée, Rémi Lacroix, Samyam Rajbhandari, Sanchit Gandhi, Shaden Smith, Stéphane Requena, Suraj Patil, Tim Dettmers, Ahmed Baruwa, Amanpreet Singh, Anastasia Cheveleva, Anne-Laure Ligozat, Arjun Subramonian, Aurélie Névéol, Charles Lovering, Dan Garrette, Deepak Tunuguntla, Ehud Reiter, Ekaterina Taktasheva, Ekaterina Voloshina, Eli Bogdanov, Genta Indra Winata, Hailey Schoelkopf, Jan-Christoph Kalo, Jekaterina Novikova, Jessica Zosa Forde, Jordan Clive, Jungo Kasai, Ken Kawamura, Liam Hazan, Marine Carpuat, Miruna Clinciu, Najoung Kim, Newton Cheng, Oleg Serikov, Omer Antverg, Oskar van der Wal, Rui Zhang, Ruochen Zhang, Sebastian Gehrmann, Shachar Mirkin, Shani Pais, Tatiana Shavrina, Thomas Scialom, Tian Yun, Tomasz Limisiewicz, Verena Rieser, Vitaly Protasov, Vladislav Mikhailov, Yada Pruksachatkun, Yonatan Belinkov, Zachary Bamberger, Zdeněk Kasner, Alice Rueda, Amanda Pestana, Amir Feizpour, Ammar Khan, Amy Faranak, Ana Santos, Anthony Hevia, Antigona Unldreaj, Arash Aghagol, Arezoo Abdollahi, Aycha Tammour, Azadeh HajiHosseini, Bahareh Behroozi, Benjamin Ajibade, Bharat Saxena, Carlos Muñoz Ferrandis, Danish Contractor, David Lansky, Davis David, Douwe Kiela, Duong A.
+Nguyen, Edward Tan, Emi Baylor, Ezinwanne Ozoani, Fatima Mirza, Frankline Ononiwu, Habib Rezanejad, Hessie Jones, Indrani Bhattacharya, Irene Solaiman, Irina Sedenko, Isar Nejadgholi, Jesse Passmore, Josh Seltzer, Julio Bonis Sanz, Livia Dutra, Mairon Samagaio, Maraim Elbadri, Margot Mieskes, Marissa Gerchick, Martha Akinlolu, Michael McKenna, Mike Qiu, Muhammed Ghauri, Mykola Burynok, Nafis Abrar, Nazneen Rajani, Nour Elkott, Nour Fahmy, Olanrewaju Samuel, Ran An, Rasmus Kromann, Ryan Hao, Samira Alizadeh, Sarmad Shubber, Silas Wang, Sourav Roy, Sylvain Viguier, Thanh Le, Tobi Oyebade, Trieu Le, Yoyo Yang, Zach Nguyen, Abhinav Ramesh Kashyap, Alfredo Palasciano, Alison Callahan, Anima Shukla, Antonio Miranda-Escalada, Ayush Singh, Benjamin Beilharz, Bo Wang, Caio Brito, Chenxi Zhou, Chirag Jain, Chuxin Xu, Clémentine Fourrier, Daniel León Periñán, Daniel Molano, Dian Yu, Enrique Manjavacas, Fabio Barth, Florian Fuhrimann, Gabriel Altay, Giyaseddin Bayrak, Gully Burns, Helena U.
+Vrabec, Imane Bello, Ishani Dash, Jihyun Kang, John Giorgi, Jonas Golde, Jose David Posada, Karthik Rangasai Sivaraman, Lokesh Bulchandani, Lu Liu, Luisa Shinzato, Madeleine Hahn de Bykhovetz, Maiko Takeuchi, Marc Pàmies, Maria A Castillo, Marianna Nezhurina, Mario Sänger, Matthias Samwald, Michael Cullan, Michael Weinberg, Michiel De Wolf, Mina Mihaljcic, Minna Liu, Moritz Freidank, Myungsun Kang, Natasha Seelam, Nathan Dahlberg, Nicholas Michio Broad, Nikolaus Muellner, Pascale Fung, Patrick Haller, Ramya Chandrasekhar, Renata Eisenberg, Robert Martin, Rodrigo Canalli, Rosaline Su, Ruisi Su, Samuel Cahyawijaya, Samuele Garda, Shlok S Deshmukh, Shubhanshu Mishra, Sid Kiblawi, Simon Ott, Sinee Sang-aroonsiri, Srishti Kumar, Stefan Schweter, Sushil Bharati, Tanmay Laud, Théo Gigant, Tomoya Kainuma, Wojciech Kusa, Yanis Labrak, Yash Shailesh Bajaj, Yash Venkatraman, Yifan Xu, Yingxin Xu, Yu Xu, Zhe Tan, Zhongli Xie, Zifan Ye, Mathilde Bras, Younes Belkada, and Thomas Wolf.
+2022.
+BLOOM: a 176b-parameter open-access multilingual language model.
+  • Blaise Cruz and Cheng (2020) Jan Christian Blaise Cruz and Charibeth Cheng. 2020. Establishing baselines for text classification in low-resource languages. arXiv e-prints, pages arXiv–2005.
+  • Blasi et al.
+(2022) Damian Blasi, Antonios Anastasopoulos, and Graham Neubig.
+2022.
+Systematic inequalities in language technology performance across the world’s languages.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 5486–5505.
+Association for Computational Linguistics.
+  • Blevins et al.
+(2022) Terra Blevins, Hila Gonen, and Luke Zettlemoyer.
+2022.
+Analyzing the mono- and cross-lingual pretraining dynamics of multilingual language models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 3575–3590.
+Association for Computational Linguistics.
+  • Blevins et al.
+(2024) Terra Blevins, Tomasz Limisiewicz, Suchin Gururangan, Margaret Li, Hila Gonen, Noah A.
+Smith, and Luke Zettlemoyer.
+2024.
+Breaking the curse of multilinguality with cross-lingual expert language models.
+arXiv.
+  • Blevins and Zettlemoyer (2022) Terra Blevins and Luke Zettlemoyer.
+2022.
+Language contamination helps explains the cross-lingual capabilities of English pretrained models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 3563–3574, Abu Dhabi, United Arab Emirates.
+Association for Computational Linguistics.
+  • Bonferroni (1936) C.E. Bonferroni. 1936. Teoria statistica delle classi ecalcolo delle probabilità. Pubblicazioni del R Instituto Superiore di Scienze Economiche e Commerciali di Firenze, 8:3–62.
+  • Borcard et al. (1992) Daniel Borcard, Pierre Legendre, and Pierre Drapeau. 1992. Partialling out the spatial component of ecological variation. Ecology, 73(3):1045–1055.
+  • Brants et al.
+(2007) Thorsten Brants, Ashok C.
+Popat, Peng Xu, Franz J.
+Och, and Jeffrey Dean.
+2007.
+Large language models in machine translation.
+In Proceedings of the 2007 Joint Conference on Empirical Methods in Natural Language Processing and Computational Natural Language Learning (EMNLP-CoNLL), pages 858–867, Prague, Czech Republic.
+Association for Computational Linguistics.
+  • Brown et al.
+(2020a) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel Ziegler, Jeffrey Wu, Clemens Winter, Chris Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei.
+2020a.
+Language models are few-shot learners.
+In Advances in Neural Information Processing Systems, volume 33, pages 1877–1901.
+  • Brown et al.
+(2020b) Tom Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared D Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, et al.
+2020b.
+Language models are few-shot learners.
+Advances in neural information processing systems, 33:1877–1901.
+  • Camacho-Collados et al.
+(2016) José Camacho-Collados, Claudio Delli Bovi, Alessandro Raganato, and Roberto Navigli.
+2016.
+A large-scale multilingual disambiguation of glosses.
+In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), pages 1701–1708, Portorož, Slovenia.
+European Language Resources Association (ELRA).
+  • Campos et al.
+(2020) José Ramom Pichel Campos, Pablo Gamallo Otero, and Iñaki Alegria Loinaz.
+2020.
+Measuring diachronic language distance using perplexity: Application to english, portuguese, and spanish.
+Natural Language Engineering, 26(4):433–454.
+  • Carey (2024) Lily Carey. 2024. Goldfish may have a longer memory span than just three seconds. Discover Magazine.
+  • Carlini et al.
+(2023) Nicholas Carlini, Daphne Ippolito, Matthew Jagielski, Katherine Lee, Florian Tramer, and Chiyuan Zhang.
+2023.
+Quantifying memorization across neural language models.
+In International Conference on Learning Representations.
+  • Caswell et al.
+(2020) Isaac Caswell, Theresa Breiner, Daan van Esch, and Ankur Bapna.
+2020.
+Language ID in the wild: Unexpected challenges on the path to a thousand-language web text corpus.
+In Proceedings of the 28th International Conference on Computational Linguistics, pages 6588–6608, Barcelona, Spain (Online).
+International Committee on Computational Linguistics.
+  • Cawoylel (2023) Cawoylel. 2023. Fula speech corpus.
+  • Chai et al.
+(2022) Yuan Chai, Yaobo Liang, and Nan Duan.
+2022.
+Cross-lingual ability of multilingual masked language models: A study of language structure.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 4702–4712.
+Association for Computational Linguistics.
+  • Chandra et al. (1981) Ashok K. Chandra, Dexter C. Kozen, and Larry J. Stockmeyer. 1981. Alternation. Journal of the Association for Computing Machinery, 28(1):114–133.
+  • Chang et al.
+(2024a) Tyler A.
+Chang, Catherine Arnett, Zhuowen Tu, and Benjamin K.
+Bergen.
+2024a.
+When is multilinguality a curse?
+language modeling for 250 high- and low-resource languages.
+In Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing, pages 4074–4096, Miami, Florida, USA.
+Association for Computational Linguistics.
+  • Chang and Bergen (2022) Tyler A. Chang and Benjamin K. Bergen. 2022. Word acquisition in neural language models. Transactions of the Association for Computational Linguistics, 10:1–16.
+  • Chang et al.
+(2024b) Tyler A.
+Chang, Zhuowen Tu, and Benjamin K.
+Bergen.
+2024b.
+Characterizing learning curves during language model pre-training: Learning, forgetting, and stability.
+Transactions of the Association for Computational Linguistics.
+  • Cherokee Corpus (2023) Cherokee Corpus. 2023. Cherokee corpus and Cherokee-English Dictionary.
+  • Chi et al.
+(2020) Ethan A Chi, John Hewitt, and Christopher D Manning.
+2020.
+Finding universal grammatical relations in multilingual bert.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 5564–5577.
+  • Chi et al.
+(2021) Zewen Chi, Li Dong, Bo Zheng, Shaohan Huang, Xian-Ling Mao, Heyan Huang, and Furu Wei.
+2021.
+Improving pretrained cross-lingual language models via self-labeled word alignment.
+In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 3418–3430, Online.
+Association for Computational Linguistics.
+  • Choshen et al.
+(2022) Leshem Choshen, Guy Hacohen, Daphna Weinshall, and Omri Abend.
+2022.
+The grammar-learning trajectories of neural language models.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 8281–8297, Dublin, Ireland.
+Association for Computational Linguistics.
+  • Chowdhery et al.
+(2022) Aakanksha Chowdhery, Sharan Narang, Jacob Devlin, Maarten Bosma, Gaurav Mishra, Adam Roberts, Paul Barham, Hyung Won Chung, Charles Sutton, Sebastian Gehrmann, et al.
+2022.
+Palm: Scaling language modeling with pathways.
+arXiv preprint arXiv:2204.02311.
+  • Cieri et al.
+(2016) Christopher Cieri, Mike Maxwell, Stephanie Strassel, and Jennifer Tracey.
+2016.
+Selection criteria for low resource language programs.
+In Proceedings of the Tenth International Conference on Language Resources and Evaluation (LREC’16), pages 4543–4549, Portorož, Slovenia.
+European Language Resources Association (ELRA).
+  • Clarin (2023) Clarin. 2023. Clarin.si. Dataset.
+  • CMU (2010) CMU. 2010. Haitian Creole language data. http://www.speech.cs.cmu.edu/haitian/.
+  • Common Crawl (2022) Common Crawl. 2022. Common crawl. Dataset.
+  • Conneau et al.
+(2020a) Alexis Conneau, Alexei Baevski, Ronan Collobert, Abdelrahman Mohamed, and Michael Auli.
+2020a.
+Unsupervised cross-lingual representation learning for speech recognition.
+arXiv preprint arXiv:2006.13979.
+  • Conneau et al.
+(2020b) Alexis Conneau, Kartikay Khandelwal, Naman Goyal, Vishrav Chaudhary, Guillaume Wenzek, Francisco Guzmán, Edouard Grave, Myle Ott, Luke Zettlemoyer, and Veselin Stoyanov.
+2020b.
+Unsupervised cross-lingual representation learning at scale.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 8440–8451, Online.
+Association for Computational Linguistics.
+  • Conneau and Lample (2019) Alexis Conneau and Guillaume Lample. 2019. Cross-lingual language model pretraining. Advances in neural information processing systems, 32.
+  • Conneau et al.
+(2020c) Alexis Conneau, Shijie Wu, Haoran Li, Luke Zettlemoyer, and Veselin Stoyanov.
+2020c.
+Emerging cross-lingual structure in pretrained language models.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 6022–6034.
+Association for Computational Linguistics.
+  • Cooley and Tukey (1965) James W. Cooley and John W. Tukey. 1965. An algorithm for the machine calculation of complex Fourier series. Mathematics of Computation, 19(90):297–301.
+  • Costa-jussà et al.
+(2022) Marta R.
+Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al Youngblood, Bapi Akula, Loic Barrault, Gabriel Mejia Gonzalez, Prangthip Hansanti, John Hoffman, Semarley Jarrett, Kaushik Ram Sadagopan, Dirk Rowe, Shannon Spruit, Chau Tran, Pierre Andrews, Necip Fazil Ayan, Shruti Bhosale, Sergey Edunov, Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn, Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and Jeff Wang.
+2022.
+No language left behind: Scaling human-centered machine translation.
+arXiv.
+  • Dang et al.
+(2024) John Dang, Shivalika Singh, Daniel D’souza, Arash Ahmadian, Alejandro Salamanca, Madeline Smith, Aidan Peppin, Sungjin Hong, Manoj Govindassamy, Terrence Zhao, Sandra Kublik, Meor Amer, Viraat Aryabumi, Jon Ander Campos, Yi-Chern Tan, Tom Kocmi, Florian Strub, Nathan Grinsztajn, Yannis Flet-Berliac, Acyr Locatelli, Hangyu Lin, Dwarak Talupuru, Bharat Venkitesh, David Cairuz, Bowen Yang, Tim Chung, Wei-Yin Ko, Sylvie Shang Shi, Amir Shukayev, Sammie Bae, Aleksandra Piktus, Roman Castagné, Felipe Cruz-Salinas, Eddie Kim, Lucas Crawhall-Stein, Adrien Morisot, Sudip Roy, Phil Blunsom, Ivan Zhang, Aidan Gomez, Nick Frosst, Marzieh Fadaee, Beyza Ermis, Ahmet Üstün, and Sara Hooker.
+2024.
+Aya Expanse: Combining research breakthroughs for a new multilingual frontier.
+arXiv.
+  • de Marneffe et al. (2021) Marie-Catherine de Marneffe, Christopher D. Manning, Joakim Nivre, and Daniel Zeman. 2021. Universal dependencies. Computational Linguistics, 47(2):255–308.
+  • Devlin et al.
+(2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova.
+2019.
+BERT: Pre-training of deep bidirectional Transformers for language understanding.
+In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pages 4171–4186.
+Association for Computational Linguistics.
+  • Dryer and Haspelmath (2013) Matthew S. Dryer and Martin Haspelmath, editors. 2013. WALS Online (v2020.3). Zenodo.
+  • Dunn (2020) Jonathan Dunn. 2020. Mapping languages: the corpus of global language use. Lang. Resour. Evaluation, 54(4):999–1018.
+  • Eberhard et al. (2022) Eberhard, David M., Gary F. Simons, and Charles D. Fennig (eds.). 2022. Ethnologue: Languages of the world. twenty-fifth edition.
+  • eBible (2023) eBible. 2023. eBible. Dataset.
+  • Ebrahimi et al.
+(2022) Abteen Ebrahimi, Manuel Mager, Arturo Oncevay, Vishrav Chaudhary, Luis Chiruzzo, Angela Fan, John Ortega, Ricardo Ramos, Annette Rios, Ivan Vladimir Meza Ruiz, Gustavo Giménez-Lugo, Elisabeth Mager, Graham Neubig, Alexis Palmer, Rolando Coto-Solano, Thang Vu, and Katharina Kann.
+2022.
+AmericasNLI: Evaluating zero-shot natural language understanding of pretrained multilingual models in truly low-resource languages.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 6279–6299.
+Association for Computational Linguistics.
+  • El-Haj (2020) Mahmoud El-Haj.
+2020.
+Habibi - a multi dialect multi national Arabic song lyrics corpus.
+In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 1318–1326, Marseille, France.
+European Language Resources Association.
+  • El-Haj et al.
+(2018) Mahmoud El-Haj, Paul Rayson, and Mariam Aboelezz.
+2018.
+Arabic dialect identification in the context of bivalency and code-switching.
+In Proceedings of the Eleventh International Conference on Language Resources and Evaluation (LREC 2018), Miyazaki, Japan.
+European Language Resources Association (ELRA).
+  • ElJundi et al.
+(2019) Obeida ElJundi, Wissam Antoun, Nour El Droubi, Hazem Hajj, Wassim El-Hajj, and Khaled Shaban.
+2019.
+hULMonA: The universal language model in Arabic.
+In Proceedings of the Fourth Arabic Natural Language Processing Workshop, pages 68–77, Florence, Italy.
+Association for Computational Linguistics.
+  • Eronen et al.
+(2023) Juuso Eronen, Michal Ptaszynski, and Fumito Masui.
+2023.
+Zero-shot cross-lingual transfer language selection using linguistic similarity.
+Information Processing & Management, 60(3):103250.
+  • Ethnologue (2024) Ethnologue. 2024. Ethnologue, Languages of the World. SIL International.
+  • Fan et al.
+(2021) Angela Fan, Shruti Bhosale, Holger Schwenk, Zhiyi Ma, Ahmed El-Kishky, Siddharth Goyal, Mandeep Baines, Onur Celebi, Guillaume Wenzek, Vishrav Chaudhary, Naman Goyal, Tom Birch, Vitaliy Liptchinsky, Sergey Edunov, Michael Auli, and Armand Joulin.
+2021.
+Beyond english-centric multilingual machine translation.
+J.
+Mach.
+Learn.
+Res., 22:107:1–107:48.
+  • FFR Dataset (2023) FFR Dataset. 2023. Fon and french dataset. Dataset.
+  • FitzGerald et al.
+(2022) Jack FitzGerald, Christopher Hench, Charith Peris, Scott Mackie, Kay Rottmann, Ana Sanchez, Aaron Nash, Liam Urbach, Vishesh Kakarala, Richa Singh, et al.
+2022.
+Massive: A 1m-example multilingual natural language understanding dataset with 51 typologically-diverse languages.
+arXiv preprint arXiv:2204.08582.
+  • Foroutan et al.
+(2022) Negar Foroutan, Mohammadreza Banaei, Remi Lebret, Antoine Bosselut, and Karl Aberer.
+2022.
+Discovering language-neutral sub-networks in multilingual language models.
+arXiv preprint arXiv:2205.12672.
+  • Fujinuma et al.
+(2022) Yoshinari Fujinuma, Jordan Boyd-Graber, and Katharina Kann.
+2022.
+Match the script, adapt if multilingual: Analyzing the effect of multilingual pretraining on cross-lingual transferability.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1500–1512.
+Association for Computational Linguistics.
+  • Gaim et al. (2021) Fitsum Gaim, Wonsuk Yang, and Jong Park. 2021. Monolingual pre-trained language models for Tigrinya. Widening NLP Workshop (WiNLP).
+  • Gao et al.
+(2023) Leo Gao, Jonathan Tow, Baber Abbasi, Stella Biderman, Sid Black, Anthony DiPofi, Charles Foster, Laurence Golding, Jeffrey Hsu, Alain Le Noac’h, Haonan Li, Kyle McDonell, Niklas Muennighoff, Chris Ociepa, Jason Phang, Laria Reynolds, Hailey Schoelkopf, Aviya Skowron, Lintang Sutawika, Eric Tang, Anish Thite, Ben Wang, Kevin Wang, and Andy Zou.
+2023.
+Language model evaluation harness: A framework for few-shot language model evaluation.
+  • Gbedevi Akouyo et al.
+(2021) Yvette Gbedevi Akouyo, Kevin Zhang, and Tchaye-Kondi Jude.
+2021.
+GELR: A bilingual Ewe-English corpus building and evaluation.
+International Journal of Engineering Research and Technology (IJERT), 10.
+  • Gemma Team (2025) Google DeepMind Gemma Team. 2025. Gemma 3 technical report. arXiv.
+  • Gerz et al.
+(2018) Daniela Gerz, Ivan Vulić, Edoardo Maria Ponti, Roi Reichart, and Anna Korhonen.
+2018.
+On the relation between linguistic typology and (limitations of) multilingual language modeling.
+In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing, pages 316–327.
+Association for Computational Linguistics.
+  • Goldhahn et al.
+(2012) Dirk Goldhahn, Thomas Eckart, and Uwe Quasthoff.
+2012.
+Building large monolingual dictionaries at the Leipzig corpora collection: From 100 to 200 languages.
+In Proceedings of the Eighth International Conference on Language Resources and Evaluation (LREC’12), pages 759–765.
+European Language Resources Association (ELRA).
+  • Góngora et al.
+(2021) Santiago Góngora, Nicolás Giossa, and Luis Chiruzzo.
+2021.
+Experiments on a Guarani corpus of news and social media.
+In Proceedings of the First Workshop on Natural Language Processing for Indigenous Languages of the Americas, pages 153–158, Online.
+Association for Computational Linguistics.
+  • Góngora et al.
+(2022) Santiago Góngora, Nicolás Giossa, and Luis Chiruzzo.
+2022.
+Can we use word embeddings for enhancing Guarani-Spanish machine translation?
+In Proceedings of the Fifth Workshop on the Use of Computational Methods in the Study of Endangered Languages, pages 127–132, Dublin, Ireland.
+Association for Computational Linguistics.
+  • Google DeepMind (2023) Google DeepMind. 2023. Gemini: A family of highly capable multimodal models. Google.
+  • Gowda et al.
+(2021) Thamme Gowda, Zhao Zhang, Chris Mattmann, and Jonathan May.
+2021.
+Many-to-English machine translation tools, data, and pretrained models.
+In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing: System Demonstrations, pages 306–316, Online.
+Association for Computational Linguistics.
+  • Goyal et al.
+(2021) Naman Goyal, Jingfei Du, Myle Ott, Giri Anantharaman, and Alexis Conneau.
+2021.
+Larger-scale transformers for multilingual masked language modeling.
+In Proceedings of the 6th Workshop on Representation Learning for NLP (RepL4NLP-2021), pages 29–33.
+  • Guarasci et al.
+(2022) Raffaele Guarasci, Stefano Silvestri, Giuseppe De Pietro, Hamido Fujita, and Massimo Esposito.
+2022.
+BERT syntactic transfer: A computational experiment on Italian, French and English languages.
+Computer Speech & Language, 71:101261.
+  • Gusfield (1997) Dan Gusfield. 1997. Algorithms on Strings, Trees and Sequences. Cambridge University Press, Cambridge, UK.
+  • Gutiérrez-Fandiño et al.
+(2021) Asier Gutiérrez-Fandiño, Jordi Armengol-Estapé, Marc Pàmies, Joan Llop-Palao, Joaquin Silveira-Ocampo, Casimiro Pio Carrino, Aitor Gonzalez-Agirre, Carme Armentano-Oller, Carlos Rodriguez-Penagos, and Marta Villegas.
+2021.
+MarIA: Spanish language models.
+arXiv.
+  • Hammarström et al. (2023) Harald Hammarström, Robert Forkel, Martin Haspelmath, and Sebastian Bank. 2023. Glottolog 4.8. Max Planck Institute for Evolutionary Anthropology, Leipzig.
+  • Hämmerl et al.
+(2022) Katharina Hämmerl, Björn Deiseroth, Patrick Schramowski, Jindřich Libovickỳ, Constantin A Rothkopf, Alexander Fraser, and Kristian Kersting.
+2022.
+Speaking multiple languages affects the moral bias of language models.
+arXiv preprint arXiv:2211.07733.
+  • Hangya et al.
+(2022) Viktor Hangya, Hossain Shaikh Saadi, and Alexander Fraser.
+2022.
+Improving low-resource languages in pre-trained multilingual language models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 11993–12006.
+  • Hasan et al.
+(2021) Tahmid Hasan, Abhik Bhattacharjee, Md.
+Saiful Islam, Kazi Samin Mubasshir, Yuan-Fang Li, Yong-Bin Kang, M.
+Sohel Rahman, and Rifat Shahriyar.
+2021.
+XL-Sum: Large-scale multilingual abstractive summarization for 44 languages.
+In Findings of the Association for Computational Linguistics: ACL/IJCNLP 2021, Online Event, August 1-6, 2021, volume ACL/IJCNLP 2021 of Findings of ACL, pages 4693–4703.
+Association for Computational Linguistics.
+  • Hoffmann et al.
+(2022) Jordan Hoffmann, Sebastian Borgeaud, Arthur Mensch, Elena Buchatskaya, Trevor Cai, Eliza Rutherford, Diego de las Casas, Lisa Anne Hendricks, Johannes Welbl, Aidan Clark, Tom Hennigan, Eric Noland, Katherine Millican, George van den Driessche, Bogdan Damoc, Aurelia Guy, Simon Osindero, Karen Simonyan, Erich Elsen, Oriol Vinyals, Jack William Rae, and Laurent Sifre.
+2022.
+Training compute-optimal large language models.
+In Advances in Neural Information Processing Systems, volume 35, pages 30016–30030.
+  • Honnibal et al. (2020) Matthew Honnibal, Ines Montani, Sofie Van Landeghem, and Adriane Boyd. 2020. spaCy: Industrial-strength natural language processing in python. SpaCy.
+  • HornMT (2023) HornMT. 2023. Machine translation benchmark dataset for languages in the horn of africa. Dataset.
+  • Hu et al.
+(2020a) Jennifer Hu, Jon Gauthier, Peng Qian, Ethan Wilcox, and Roger Levy.
+2020a.
+A systematic assessment of syntactic generalization in neural language models.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 1725–1744, Online.
+Association for Computational Linguistics.
+  • Hu and Levy (2023) Jennifer Hu and Roger Levy. 2023. Prompt-based methods may underestimate large language models’ linguistic generalizations. arXiv preprint arXiv:2305.13264.
+  • Hu et al.
+(2020b) Junjie Hu, Sebastian Ruder, Aditya Siddhant, Graham Neubig, Orhan Firat, and Melvin Johnson.
+2020b.
+XTREME: A massively multilingual multi-task benchmark for evaluating cross-lingual generalisation.
+In International Conference on Machine Learning, pages 4411–4421.
+  • Hu et al.
+(2020c) Junjie Hu, Sebastian Ruder, Aditya Siddhant, Graham Neubig, Orhan Firat, and Melvin Johnson.
+2020c.
+XTREME: A massively multilingual multi-task benchmark for evaluating cross-lingual generalisation.
+In Proceedings of the 37th International Conference on Machine Learning, volume 119 of Proceedings of Machine Learning Research, pages 4411–4421.
+PMLR.
+  • Imani et al.
+(2023) Ayyoob Imani, Peiqin Lin, Amir Hossein Kargaran, Silvia Severini, Masoud Jalili Sabet, Nora Kassner, Chunlan Ma, Helmut Schmid, André Martins, François Yvon, and Hinrich Schütze.
+2023.
+Glot500: Scaling multilingual corpora and language models to 500 languages.
+In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1082–1117, Toronto, Canada.
+Association for Computational Linguistics.
+  • ImaniGooghari et al.
+(2022) Ayyoob ImaniGooghari, Silvia Severini, Masoud Jalili Sabet, François Yvon, and Hinrich Schütze.
+2022.
+Graph-based multilingual label propagation for low-resource part-of-speech tagging.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 1577–1589, Abu Dhabi, United Arab Emirates.
+Association for Computational Linguistics.
+  • Jill Levine and Lateef-Jan (2018) Suzanne Jill Levine and Katie Lateef-Jan. 2018. Untranslatability Goes Global. Routledge.
+  • Joanis et al.
+(2020) Eric Joanis, Rebecca Knowles, Roland Kuhn, Samuel Larkin, Patrick Littell, Chi-kiu Lo, Darlene Stewart, and Jeffrey Micher.
+2020.
+The Nunavut Hansard Inuktitut–English parallel corpus 3.0 with preliminary machine translation results.
+  • Jones et al.
+(2021) Alexander Jones, William Yang Wang, and Kyle Mahowald.
+2021.
+A massively multilingual analysis of cross-linguality in shared embedding space.
+In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 5833–5847.
+  • Joshi et al.
+(2019) Pratik Joshi, Christain Barnes, Sebastin Santy, Simran Khanuja, Sanket Shah, Anirudh Srinivasan, Satwik Bhattamishra, Sunayana Sitaram, Monojit Choudhury, and Kalika Bali.
+2019.
+Unsung challenges of building and deploying language technologies for low resource language communities.
+arXiv preprint arXiv:1912.03457.
+  • Joshi et al.
+(2020a) Pratik Joshi, Sebastin Santy, Amar Budhiraja, Kalika Bali, and Monojit Choudhury.
+2020a.
+The state and fate of linguistic diversity and inclusion in the NLP world.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 6282–6293.
+  • Joshi et al.
+(2020b) Pratik Joshi, Sebastin Santy, Amar Budhiraja, Kalika Bali, and Monojit Choudhury.
+2020b.
+The state and fate of linguistic diversity and inclusion in the nlp world.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 6282–6293.
+  • Jumelet et al.
+(2025) Jaap Jumelet, Leonie Weissweiler, and Arianna Bisazza.
+2025.
+MultiBLiMP 1.0: A massively multilingual benchmark of linguistic minimal pairs.
+arXiv.
+URL removed because this paper includes un-anonymized links to the current paper.
+This paper is published by authors unaffiliated with the authors of the current paper.
+  • Kakwani et al.
+(2020) Divyanshu Kakwani, Anoop Kunchukuttan, Satish Golla, Gokul N.C., Avik Bhattacharyya, Mitesh M.
+Khapra, and Pratyush Kumar.
+2020.
+IndicNLPSuite: Monolingual corpora, evaluation benchmarks and pre-trained multilingual language models for Indian languages.
+In Findings of the Association for Computational Linguistics: EMNLP 2020, pages 4948–4961.
+Association for Computational Linguistics.
+  • Kaplan et al.
+(2020) Jared Kaplan, Sam McCandlish, Tom Henighan, Tom B.
+Brown, Benjamin Chess, Rewon Child, Scott Gray, Alec Radford, Jeff Wu, and Dario Amodei.
+2020.
+Scaling laws for neural language models.
+arXiv.
+  • Karthikeyan et al.
+(2020) Karthikeyan, Zihan Wang, Stephen Mayhew, and Dan Roth.
+2020.
+Cross-lingual ability of multilingual BERT: An empirical study.
+In International Conference on Learning Representations.
+  • Katz (1987) Slava Katz.
+1987.
+Estimation of probabilities from sparse data for the language model component of a speech recognizer.
+IEEE Transactions on Acoustics, Speech, and Signal Processing, 35(3):400–401.
+  • Koehn (2023) Philipp Koehn. 2023. Statistical and neural machine translation. Dataset.
+  • Koto and Koto (2020) Fajri Koto and Ikhwan Koto.
+2020.
+Towards computational linguistics in Minangkabau language: Studies on sentiment analysis and machine translation.
+In Proceedings of the 34th Pacific Asia Conference on Language, Information and Computation, pages 138–148, Hanoi, Vietnam.
+Association for Computational Linguistics.
+  • Kreutzer et al.
+(2022) Julia Kreutzer, Isaac Caswell, Lisa Wang, Ahsan Wahab, Daan van Esch, Nasanbayar Ulzii-Orshikh, Allahsera Tapo, Nishant Subramani, Artem Sokolov, Claytone Sikasote, Monang Setyawan, Supheakmungkol Sarin, Sokhar Samb, Benoît Sagot, Clara Rivera, Annette Rios, Isabel Papadimitriou, Salomey Osei, Pedro Ortiz Suarez, Iroro Orife, Kelechi Ogueji, Andre Niyongabo Rubungo, Toan Q.
+Nguyen, Mathias Müller, André Müller, Shamsuddeen Hassan Muhammad, Nanda Muhammad, Ayanda Mnyakeni, Jamshidbek Mirzakhalov, Tapiwanashe Matangira, Colin Leong, Nze Lawson, Sneha Kudugunta, Yacine Jernite, Mathias Jenny, Orhan Firat, Bonaventure F.
+P.
+Dossou, Sakhile Dlamini, Nisansa de Silva, Sakine Çabuk Ballı, Stella Biderman, Alessia Battisti, Ahmed Baruwa, Ankur Bapna, Pallavi Baljekar, Israel Abebe Azime, Ayodele Awokoya, Duygu Ataman, Orevaoghene Ahia, Oghenefego Ahia, Sweta Agrawal, and Mofetoluwa Adeyemi.
+2022.
+Quality at a glance: An audit of web-crawled multilingual datasets.
+Transactions of the Association for Computational Linguistics, 10:50–72.
+  • Kudo and Richardson (2018) Taku Kudo and John Richardson.
+2018.
+SentencePiece: A simple and language independent subword tokenizer and detokenizer for neural text processing.
+In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 66–71.
+Association for Computational Linguistics.
+  • Kudugunta et al.
+(2023) Sneha Kudugunta, Isaac Caswell, Biao Zhang, Xavier Garcia, Christopher A.
+Choquette-Choo, Katherine Lee, Derrick Xin, Aditya Kusupati, Romi Stella, Ankur Bapna, and Orhan Firat.
+2023.
+Madlad-400: A multilingual and document-level large audited dataset.
+arXiv.
+  • Kunchukuttan et al.
+(2018) Anoop Kunchukuttan, Pratik Mehta, and Pushpak Bhattacharyya.
+2018.
+The IIT Bombay English-Hindi parallel corpus.
+In Proceedings of the Eleventh International Conference on Language Resources and Evaluation (LREC 2018), Miyazaki, Japan.
+European Language Resources Association (ELRA).
+  • Laurençon et al.
+(2022) Hugo Laurençon, Lucile Saulnier, Thomas Wang, Christopher Akiki, Albert Villanova del Moral, Teven Le Scao, Leandro Von Werra, Chenghao Mou, Eduardo González Ponferrada, Huu Nguyen, et al.
+2022.
+The BigScience ROOTS Corpus: A 1.6 TB Composite Multilingual Dataset.
+In Thirty-sixth Conference on Neural Information Processing Systems Datasets and Benchmarks Track.
+  • Lee et al.
+(2022) Katherine Lee, Daphne Ippolito, Andrew Nystrom, Chiyuan Zhang, Douglas Eck, Chris Callison-Burch, and Nicholas Carlini.
+2022.
+Deduplicating training data makes language models better.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics, pages 8424–8445.
+Association for Computational Linguistics.
+  • Leong et al.
+(2022) Colin Leong, Joshua Nemecek, Jacob Mansdorfer, Anna Filighera, Abraham Owodunni, and Daniel Whitenack.
+2022.
+Bloom library: Multimodal datasets in 300+ languages for a variety of downstream tasks.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, EMNLP 2022, Abu Dhabi, United Arab Emirates, December 7-11, 2022, pages 8608–8621.
+Association for Computational Linguistics.
+  • Levy et al.
+(2024) Mosh Levy, Alon Jacoby, and Yoav Goldberg.
+2024.
+Same task, more tokens: the impact of input length on the reasoning performance of large language models.
+In Annual Meeting of the Association for Computational Linguistics.
+  • Libovickỳ et al.
+(2020) Jindřich Libovickỳ, Rudolf Rosa, and Alexander Fraser.
+2020.
+On the language neutrality of pre-trained multilingual representations.
+In Findings of the Association for Computational Linguistics: EMNLP 2020, pages 1663–1674.
+  • Lin et al. (2024) Peiqin Lin, Shaoxiong Ji, Jörg Tiedemann, André FT Martins, and Hinrich Schütze. 2024. MaLA-500: Massive language adaptation of large language models. arXiv.
+  • Lin et al.
+(2022) Xi Victoria Lin, Todor Mihaylov, Mikel Artetxe, Tianlu Wang, Shuohui Chen, Daniel Simig, Myle Ott, Naman Goyal, Shruti Bhosale, Jingfei Du, Ramakanth Pasunuru, Sam Shleifer, Punit Singh Koura, Vishrav Chaudhary, Brian O’Horo, Jeff Wang, Luke Zettlemoyer, Zornitsa Kozareva, Mona Diab, Veselin Stoyanov, and Xian Li.
+2022.
+Few-shot learning with multilingual generative language models.
+In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 9019–9052.
+Association for Computational Linguistics.
+  • Lin et al.
+(2021) Xi Victoria Lin, Todor Mihaylov, Mikel Artetxe, Tianlu Wang, Shuohui Chen, Daniel Simig, Myle Ott, Naman Goyal, Shruti Bhosale, Jingfei Du, et al.
+2021.
+Few-shot learning with multilingual language models.
+arXiv preprint arXiv:2112.10668.
+  • LINDAT (2023) LINDAT. 2023. Lindat/clariah-cz repository. Dataset.
+  • Lingala Songs (2023) Lingala Songs. 2023. Lingala song lyrics. Dataset.
+  • Littell et al.
+(2017) Patrick Littell, David R.
+Mortensen, Ke Lin, Katherine Kairis, Carlisle Turner, and Lori Levin.
+2017.
+URIEL and lang2vec: Representing languages as typological, geographical, and phylogenetic vectors.
+In Proceedings of the 15th Conference of the European Chapter of the Association for Computational Linguistics: Volume 2, Short Papers, pages 8–14.
+Association for Computational Linguistics.
+  • Liu et al.
+(2019) Yinhan Liu, Myle Ott, Naman Goyal, Jingfei Du, Mandar Joshi, Danqi Chen, Omer Levy, Mike Lewis, Luke Zettlemoyer, and Veselin Stoyanov.
+2019.
+Roberta: A robustly optimized bert pretraining approach.
+arXiv.
+  • Liu et al.
+(2021) Zeyu Liu, Yizhong Wang, Jungo Kasai, Hannaneh Hajishirzi, and Noah A.
+Smith.
+2021.
+Probing across time: What does RoBERTa know and when?
+In Findings of the Association for Computational Linguistics: EMNLP 2021, pages 820–842, Punta Cana, Dominican Republic.
+Association for Computational Linguistics.
+  • Luukkonen et al.
+(2023) Risto Luukkonen, Ville Komulainen, Jouni Luoma, Anni Eskelinen, Jenna Kanerva, Hanna-Mari Kupari, Filip Ginter, Veronika Laippala, Niklas Muennighoff, Aleksandra Piktus, Thomas Wang, Nouamane Tazi, Teven Scao, Thomas Wolf, Osma Suominen, Samuli Sairanen, Mikko Merioksa, Jyrki Heinonen, Aija Vahtola, Samuel Antao, and Sampo Pyysalo.
+2023.
+FinGPT: Large generative models for a small language.
+In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing, pages 2710–2726, Singapore.
+Association for Computational Linguistics.
+  • LyricsTranslate (2023) LyricsTranslate. 2023. Lyricstranslate. Dataset.
+  • Ma et al. (2023) Chunlan Ma, Ayyoob ImaniGooghari, Haotian Ye, Ehsaneddin Asgari, and Hinrich Schütze. 2023. Taxi1500: A multilingual dataset for text classification in 1500 languages.
+  • Mabuya et al. (2023) Rooweither Mabuya, Jade Abbott, and Vukosi Marivate. 2023. Umsuka isizuluparallel corpus. Dataset.
+  • Mager et al.
+(2021) Manuel Mager, Arturo Oncevay, Abteen Ebrahimi, John Ortega, Annette Rios, Angela Fan, Ximena Gutierrez-Vasques, Luis Chiruzzo, Gustavo Giménez-Lugo, Ricardo Ramos, Ivan Vladimir Meza Ruiz, Rolando Coto-Solano, Alexis Palmer, Elisabeth Mager-Hois, Vishrav Chaudhary, Graham Neubig, Ngoc Thang Vu, and Katharina Kann.
+2021.
+Findings of the AmericasNLP 2021 shared task on open machine translation for indigenous languages of the Americas.
+In Proceedings of the First Workshop on Natural Language Processing for Indigenous Languages of the Americas, pages 202–217.
+Association for Computational Linguistics.
+  • Magueresse et al. (2020) Alexandre Magueresse, Vincent Carles, and Evan Heetderks. 2020. Low-resource languages: A review of past work and future challenges. arXiv preprint arXiv:2006.07264.
+  • Majliš (2011) Martin Majliš.
+2011.
+W2C – web to corpus – corpora.
+LINDAT/CLARIAH-CZ digital library at the Institute of Formal and Applied Linguistics (ÚFAL), Faculty of Mathematics and Physics, Charles University.
+  • Malkin et al.
+(2022) Dan Malkin, Tomasz Limisiewicz, and Gabriel Stanovsky.
+2022.
+A balanced data approach for evaluating cross-lingual transfer: Mapping the linguistic blood bank.
+In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 4903–4915, Seattle, United States.
+Association for Computational Linguistics.
+  • Martin et al.
+(2020) Louis Martin, Benjamin Muller, Pedro Javier Ortiz Suárez, Yoann Dupont, Laurent Romary, Éric de la Clergerie, Djamé Seddah, and Benoît Sagot.
+2020.
+CamemBERT: a tasty French language model.
+In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics, pages 7203–7219, Online.
+Association for Computational Linguistics.
+  • Masakhane (2023) Masakhane. 2023. Masakhane: A living collection of NLP projects for Africans, by Africans. Dataset.
+  • Mayer and Cysouw (2014) Thomas Mayer and Michael Cysouw.
+2014.
+Creating a massively parallel Bible corpus.
+In Proceedings of the Ninth International Conference on Language Resources and Evaluation, LREC 2014, Reykjavik, Iceland, May 26-31, 2014, pages 3158–3163.
+European Language Resources Association (ELRA).
+  • Meta (2024) Meta. 2024. The Llama 3 herd of models. arXiv.
+  • Meta AI (2024) Meta AI. 2024. The Llama 3 Herd of Models. arXiv preprint arXiv:2407.21783.
+  • Mirzakhalov et al.
+(2021) Jamshidbek Mirzakhalov, Anoop Babu, Duygu Ataman, Sherzod Kariev, Francis Tyers, Otabek Abduraufov, Mammad Hajili, Sardana Ivanova, Abror Khaytbaev, Antonio Laverghetta Jr., Bekhzodbek Moydinboyev, Esra Onal, Shaxnoza Pulatova, Ahsan Wahab, Orhan Firat, and Sriram Chellappan.
+2021.
+A large-scale study of machine translation in Turkic languages.
+In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 5876–5890, Online and Punta Cana, Dominican Republic.
+Association for Computational Linguistics.
+  • Moran et al.
+(2022) Steven Moran, Christian Bentz, Ximena Gutierrez-Vasques, Olga Pelloni, and Tanja Samardzic.
+2022.
+TeDDi sample: Text data diversity sample for language comparison and multilingual NLP.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 1150–1158, Marseille, France.
+European Language Resources Association.
+  • Morishita et al.
+(2020) Makoto Morishita, Jun Suzuki, and Masaaki Nagata.
+2020.
+JParaCrawl: A large scale web-based English-Japanese parallel corpus.
+In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 3603–3609, Marseille, France.
+European Language Resources Association.
+  • Muennighoff et al.
+(2023) Niklas Muennighoff, Alexander M.
+Rush, Boaz Barak, Teven Le Scao, Nouamane Tazi, Aleksandra Piktus, Sampo Pyysalo, Thomas Wolf, and Colin Raffel.
+2023.
+Scaling data-constrained language models.
+In Advances in Neural Information Processing Systems.
+  • Muennighoff et al.
+(2022) Niklas Muennighoff, Thomas Wang, Lintang Sutawika, Adam Roberts, Stella Biderman, Teven Le Scao, M Saiful Bari, Sheng Shen, Zheng-Xin Yong, Hailey Schoelkopf, et al.
+2022.
+Crosslingual generalization through multitask finetuning.
+arXiv preprint arXiv:2211.01786.
+  • Mukiibi et al.
+(2022) Jonathan Mukiibi, Andrew Katumba, Joyce Nakatumba-Nabende, Ali Hussein, and Joshua Meyer.
+2022.
+The makerere radio speech corpus: A Luganda radio corpus for automatic speech recognition.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 1945–1954.
+European Language Resources Association.
+  • Nakazawa et al.
+(2022) Toshiaki Nakazawa, Hideya Mino, Isao Goto, Raj Dabre, Shohei Higashiyama, Shantipriya Parida, Anoop Kunchukuttan, Makoto Morishita, Ondřej Bojar, Chenhui Chu, Akiko Eriguchi, Kaori Abe, Yusuke Oda, and Sadao Kurohashi.
+2022.
+Overview of the 9th workshop on Asian translation.
+In Proceedings of the 9th Workshop on Asian Translation, pages 1–36, Gyeongju, Republic of Korea.
+International Conference on Computational Linguistics.
+  • Nakazawa et al.
+(2021) Toshiaki Nakazawa, Hideki Nakayama, Chenchen Ding, Raj Dabre, Shohei Higashiyama, Hideya Mino, Isao Goto, Win Pa Pa, Anoop Kunchukuttan, Shantipriya Parida, Ondřej Bojar, Chenhui Chu, Akiko Eriguchi, Kaori Abe, Yusuke Oda, and Sadao Kurohashi.
+2021.
+Overview of the 8th workshop on Asian translation.
+In Proceedings of the 8th Workshop on Asian Translation (WAT2021), pages 1–45, Online.
+Association for Computational Linguistics.
+  • Nart (2023) Nart. 2023. Abkhaz text. Dataset.
+  • Neubig (2011) Graham Neubig. 2011. The Kyoto free translation task. http://www.phontron.com/kftt.
+  • Niyongabo (2023) Patrick Niyongabo. 2023. An english-kinyarwanda statistical machine translation (SMT) model. Dataset.
+  • Ogueji et al.
+(2021) Kelechi Ogueji, Yuxin Zhu, and Jimmy Lin.
+2021.
+Small data?
+no problem!
+exploring the viability of pretrained multilingual language models for low-resourced languages.
+In Proceedings of the 1st Workshop on Multilingual Representation Learning, pages 116–126.
+Association for Computational Linguistics.
+  • Ogunremi et al.
+(2023) Tolulope Ogunremi, Dan Jurafsky, and Christopher Manning.
+2023.
+Mini but mighty: Efficient multilingual pretraining with linguistically-informed data selection.
+In Findings of the Association for Computational Linguistics: EACL 2023, pages 1251–1266.
+  • Oladipo et al.
+(2022) Akintunde Oladipo, Odunayo Ogundepo, Kelechi Ogueji, and Jimmy Lin.
+2022.
+An exploration of vocabulary size and transfer effects in multilingual language models for African languages.
+In 3rd Workshop on African Natural Language Processing.
+  • OpenAI (2023) OpenAI. 2023. GPT-4 technical report. OpenAI.
+  • Ortiz Suarez et al.
+(2026) Pedro Ortiz Suarez, Laurie Burchell, Catherine Arnett, Rafael Mosquera-Gómez, Sara Hincapie-Monsalve, Thom Vaughan, Damian Stewart, Malte Ostendorff, Idris Abdulmumin, Vukosi Marivate, et al.
+2026.
+Commonlid: Re-evaluating state-of-the-art language identification performance on web data.
+arXiv preprint arXiv:2601.18026.
+  • Ortiz Suárez et al.
+(2019) Pedro Javier Ortiz Suárez, Benoît Sagot, and Laurent Romary.
+2019.
+Asynchronous pipeline for processing huge corpora on medium to low resource infrastructures.
+In 7th Workshop on the Challenges in the Management of Large Corpora (CMLC-7).
+Leibniz-Institut für Deutsche Sprache.
+  • Palen-Michel et al.
+(2022) Chester Palen-Michel, June Kim, and Constantine Lignos.
+2022.
+Multilingual open text release 1: Public domain news in 44 languages.
+In Proceedings of the Thirteenth Language Resources and Evaluation Conference, pages 2080–2089, Marseille, France.
+European Language Resources Association.
+  • Park et al.
+(2021) Sungjoon Park, Sungdong Kim, Jihyung Moon, Won Ik Cho, Kyunghyun Cho, Jiyoon Han, Jangwon Park, Chisung Song, Junseong Kim, Yongsook Song, et al.
+2021.
+Klue: Korean language understanding evaluation.
+In Thirty-fifth Conference on Neural Information Processing Systems (NeurIPS 2021).
+Advances in Neural Information Processing Systems.
+  • Patel et al.
+(2022) Ajay Patel, Bryan Li, Mohammad Sadegh Rasooli, Noah Constant, Colin Raffel, and Chris Callison-Burch.
+2022.
+Bidirectional language models are also few-shot learners.
+arXiv preprint arXiv:2209.14500.
+  • Penedo et al.
+(2025) Guilherme Penedo, Hynek Kydlíček, Vinko Sabolčec, Bettina Messmer, Negar Foroutan, Amir Hossein Kargaran, Colin Raffel, Martin Jaggi, Leandro Von Werra, and Thomas Wolf.
+2025.
+Fineweb2: One pipeline to scale them all — adapting pre-training data processing to every language.
+In Second Conference on Language Modeling.
+  • Pepperidge Farm (2024) Pepperidge Farm. 2024. Goldfish® Crackers. [Online; accessed 9-July-2024].
+  • Pfeiffer et al.
+(2021) Jonas Pfeiffer, Ivan Vulić, Iryna Gurevych, and Sebastian Ruder.
+2021.
+Unks everywhere: Adapting multilingual language models to new scripts.
+In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, pages 10186–10203.
+  • Pires et al.
+(2019) Telmo Pires, Eva Schlinger, and Dan Garrette.
+2019.
+How multilingual is multilingual BERT?
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 4996–5001.
+Association for Computational Linguistics.
+  • Podile and Eiselen (2016) Kholisa Podile and Roald Eiselen. 2016. NCHLT isiXhosa Named Entity Annotated Corpus.
+  • Polignano et al.
+(2019) Marco Polignano, Valerio Basile, Pierpaolo Basile, Marco de Gemmis, and Giovanni Semeraro.
+2019.
+Alberto: Modeling italian social media language with bert.
+IJCoL.
+Italian Journal of Computational Linguistics, 5(5-2):11–31.
+  • Ponti et al.
+(2020) Edoardo Maria Ponti, Goran Glavaš, Olga Majewska, Qianchu Liu, Ivan Vulić, and Anna Korhonen.
+2020.
+XCOPA: A multilingual dataset for causal commonsense reasoning.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 2362–2376, Online.
+Association for Computational Linguistics.
+  • Pyysalo et al.
+(2021) Sampo Pyysalo, Jenna Kanerva, Antti Virtanen, and Filip Ginter.
+2021.
+WikiBERT models: Deep transfer learning for many languages.
+In Proceedings of the 23rd Nordic Conference on Computational Linguistics (NoDaLiDa), pages 1–10, Reykjavik, Iceland (Online).
+Linköping University Electronic Press, Sweden.
+  • QCBS (2023) QCBS. 2023. Advanced Multivariate Analyses in R: Variation Partitioning. In QCBS R Workshop Series. Québec Centre for Biodiversity Science.
+  • Radford et al. (2018) Alec Radford, Karthik Narasimhan, Tim Salimans, and Ilya Sutskever. 2018. Improving language understanding by generative pre-training. OpenAI.
+  • Radford et al. (2019) Alec Radford, Jeff Wu, Rewon Child, David Luan, Dario Amodei, and Ilya Sutskever. 2019. Language models are unsupervised multitask learners. OpenAI.
+  • Raffel et al.
+(2020) Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter J.
+Liu.
+2020.
+Exploring the limits of transfer learning with a unified text-to-text transformer.
+J.
+Mach.
+Learn.
+Res., 21:140:1–140:67.
+  • Rama and Kolachina (2012) Taraka Rama and Prasanth Kolachina.
+2012.
+How good are typological distances for determining genealogical relationships among languages?
+In Proceedings of COLING 2012, pages 975–984.
+The COLING 2012 Organizing Committee.
+  • Rasooli and Tetreault (2015) Mohammad Sadegh Rasooli and Joel R. Tetreault. 2015. Yara parser: A fast and accurate dependency parser. Computing Research Repository, arXiv:1503.06733. Version 2.
+  • Reid and Artetxe (2022) Machel Reid and Mikel Artetxe. 2022. On the role of parallel data in cross-lingual transfer learning. arXiv preprint arXiv:2212.10173.
+  • Rozis and Skadiņš (2017) Roberts Rozis and Raivis Skadiņš.
+2017.
+Tilde MODEL - multilingual open data for EU languages.
+In Proceedings of the 21st Nordic Conference on Computational Linguistics, pages 263–265, Gothenburg, Sweden.
+Association for Computational Linguistics.
+  • Rust et al.
+(2021) Phillip Rust, Jonas Pfeiffer, Ivan Vulić, Sebastian Ruder, and Iryna Gurevych.
+2021.
+How good is your tokenizer?
+on the monolingual performance of multilingual language models.
+In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 3118–3135.
+Association for Computational Linguistics.
+  • SADiLaR (2023a) SADiLaR. 2023a. Mburisano covid-19 multilingual corpus. Dataset.
+  • SADiLaR (2023b) SADiLaR. 2023b. South african centre for digital language resources, nchlt corpus. Dataset.
+  • Sajjad et al.
+(2020) Hassan Sajjad, Ahmed Abdelali, Nadir Durrani, and Fahim Dalvi.
+2020.
+AraBench: Benchmarking dialectal Arabic-English machine translation.
+In Proceedings of the 28th International Conference on Computational Linguistics, pages 5094–5107, Barcelona, Spain (Online).
+International Committee on Computational Linguistics.
+  • Scao et al.
+(2022) Teven Le Scao, Angela Fan, Christopher Akiki, Elizabeth-Jane Pavlick, Suzana Ili’c, Daniel Hesslow, Roman Castagn’e, Alexandra Sasha Luccioni, Franccois Yvon, Matthias Gallé, Jonathan Tow, Alexander M.
+Rush, Stella Rose Biderman, Albert Webson, Pawan Sasanka Ammanamanchi, Thomas Wang, Benoît Sagot, Niklas Muennighoff, Albert Villanova del Moral, Olatunji Ruwase, et al.
+2022.
+Bloom: A 176b-parameter open-access multilingual language model.
+arXiv.
+  • Schwenk et al.
+(2021) Holger Schwenk, Vishrav Chaudhary, Shuo Sun, Hongyu Gong, and Francisco Guzmán.
+2021.
+WikiMatrix: Mining 135M parallel sentences in 1620 language pairs from Wikipedia.
+In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 1351–1361, Online.
+Association for Computational Linguistics.
+  • Severini et al.
+(2022) Silvia Severini, Ayyoob Imani, Philipp Dufter, and Hinrich Schütze.
+2022.
+Towards a broad coverage named entity resource: A data-efficient approach for many diverse languages.
+arXiv preprint arXiv:2201.12219.
+  • Siddhant et al.
+(2022) Aditya Siddhant, Ankur Bapna, Orhan Firat, Yuan Cao, Mia Xu Chen, Isaac Caswell, and Xavier Garcia.
+2022.
+Towards the next 1000 languages in multilingual machine translation: Exploring the synergy between supervised and self-supervised learning.
+arXiv preprint arXiv:2201.03110.
+  • Singh (2008a) Anil Kumar Singh.
+2008a.
+Named entity recognition for south and south East Asian languages: Taking stock.
+In Proceedings of the IJCNLP-08 Workshop on Named Entity Recognition for South and South East Asian Languages.
+  • Singh (2008b) Anil Kumar Singh.
+2008b.
+Natural language processing for less privileged languages: Where do we come from?
+where are we going?
+In Proceedings of the IJCNLP-08 Workshop on NLP for Less Privileged Languages.
+  • Souza et al.
+(2020) Fábio Souza, Rodrigo Nogueira, and Roberto Lotufo.
+2020.
+Bertimbau: pretrained bert models for brazilian portuguese.
+In Intelligent Systems: 9th Brazilian Conference, BRACIS 2020, Rio Grande, Brazil, October 20–23, 2020, Proceedings, Part I 9, pages 403–417.
+Springer.
+  • Stanford (2023) Stanford. 2023. Stanford nlp group datasets. Dataset.
+  • Strubell et al.
+(2019) Emma Strubell, Ananya Ganesh, and Andrew McCallum.
+2019.
+Energy and policy considerations for deep learning in NLP.
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 3645–3650.
+Association for Computational Linguistics.
+  • Team Gemma et al.
+(2025) Team Gemma, Aishwarya Kamath, Johan Ferret, Shreya Pathak, Nino Vieillard, Ramona Merhej, Sarah Perrin, Tatiana Matejovicova, Alexandre Ramé, Morgane Rivière, et al.
+2025.
+Gemma 3 technical report.
+arXiv preprint arXiv:2503.19786.
+  • Teferra Abate et al.
+(2018) Solomon Teferra Abate, Michael Melese, Martha Yifiru Tachbelie, Million Meshesha, Solomon Atinafu, Wondwossen Mulugeta, Yaregal Assabie, Hafte Abera, Binyam Ephrem, Tewodros Abebe, Wondimagegnhue Tsegaye, Amanuel Lemma, Tsegaye Andargie, and Seifedin Shifaw.
+2018.
+Parallel corpora for bi-directional statistical machine translation for seven Ethiopian language pairs.
+In Proceedings of the First Workshop on Linguistic Resources for Natural Language Processing, pages 83–90, Santa Fe, New Mexico, USA.
+Association for Computational Linguistics.
+  • Tenney et al.
+(2019a) Ian Tenney, Dipanjan Das, and Ellie Pavlick.
+2019a.
+Bert rediscovers the classical nlp pipeline.
+In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics, pages 4593–4601.
+  • Tenney et al.
+(2019b) Ian Tenney, Patrick Xia, Berlin Chen, Alex Wang, Adam Poliak, R Thomas McCoy, Najoung Kim, Benjamin Van Durme, Samuel R Bowman, Dipanjan Das, et al.
+2019b.
+What do you learn from context?
+probing for sentence structure in contextualized word representations.
+In International Conference on Learning Representations.
+  • Teodorescu et al.
+(2022) Daniela Teodorescu, Josie Matalski, Delaney Lothian, Denilson Barbosa, and Carrie Demmans Epp.
+2022.
+Cree corpus: A collection of nêhiyawêwin resources.
+In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 6354–6364.
+Association for Computational Linguistics.
+  • Tiedemann (2012) Jörg Tiedemann.
+2012.
+Parallel data, tools and interfaces in OPUS.
+In Proceedings of the Eighth International Conference on Language Resources and Evaluation (LREC’12), pages 2214–2218.
+European Language Resources Association (ELRA).
+  • Tiedemann (2020) Jörg Tiedemann.
+2020.
+The Tatoeba Translation Challenge – Realistic Data Sets for Low Resource and Multilingual MT.
+In Proceedings of the Fifth Conference on Machine Translation, pages 1174–1182.
+Association for Computational Linguistics.
+  • Turc et al. (2019) Iulia Turc, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. Well-read students learn better: On the importance of pre-training compact models. arXiv.
+  • Turc et al.
+(2021) Iulia Turc, Kenton Lee, Jacob Eisenstein, Ming-Wei Chang, and Kristina Toutanova.
+2021.
+Revisiting the primacy of english in zero-shot cross-lingual transfer.
+CoRR, abs/2106.16171.
+  • Ulukau (2023) Ulukau. 2023. Ulukau: The Hawaiian Electronic Library. https://ulukau.org/index.php?l=en.
+  • Virtanen et al.
+(2020) Pauli Virtanen, Ralf Gommers, Travis E.
+Oliphant, Matt Haberland, Tyler Reddy, David Cournapeau, Evgeni Burovski, Pearu Peterson, Warren Weckesser, Jonathan Bright, Stéfan J.
+van der Walt, Matthew Brett, Joshua Wilson, K.
+Jarrod Millman, Nikolay Mayorov, Andrew R.
+J.
+Nelson, Eric Jones, Robert Kern, Eric Larson, C J Carey, İlhan Polat, Yu Feng, Eric W.
+Moore, Jake VanderPlas, Denis Laxalde, Josef Perktold, Robert Cimrman, Ian Henriksen, E.
+A.
+Quintero, Charles R.
+Harris, Anne M.
+Archibald, Antônio H.
+Ribeiro, Fabian Pedregosa, Paul van Mulbregt, and SciPy 1.0 Contributors.
+2020.
+SciPy 1.0: Fundamental Algorithms for Scientific Computing in Python.
+Nature Methods, 17:261–272.
+  • Wang et al.
+(2023) Mingyang Wang, Heike Adel, Lukas Lange, Jannik Strötgen, and Hinrich Schütze.
+2023.
+NLNDE at semeval-2023 task 12: Adaptive pretraining and source language selection for low-resource multilingual sentiment analysis.
+CoRR, abs/2305.00090.
+  • Wang et al.
+(2020) Zirui Wang, Zachary C.
+Lipton, and Yulia Tsvetkov.
+2020.
+On negative interference in multilingual models: Findings and a meta-learning treatment.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 4438–4450.
+Association for Computational Linguistics.
+  • Warstadt et al.
+(2023) Alex Warstadt, Aaron Mueller, Leshem Choshen, Ethan Wilcox, Chengxu Zhuang, Juan Ciro, Rafael Mosquera, Bhargavi Paranjabe, Adina Williams, Tal Linzen, and Ryan Cotterell.
+2023.
+Findings of the BabyLM challenge: Sample-efficient pretraining on developmentally plausible corpora.
+In Proceedings of the BabyLM Challenge at the 27th Conference on Computational Natural Language Learning, pages 1–34, Singapore.
+Association for Computational Linguistics.
+  • Wenzek et al.
+(2020) Guillaume Wenzek, Marie-Anne Lachaux, Alexis Conneau, Vishrav Chaudhary, Francisco Guzmán, Armand Joulin, and Edouard Grave.
+2020.
+CCNet: Extracting high quality monolingual datasets from web crawl data.
+In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 4003–4012, Marseille, France.
+European Language Resources Association.
+  • Wikimedia (2023) Wikimedia. 2023. Wikimedia dumps.
+  • Wikipedia (2024) Wikipedia. 2024. Wikipedia.
+  • Wilie et al.
+(2020) Bryan Wilie, Karissa Vincentio, Genta Indra Winata, Samuel Cahyawijaya, Xiaohong Li, Zhi Yuan Lim, Sidik Soleman, Rahmad Mahendra, Pascale Fung, Syafri Bahar, and Ayu Purwarianti.
+2020.
+IndoNLU: Benchmark and resources for evaluating Indonesian natural language understanding.
+In Proceedings of the 1st Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics and the 10th International Joint Conference on Natural Language Processing, pages 843–857, Suzhou, China.
+Association for Computational Linguistics.
+  • Winata et al.
+(2022) Genta Winata, Shijie Wu, Mayank Kulkarni, Thamar Solorio, and Daniel Preotiuc-Pietro.
+2022.
+Cross-lingual few-shot learning on unseen languages.
+In Proceedings of the 2nd Conference of the Asia-Pacific Chapter of the Association for Computational Linguistics and the 12th International Joint Conference on Natural Language Processing (Volume 1: Long Papers), pages 777–791.
+Association for Computational Linguistics.
+  • Winata et al.
+(2023) Genta Indra Winata, Alham Fikri Aji, Samuel Cahyawijaya, Rahmad Mahendra, Fajri Koto, Ade Romadhony, Kemal Kurniawan, David Moeljadi, Radityo Eko Prasojo, Pascale Fung, Timothy Baldwin, Jey Han Lau, Rico Sennrich, and Sebastian Ruder.
+2023.
+NusaX: Multilingual parallel sentiment dataset for 10 Indonesian local languages.
+In Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics, pages 815–834.
+Association for Computational Linguistics.
+  • Winata et al.
+(2021) Genta Indra Winata, Andrea Madotto, Zhaojiang Lin, Rosanne Liu, Jason Yosinski, and Pascale Fung.
+2021.
+Language models are few-shot multilingual learners.
+In Proceedings of the 1st Workshop on Multilingual Representation Learning, pages 1–15.
+Association for Computational Linguistics.
+  • Wolf et al.
+(2020) Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Remi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander Rush.
+2020.
+Transformers: State-of-the-art natural language processing.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 38–45.
+Association for Computational Linguistics.
+  • Workshop et al.
+(2022) BigScience Workshop, Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilić, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, et al.
+2022.
+Bloom: A 176b-parameter open-access multilingual language model.
+arXiv preprint arXiv:2211.05100.
+  • Wu and Dredze (2020) Shijie Wu and Mark Dredze.
+2020.
+Are all languages created equal in multilingual BERT?
+In Proceedings of the 5th Workshop on Representation Learning for NLP, pages 120–130.
+Association for Computational Linguistics.
+  • Wu et al. (2022) Zhengxuan Wu, Isabel Papadimitriou, and Alex Tamkin. 2022. Oolong: Investigating what makes crosslingual transfer hard with controlled studies. arXiv.
+  • Xia et al.
+(2023) Mengzhou Xia, Mikel Artetxe, Chunting Zhou, Xi Victoria Lin, Ramakanth Pasunuru, Danqi Chen, Luke Zettlemoyer, and Veselin Stoyanov.
+2023.
+Training trajectories of language models across scales.
+In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 13711–13738.
+Association for Computational Linguistics.
+  • Xu et al.
+(2020) Liang Xu, Hai Hu, Xuanwei Zhang, Lu Li, Chenjie Cao, Yudong Li, Yechen Xu, Kai Sun, Dian Yu, Cong Yu, Yin Tian, Qianqian Dong, Weitang Liu, Bo Shi, Yiming Cui, Junyi Li, Jun Zeng, Rongzhao Wang, Weijian Xie, Yanting Li, Yina Patterson, Zuoyu Tian, Yiwen Zhang, He Zhou, Shaoweihua Liu, Zhe Zhao, Qipeng Zhao, Cong Yue, Xinrui Zhang, Zhengliang Yang, Kyle Richardson, and Zhenzhong Lan.
+2020.
+CLUE: A Chinese language understanding evaluation benchmark.
+In Proceedings of the 28th International Conference on Computational Linguistics, pages 4762–4772, Barcelona, Spain (Online).
+International Committee on Computational Linguistics.
+  • Xue et al.
+(2021) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel.
+2021.
+mT5: A massively multilingual pre-trained text-to-text transformer.
+In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 483–498, Online.
+Association for Computational Linguistics.
+  • Yang et al.
+(2020) Jian Yang, Shuming Ma, Dongdong Zhang, Shuangzhi Wu, Zhoujun Li, and Ming Zhou.
+2020.
+Alternating language modeling for cross-lingual pre-training.
+In Proceedings of the AAAI Conference on Artificial Intelligence, volume 34, pages 9386–9393.
+  • Zaydelman et al.
+(2016) Lyudmila Zaydelman, Irina Krylova, and Boris Orekhov.
+2016.
+The technology of web-texts collection of Russian minor languages.
+In Proceedings of the International Scientific Conference CPT2015, pages 179–181.
+  • Zevallos et al.
+(2022) Rodolfo Zevallos, John Ortega, William Chen, Richard Castro, Nuria Bel, Cesar Toshio, Renzo Venturas, Hilario Aradiel, and Nelsi Melgarejo.
+2022.
+Introducing qubert: A large monolingual corpus and bert model for southern quechua.
+In Proceedings of the Third Workshop on Deep Learning for Low-Resource Natural Language Processing, pages 1–13.
+  • Zhang et al.
+(2020) Shiyue Zhang, Benjamin Frey, and Mohit Bansal.
+2020.
+ChrEn: Cherokee-English machine translation for endangered language revitalization.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 577–595.
+Association for Computational Linguistics.
+  • Zhao et al.
+(2020) Mengjie Zhao, Tao Lin, Fei Mi, Martin Jaggi, and Hinrich Schütze.
+2020.
+Masking as an efficient alternative to finetuning for pretrained language models.
+In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP), pages 2226–2241.
+  • Zhou and Waibel (2021) Zhong Zhou and Alex Waibel.
+2021.
+Family of origin and family of choice: Massively parallel lexiconized iterative pretraining for severely low resource machine translation.
+arXiv preprint arXiv:2104.05848.
+  • Zueva et al.
+(2020) Anna Zueva, Anastasia Kuznetsova, and Francis Tyers.
+2020.
+A finite-state morphological analyser for Evenki.
+In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 2581–2589.
+European Language Resources Association.
 Appendix A Appendix
-A.1 Training Dataset Details
+A.1.  Training Dataset Details
 Data sources.
 As described in §3.1, we merge the text datasets compiled in Chang et al.
-(2023), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023).
+(2024a), Glot500 (Imani et al., 2023), and MADLAD-400 (clean split; Kudugunta et al., 2023).
 These datasets include popular multilingual corpora such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), and others.
@@
 Comprehensively, the Goldfish dataset includes:
-  • Chang et al. (2023):
+  • Chang et al. (2024a):
     OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021), Wikipedia (Wikipedia, 2024), No Language Left Behind (Costa-jussà et al., 2022), Leipzig Corpora Collection (Goldhahn et al., 2012), eBible translations (eBible, 2023), Tatoeba (Tiedemann, 2012, 2020), AfriBERTa (Ogueji et al., 2021), NusaX (Winata et al., 2023), AmericasNLP (Mager et al., 2021), Nunavut Hansard Inuktitut–English Parallel Corpus (Joanis et al., 2020), Cherokee-English ChrEn dataset (Zhang et al., 2020), Cherokee Corpus (Cherokee Corpus, 2023), Cree Corpus (Teodorescu et al., 2022), Languages of Russia (Zaydelman et al., 2016), Evenki Life newspaper (Zueva et al., 2020), transcribed Fula Speech Corpora (Cawoylel, 2023), IsiXhosa (Podile and Eiselen, 2016), Ewe Language Corpus (Gbedevi Akouyo et al., 2021), Makerere Luganda Corpora (Mukiibi et al., 2022), CMU Haitian Creole dataset (CMU, 2010), Tigrinya Language Modeling Dataset (Gaim et al., 2021), and Ulukau (Ulukau, 2023).
   • Glot500 (Imani et al., 2023):
-    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008), and XLSum (Hasan et al., 2021).
+    AI4Bharat (AI4Bharat, 2023), AI FOR THAI LotusCorpus (AI FOR THAI, 2023), Arabic Dialects Dataset (El-Haj et al., 2018), AfriBERTa (Ogueji et al., 2021), AfroMAFT (Adelani et al., 2022; Xue et al., 2021), Anuvaad (Anuvaad, 2023), AraBench (Sajjad et al., 2020), Autshumato (Autshumato, 2023) Bloom Library (Leong et al., 2022), CC100 (Conneau et al., 2020b), CCNet (Wenzek et al., 2020), CMU Haitian Creole (CMU, 2010), SADiLaR NCHLT corpus (SADiLaR, 2023b), Clarin (Clarin, 2023), DART (Alsarsour et al., 2018), Earthlings (Dunn, 2020), FFR Dataset (FFR Dataset, 2023), GiossaMedia (Góngora et al., 2022, 2021), Glosses (Camacho-Collados et al., 2016), Habibi (El-Haj, 2020), HinDialect (Bafna, 2022), HornMT (HornMT, 2023), IITB (Kunchukuttan et al., 2018), IndicNLP (Nakazawa et al., 2021), Indiccorp (Kakwani et al., 2020), isiZulu (Mabuya et al., 2023), JParaCrawl (Morishita et al., 2020), kinyarwandaSMT (Niyongabo, 2023), LeipzigData (Goldhahn et al., 2012), LINDAT (LINDAT, 2023), Lingala Song Lyrics (Lingala Songs, 2023), LyricsTranslate (LyricsTranslate, 2023), mC4 (Raffel et al., 2020), MTData (Gowda et al., 2021), MaCoCu (Bañón et al., 2022), Makerere MT Corpus (Mukiibi et al., 2022), Masakhane Community (Masakhane, 2023), Mburisano Covid Corpus (SADiLaR, 2023a), Menyo20K (Adelani et al., 2021a), Minangkabau corpora (Koto and Koto, 2020), MoT (Palen-Michel et al., 2022), NLLB seed (Costa-jussà et al., 2022), Nart Abkhaz text (Nart, 2023), OPUS (Tiedemann, 2012), OSCAR (Ortiz Suárez et al., 2019), ParaCrawl (Bañón et al., 2020), Parallel Corpora for Ethiopian Languages (Teferra Abate et al., 2018), Phontron (Neubig, 2011), QADI (Abdelali et al., 2021), Quechua-IIC (Zevallos et al., 2022), SLI GalWeb.1.0 (Agerri et al., 2018), Shami (Abu Kwaik et al., 2018), Stanford NLP (Stanford, 2023), StatMT (Koehn, 2023), TICO (Anastasopoulos et al., 2020), TIL (Mirzakhalov et al., 2021), Tatoeba (Tiedemann, 2020), TeDDi (Moran et al., 2022), Tilde (Rozis and Skadiņš, 2017), W2C (Majliš, 2011), WAT (Nakazawa et al., 2022), WikiMatrix (Schwenk et al., 2021), Wikipedia (Wikipedia, 2024), Workshop on NER for South and South East Asian Languages (Singh, 2008a), and XLSum (Hasan et al., 2021).
   • MADLAD-400 (Kudugunta et al., 2023):
@@
 We start with the corpus from Chang et al.
-(2023).
+(2024a).
 We then merge the dataset per language with Glot500 for languages that have not yet reached our 1GB maximum (after byte premium scaling).
@@
 These ambiguities primarily appear for lower-resource languages.
-Additionally, we drop several redundant language codes:
+We also drop several redundant language codes:
   • We drop ory_orya (Odia) in favor of the macrocode ori_orya because ory_orya is the only individual language within ori_orya for which we have any data.
@@
 Reported token counts use the tokenizer for the largest Goldfish model for that language.
-All dataset statistics can be downloaded at https://github.com/tylerachang/goldfish.
-A.2 Pre-Training Details
-As described in §3.2, we train monolingual language models for five dataset sizes when available after byte premium scaling: 5MB, 10MB, 100MB, 1GB, and full.
-The full dataset size (including all available data) is only included if a 1GB dataset is not available for a language.
-In total, the Goldfish include 350 5MB-dataset models, 288 10MB-dataset models, 166 100MB-dataset models, 83 1GB-dataset models, and 267 full-dataset models (1154 models total).
-Full hyperparameters are reported in Table 4.
-Hyperparameter 5MB,10MB 100MB,1GB,full
+Datasets can be downloaded at: https://huggingface.co/datasets/goldfish-models/fish-food.
+A.2.  Contamination Analysis
+While we aim to exclude the FLORES dataset from our training datasets, there is still the possibility that FLORES could be contaminated inadvertently due to the presence of web data.
+To check for any potential contamination of FLORES in the datasets, we tokenize each FLORES sequence and the entire training dataset for each language that is in both Goldfish and FLORES (204 languages).
+We compute the total number of times that the first 10 tokens of any FLORES example appears in the training dataset for the language.
+For 72% of languages, no FLORES examples appear in the training dataset at all.
+For 98% of languages, less than 10 FLORES examples appear in the training dataset (out of over 2000 FLORES examples total).
+The only two languages with notable FLORES contamination are smo_latn (Samoan; 7155 occurrences of FLORES examples in the training dataset) and knc_arab (Central Kanuri; 371 occurrences of FLORES examples in the training dataset).
+A.3.  Pretraining Details
+Hyperparameter 5MB, 10MB full, 100MB+
 Total parameters 39M 125M
@@
 Hidden size 512 768
-Intermediate hidden size 2048 3072
+Intermediate hidden 2048 3072
 Attention heads 8 12
@@
 Learning rate decay Linear
-Warmup steps 10% of pre-training
+Warmup steps 10% of pretraining
 Adam $\epsilon$ 1e-6
@@
 Attention dropout 0.1
-Table 4: Pre-training hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
-Tokenizers.
-All tokenizers are trained with vocabulary size 50K (Liu et al., 2019) on the same dataset size as their corresponding model (including byte premium scaling).
-We use SentencePiece tokenizers (Kudo and Richardson, 2018) with training text randomly sampled from the dataset for the desired language.
-To avoid memory errors, we limit tokenizer training text to 100MB after byte premium scaling.
-After tokenizer training, we tokenize each training dataset, concatenating text lines such that each sequence contains exactly 512 tokens.
-We run tokenization before shuffling and sampling to the desired dataset sizes, so our sequences of 512 tokens preserve contiguous text where possible, although several of our source corpora only exist in shuffled form.
-Finally, we sample our tokenized datasets to 5MB, 10MB, 100MB, and 1GB after byte premium scaling.
-[^6] When de-tokenized, the tokenized datasets result in slightly smaller datasets than the original text datasets, because the tokenizer truncates lines to create 512-token sequences.
-All reported dataset sizes account for this truncation.
+Table 4: Pretraining hyperparameters for Goldfish trained on different dataset sizes (Devlin et al., 2019; Turc et al., 2019; Radford et al., 2018).
 Architectures.
@@
 For the 100MB-, 1GB-, and full-dataset models, we use the 125M-parameter architecture equivalent to GPT-1 (Radford et al., 2018) (similar to BERT-base and RoBERTa; Devlin et al., 2019; Liu et al., 2019).
-Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2023), we use the small model size (39M parameters) from Turc et al.
+Because smaller models perform similarly to larger models in low-resource scenarios (Chang et al., 2024a), we use the small model size (39M parameters) from Turc et al.
 (2019) for the 10MB and 5MB dataset sizes.
 Training hyperparameters.
-Language models are pre-trained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022).
+Language models are pretrained using the Hugging Face Transformers library (Wolf et al., 2020) and code from Chang and Bergen (2022).
 We refrain from extensive hyperparameter tuning to avoid biasing our hyperparameters towards English (or any other selected tuning language).
@@
 Based on initial results using randomly-sampled languages, we find that learning rate 1e-4 also works well for the 39M-parameter models; this is in line with Chang et al.
-(2023), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
-We train each model for 10 epochs of the training data; multiple epochs of pre-training is beneficial in data-constrained scenarios (Muennighoff et al., 2023), but pre-training on more than 10 epochs often leads to overfitting (increases in eval loss) in the 5MB scenarios.
-For batch sizes, following GPT-1 (most similar to our models; Radford et al., 2018), we use batch size 64 (64$\times$512 = 32K tokens) for the 1GB-dataset models.
-We find that these larger batch sizes lead to overfitting for small datasets, so we use batch sizes 4, 8, and 32 for 5MB-, 10MB-, and 100MB-dataset models respectively (determined based on initial experiments with randomly-sampled languages).
-These correspond to batches of 2K, 4K, or 16K tokens.
-For full-dataset models, we use the batch size that would be used if rounding the dataset size down to 5MB, 10MB, or 100MB (recall that we do not train a full-dataset model when the 1GB dataset is available for a language).
-Compute costs.
-All language model pre-training runs together take a total of $1.65\times 10^{20}$ FLOPs.
-This is less than 1/1900×1/1900$\times$1 / 1900 × the computation used to train the original 175B-parameter GPT-3 model (Brown et al., 2020; $3.14\times 10^{23}$ FLOPs).
+(2024a), who find that learning rate 2e-4 works well for small models, and smaller learning rates reduce the speed of any potential overfitting.
+[…]
 Models are each trained on one NVIDIA GeForce GTX TITAN X, GeForce RTX 2080 Ti, TITAN Xp, Quadro P6000, RTX A4500, RTX A5000, or RTX A6000 GPU.
-In total, Goldfish pre-training takes the equivalent of approximately 15600 A6000 GPU hours.
+In total, Goldfish pretraining takes the equivalent of approximately 15600 A6000 GPU hours.
 Inference for FLORES perplexities and reasoning benchmarks takes approximately 250 A6000 GPU hours (primarily due to the large multilingual models used for comparison).
 Dataset merging, deduplication, and tokenization takes approximately 1600 CPU core hours.
-A.3 FLORES Evaluation Details
-[…]
-FLORES log-perplexities for all models and languages are reported in Table LABEL:tab:ppl-table.
-For Goldfish models, we report the log-perplexity for the model trained on the largest dataset for the language (i.e.
+A.4.  FLORES Evaluation Details
+In §4.1, we evaluate the Goldfish models, XGLM 4.5B, XGLM 7.5B, BLOOM 7.1B, MaLA-500 10B, and bigram models on FLORES log-perplexity (negative log-likelihood).
+To avoid confounds from different tokenizers across models, we compute log-perplexities at the sequence level.
+For fair comparison with multilingual models that need to determine the input language during the early parts of a sequence, we compute log-perplexity of the second half $s_{1}$ of each sequence given the first half $s_{0}$.
+We then compute the mean over sequences:
+$\textrm{LogPPL}_{\mathcal{M}}=\textrm{mean}_{s}\Big(-\textrm{log}(P_{\mathcal{M}}(s_{1}|s_{0}))\Big)$ (2)
+A lower log-perplexity indicates better performance, where $\mathcal{M}$ assigns higher probabilities to ground truth text (FLORES sequences).
+While imperfect, perplexity does not require annotated text data, it is predictive of performance on a variety of downstream tasks (Xia et al., 2023), and it has been used to measure language model quality in previous work (Kaplan et al., 2020; Hoffmann et al., 2022; Lin et al., 2024).
+In detail, the first and second half of each sequence are determined based on number of characters, so the halfway split is the same for all models considered.
+We round to the nearest token when the halfway split is in the middle of a subword token.
+Each model $\mathcal{M}$ then assigns some probability $P_{\mathcal{M}}(s_{1}|s_{0})$ regardless of tokenization, except for rounding the halfway point to the nearest token.
+The probability for any [UNK] (unknown) token is set to random chance $1/v$ where $v$ is the tokenizer vocabulary size.
+[^5] Otherwise, for unseen writing systems (e.g.
+Tibetan script tibt in XGLM), the probability $P(\textrm{[UNK]}|\textrm{[UNK] [UNK] ...})$ is very high, resulting in artificially low perplexities.
+Setting the [UNK] token probabilities to random chance has very little effect on log-perplexity scores except for the scenario of an unseen writing system.
+As our final log-perplexity score, we compute the mean negative-log-probability over all FLORES sequences in the target language.
+Because perplexities generally use geometric means, we use arithmetic means for log-perplexities.
+The final equation is presented in Equation 2.
+The mean FLORES perplexities for each model are reported in Table 1.
+For Goldfish models, we report the perplexity for the model trained on the largest dataset for the language (i.e.
 the 1GB-dataset model when available, otherwise the full-dataset model).
-Log-perplexities of the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.
-Bigram model details.
-For each FLORES language, we train a bigram model on the entire Goldfish dataset for that language, up to 1GB after byte premium scaling §3.1.
-The bigram model computes the probability of each token $w_{i}$ as $P(w_{i}|w_{i-1})$, computed based on raw bigram counts in the tokenized Goldfish dataset.
-The tokenizer is the same as the Goldfish tokenizer for that dataset (i.e.
-the 1GB-dataset model when available, or the full-dataset model).
-When a bigram is not observed in the dataset, we use backoff to unigram probability with a penalty multiplier of $\lambda=0.40$ (i.e.
-“stupid backoff”; Brants et al., 2007).
-We do not consider $n$-grams for $n>2$ because those $n$-grams often resort to backoff and are therefore much more sensitive to the backoff penalty term $\lambda$.
+Perplexities per language for the 5MB-, 10MB-, 100MB-, and 1GB-dataset models specifically are available at https://github.com/tylerachang/goldfish.
 Ambiguous or missing languages.
@@
     nus_latn $\to$ din_latn,
-    taq_latn $\to$ kab_latn,
+    taq_latn $\to$ kab_latn.
     Here, we use the closest relative in Goldfish that uses the same script.
@@
     sin_latn $\to$ hin_latn,
-    urd_latn $\to$ hin_latn,
+    urd_latn $\to$ hin_latn.
     These are languages that are missing from Goldfish and that are written in a nonstandard script for the language (e.g.
@@
     yue_hant $\to$ zho_hant,
-    zsm_latn $\to$ msa_latn,
+    zsm_latn $\to$ msa_latn.
     These languages map to multiple different Goldfish languages or are individual languages within a macrolanguage code included in Goldfish.
 When the option is available, we use the Goldfish language with more data.
-A.4 Reasoning Task Details
-In §5, we evaluate the Goldfish models, XGLM 4.5B, XGLM 7.5B, BLOOM 7.1B, and MaLA-500 10B on:
-  • Non-English Belebele (121 languages, reading comprehension; Bandarkar et al., 2024).
-For languages that are ambiguous or missing from Goldfish, we use the same language code mapping as in §A.3.
-Each Belebele example consists of a passage, a question, and four candidate answers.
-We evaluate model accuracy in selecting the correct answer by computing text probabilities for each “[passage] [question] [answer_option]”.
-No model exceeds 41% accuracy for any language (random chance 25%).
-  • XCOPA (11 languages, commonsense reasoning; Ponti et al., 2020).
-Each example consists of a premise sentence and two possible causes or effects (i.e.
-answer options).
-We use the task format and evaluation implementation in Gao et al.
-(2023).
-This selects answers based on a model’s computed text probabilities for each “[premise] [connecting_word] [cause/effect_option]”, where the connecting word is the translation of “because” (for causes) or “therefore” (for effects).
-  • Non-English XStoryCloze (10 languages, story commonsense; Lin et al., 2022).
-Each example consists of a context story and two possible story completions.
-We use the task format and evaluation implementation in Gao et al.
-(2023).
-This selects answers based on a model’s computed text probabilities for each “[story_context] [completion_option]”.
-All models are evaluated zero shot with no fine-tuning. Results per language are available at https://github.com/tylerachang/goldfish.
-Table 5: FLORES log-perplexity score ($\downarrow$) for each model and FLORES language. Parentheses indicate that the model is not trained specifically on that language.
-Language   FLORES log-perplexity
-Goldfish Bigram XGLM 4.5B XGLM 7.5B BLOOM 7.1B MaLA-500 10B
-ace_arab (287.55) (365.59) (260.50) (263.32) (232.76) (251.92)
-ace_latn 144.62 169.33 (202.67) (208.64) (198.50) 133.10
-acm_arab (96.60) (124.48) (93.25) (88.91) (85.28) 102.65
-acq_arab (95.54) (124.94) (92.14) (88.09) (82.88) (105.38)
-aeb_arab (116.01) (140.30) (113.84) (108.98) (102.76) (120.88)
-afr_latn 79.88 115.57 79.54 (162.73) (153.05) 85.61
-ajp_arab (98.92) (125.28) (96.25) (91.49) (85.73) 103.56
-aka_latn 132.48 162.51 (234.93) (239.68) 187.66 128.37
-als_latn 77.28 119.91 76.37 (220.78) (178.58) 89.49
-amh_ethi 83.36 111.87 110.54 (266.96) (195.14) 108.99
-apc_arab 173.18 179.37 (99.97) (94.47) (87.38) 106.58
-arb_arab 82.43 117.61 79.33 75.07 69.24 96.84
-arb_latn (245.97) (346.80) 211.13 (226.14) (221.60) (229.07)
-ars_arab (83.68) (118.50) (80.72) (76.51) (70.75) (98.25)
-ary_arab (128.66) (155.55) (131.80) (125.88) (114.30) 123.30
-arz_arab 116.98 146.50 (98.72) (92.96) (87.42) 112.84
-asm_beng 93.78 118.79 135.60 (227.40) 113.74 108.20
-ast_latn 86.86 118.36 (113.30) (112.31) (99.39) 82.54
-awa_deva (128.70) (169.05) (148.81) (135.84) (123.28) (141.50)
-ayr_latn 123.33 148.81 (239.30) (231.82) (228.59) 146.85
-azb_arab 154.24 185.83 163.32 (218.12) (225.63) 165.85
-azj_latn 74.28 106.31 75.33 (183.10) (185.84) 84.13
-bak_cyrl 79.24 108.34 (259.93) (272.19) (209.63) 92.72
-bam_latn 158.88 175.25 188.31 (212.45) 203.82 143.14
-ban_latn 121.16 137.57 (154.09) (183.02) (188.39) 114.41
-bel_cyrl 79.50 124.53 81.13 (226.99) (211.85) 91.22
-bem_latn 150.39 174.20 (218.51) (188.14) (237.98) 158.61
-ben_beng 80.71 109.08 90.55 79.78 76.54 94.62
-bho_deva 121.88 138.45 (168.84) (149.88) (120.97) (130.48)
-bjn_arab (297.57) (383.82) (260.05) (261.94) (238.42) (257.15)
-bjn_latn 111.90 140.40 (151.76) (150.82) (153.76) 104.37
-bod_tibt 118.58 142.65 (134.47) (137.26) (206.06) 120.94
-bos_latn 73.84 113.42 63.67 (155.27) (135.77) 67.75
-bug_latn 172.63 181.80 (206.57) (214.07) (218.78) (211.21)
-bul_cyrl 71.36 109.94 64.51 59.03 (148.72) 70.69
-cat_latn 76.30 115.98 65.93 60.65 60.55 66.57
-ceb_latn 94.04 125.05 (111.22) (192.82) (171.75) 103.37
-ces_latn 75.07 115.11 63.68 (154.38) (140.67) 71.35
-cjk_latn 219.90 239.42 (212.05) (211.78) (220.85) 189.97
-ckb_arab 89.11 121.96 (112.62) (290.64) (199.22) 107.81
-crh_latn 105.56 127.10 (175.84) (185.40) (180.14) 119.20
-cym_latn 77.50 114.42 97.27 (226.16) (204.15) 100.22
-dan_latn 74.09 111.50 60.26 (122.87) (129.58) 67.08
-deu_latn 73.91 118.06 57.93 57.08 (98.55) 62.85
-dik_latn 152.14 169.32 (197.55) (202.82) (216.99) (225.78)
-dyu_latn 183.05 210.69 (209.72) (216.63) (209.24) 189.79
-dzo_tibt 125.44 144.63 (213.72) (217.28) (238.66) 110.58
-ell_grek 78.38 119.29 68.55 65.99 (157.23) 93.41
-eng_latn 68.73 103.16 51.10 50.39 50.56 48.43
-epo_latn 75.57 110.48 118.68 (175.58) (149.97) 88.92
-est_latn 73.18 109.48 71.72 60.94 (173.50) 96.32
-eus_latn 70.76 103.94 112.97 70.24 70.78 101.32
-ewe_latn 128.71 144.86 (181.68) (202.81) (228.49) 137.60
-fao_latn 89.22 120.77 (182.64) (216.89) (180.74) 100.02
-fij_latn 107.24 126.01 (186.08) (136.05) (215.93) (121.46)
-fin_latn 75.34 114.20 63.80 55.50 (164.25) 82.10
-fon_latn 190.75 205.52 (209.87) (255.39) 239.02 190.66
-fra_latn 70.55 116.07 56.67 55.76 53.43 59.64
-fur_latn 114.09 131.26 (203.28) (198.54) (185.86) 107.32
-fuv_latn 165.36 188.23 (188.00) (201.33) (192.94) (184.41)
-gaz_latn 120.21 202.78 (163.96) (260.02) (262.77) (144.71)
-gla_latn 97.03 133.97 172.79 (245.08) (209.10) 120.94
-gle_latn 79.48 118.25 150.60 (228.43) (204.19) 106.38
-glg_latn 76.34 114.53 86.26 (107.84) (87.42) 74.13
-grn_latn 121.55 141.64 193.46 (230.52) (225.65) 119.96
-guj_gujr 84.11 110.95 100.35 (276.27) 96.00 97.04
-hat_latn 89.82 114.98 114.15 85.51 (163.36) 102.63
-hau_latn 86.28 116.50 115.81 (215.68) (215.07) 107.27
-heb_hebr 77.17 108.77 73.71 (175.48) (147.78) 111.24
-hin_deva 76.64 110.74 79.74 71.13 70.06 88.97
-hne_deva 141.43 149.99 (164.54) (153.61) (144.13) 110.53
-hrv_latn 71.16 110.07 61.93 (153.35) (135.73) 66.55
-hun_latn 75.05 113.40 64.13 (176.50) (164.44) 78.54
-hye_armn 78.37 115.07 85.23 (259.42) (224.51) 94.83
-ibo_latn 110.84 135.25 147.33 (248.34) 149.46 123.89
-ilo_latn 111.06 129.57 (133.69) (227.74) (209.68) 115.75
-ind_latn 72.11 101.11 62.46 60.13 59.44 65.23
-isl_latn 75.36 113.10 82.56 (200.00) (179.47) 93.17
-ita_latn 75.27 116.41 60.06 59.00 (86.91) 62.98
-jav_latn 90.19 112.98 102.58 (161.56) (154.48) 102.21
-jpn_jpan 68.51 100.99 63.11 61.53 (93.28) 68.24
-kab_latn 134.05 154.27 (214.03) (256.69) (215.19) 159.99
-kac_latn 128.22 137.82 (187.08) (248.17) (239.55) 146.59
-kam_latn (240.45) (264.47) (186.89) (202.02) (200.47) 172.19
-kan_knda 76.34 103.10 92.36 (268.95) 93.76 97.03
-kas_arab (252.81) (304.08) (276.83) (267.41) (245.94) (260.94)
-kas_deva 221.04 240.29 (235.98) (228.86) (231.57) (246.51)
-kat_geor 72.44 108.39 82.24 (261.79) (236.64) 96.53
-kaz_cyrl 73.52 102.69 76.14 (192.47) (186.61) 89.71
-kbp_latn 145.05 159.84 (217.85) (264.27) (286.31) 143.65
-kea_latn 145.29 157.19 (180.88) (187.08) (182.18) 135.49
-khk_cyrl 77.40 103.60 87.49 (238.21) (192.81) 97.14
-khm_khmr 98.82 139.14 114.88 (323.85) (240.85) 117.47
-kik_latn 165.85 177.82 (222.18) (244.09) 227.12 148.48
-kin_latn 87.00 118.79 (227.82) (225.85) 135.94 113.52
-kir_cyrl 70.91 99.21 114.55 (224.69) (202.97) 94.96
-kmb_latn 179.18 199.36 (205.24) (196.76) (216.45) 168.66
-kmr_latn 99.93 130.74 155.62 (234.52) (213.35) 120.91
-knc_arab 181.38 274.16 (223.08) (222.78) (214.43) (228.52)
-knc_latn 170.17 206.34 (229.82) (227.51) (239.24) (242.18)
-kon_latn 132.91 143.02 182.94 (190.80) (189.74) (126.76)
-kor_hang 72.23 102.86 69.83 63.54 (122.61) 73.28
-lao_laoo 91.00 120.61 110.10 (268.95) (231.49) 107.59
-lij_latn 140.38 168.14 (198.93) (195.96) (193.23) 122.99
-lim_latn 123.42 154.99 (180.85) (201.88) (192.90) (116.94)
-lin_latn 106.45 122.73 144.57 (183.46) 167.57 116.17
-lit_latn 71.55 110.06 67.23 (188.51) (163.54) 92.94
-lmo_latn 162.18 201.95 (203.26) (199.20) (198.36) 152.10
-ltg_latn 121.88 138.23 (217.38) (241.02) (229.77) (227.25)
-ltz_latn 85.00 123.40 (154.55) (149.83) (188.39) 103.80
-lua_latn 152.56 162.72 (175.04) (167.40) (201.83) 147.08
-lug_latn 118.73 139.82 168.52 (213.70) 165.48 141.39
-luo_latn 139.03 155.32 (210.18) (218.56) (222.85) 163.70
-lus_latn 95.13 126.88 (157.04) (217.70) (211.72) 131.48
-lvs_latn 70.94 110.34 70.45 (187.39) (179.57) 86.42
-mag_deva 126.54 139.42 (156.88) (143.63) (128.48) (133.86)
-mai_deva 123.50 142.56 (174.69) (165.19) (126.39) 102.51
-mal_mlym 80.56 111.47 92.78 (221.87) 88.15 99.60
-mar_deva 82.32 110.97 94.21 (200.05) 92.70 100.33
-min_arab (308.31) (399.00) (269.74) (273.93) (254.70) (275.42)
-min_latn 108.38 128.37 (167.37) (164.01) (160.32) 125.86
-mkd_cyrl 73.08 109.43 73.22 (134.89) (162.92) 76.94
-mlt_latn 83.60 125.90 (280.75) (279.00) (237.59) 90.70
-mni_beng (176.58) (274.13) (279.36) (271.00) (188.18) (275.78)
-mos_latn 187.64 198.01 (228.35) (236.65) (241.18) 188.11
-mri_latn 97.39 130.22 (191.98) (187.99) (180.89) 109.67
-mya_mymr 86.45 125.64 119.52 90.49 (224.44) 121.00
-nld_latn 71.72 112.22 60.19 (111.54) (115.43) 65.13
-nno_latn 80.80 114.77 (73.23) (153.12) (151.17) 76.16
-nob_latn 76.13 109.96 64.70 (122.09) (131.71) 68.04
-npi_deva 82.42 111.11 90.72 (193.40) 86.21 86.53
-nso_latn 119.03 139.17 (166.19) (234.76) 181.13 123.81
-nus_latn (217.98) (276.12) (259.51) (266.18) (293.48) (319.71)
-nya_latn 97.55 121.06 (213.77) (202.58) 180.39 118.77
-oci_latn 97.75 135.49 (152.88) (126.52) (129.14) 93.79
-ory_orya 87.64 114.38 (113.19) (387.80) (103.65) 105.48
-pag_latn 122.99 139.15 (175.08) (184.33) (184.39) 120.19
-pan_guru 85.22 116.57 107.69 (298.41) 99.43 101.39
-pap_latn 94.63 126.00 (166.14) (177.79) (183.93) 118.24
-pbt_arab 104.87 136.61 120.56 (263.66) (210.05) 130.62
-pes_arab 75.11 112.64 70.67 (152.72) (145.67) 84.23
-plt_latn 87.89 121.62 116.37 (236.48) (176.35) 102.04
-pol_latn 72.87 114.38 61.18 (146.68) (130.92) 69.19
-por_latn 72.38 110.34 59.80 57.79 55.73 63.06
-prs_arab 96.31 120.44 (80.74) (150.59) (141.22) 83.70
-quy_latn 121.48 144.34 185.50 125.42 (196.23) 132.70
-ron_latn 75.68 118.45 62.78 (151.24) (135.97) 70.39
-run_latn 112.80 135.64 (229.50) (226.86) 152.96 127.91
-rus_cyrl 73.59 117.13 58.22 57.38 (110.95) 65.18
-sag_latn 162.70 167.88 (182.49) (171.77) (196.64) 150.49
-san_deva 134.36 156.91 167.12 (188.71) (167.33) 140.17
-sat_olck 148.03 156.26 (217.91) (217.35) (298.99) 124.11
-scn_latn 124.37 157.13 (172.60) (187.74) (175.53) 108.90
-shn_mymr 162.52 182.33 (253.11) (475.53) (373.38) (341.88)
-sin_sinh 83.19 114.20 97.91 (304.98) (235.47) 106.82
-slk_latn 72.22 113.10 63.60 (181.84) (157.25) 78.82
-slv_latn 71.40 111.05 64.99 (176.07) (147.81) 75.62
-smo_latn 103.20 142.22 (197.35) (213.38) (214.55) 120.15
-sna_latn 93.53 123.08 (228.56) (225.74) 180.82 125.19
-snd_arab 91.04 116.69 150.37 (257.35) (212.74) 117.36
-som_latn 104.72 138.92 125.40 (245.50) (229.18) 135.24
-sot_latn 99.85 142.34 (171.86) (235.46) 192.76 122.30
-spa_latn 77.45 116.81 63.10 61.91 59.82 66.71
-srd_latn 116.75 135.26 (202.61) (200.61) (191.63) 106.39
-srp_cyrl 76.66 116.32 72.33 (175.02) (149.07) 71.86
-ssw_latn 124.51 143.31 186.61 (232.90) (220.26) 133.78
-sun_latn 91.07 116.80 109.40 (167.76) (162.37) 98.90
-swe_latn 73.76 109.95 62.27 (106.34) (126.73) 65.90
-swh_latn 79.45 109.31 89.60 76.81 98.22 95.63
-szl_latn 131.79 157.70 (184.61) (229.41) (208.14) 122.68
-tam_taml 79.87 107.64 88.40 78.39 79.09 95.70
-taq_latn (241.82) (267.82) (216.97) (231.56) (222.07) (225.98)
-taq_tfng None None (301.75) (289.98) (261.59) (396.24)
-tat_cyrl 76.67 107.20 (222.35) (227.31) (192.83) 88.82
-tel_telu 80.76 106.79 89.99 77.70 88.52 95.52
-tgk_cyrl 79.61 113.99 (267.56) (281.19) (213.45) 100.89
-tgl_latn 86.11 123.04 91.13 (168.86) (162.76) 89.23
-tha_thai 73.84 102.82 71.03 64.31 (177.26) 87.04
-tir_ethi 107.10 134.77 142.22 (300.08) (258.62) 133.08
-tpi_latn 141.91 164.61 (218.27) (212.39) (197.37) 109.12
-tsn_latn 111.75 143.35 184.89 (240.07) 186.95 127.47
-tso_latn 111.93 133.30 (236.13) (239.86) 205.36 130.50
-tuk_latn 81.23 107.98 (239.28) (256.07) (198.90) 114.93
-tum_latn 132.63 154.17 (233.00) (187.16) 237.42 141.81
-tur_latn 69.75 100.52 64.44 61.13 (130.11) 86.26
-twi_latn 131.51 148.42 (211.42) (216.89) 174.63 128.30
-tzm_tfng None None (206.77) (206.42) (243.59) (332.15)
-uig_arab 75.68 105.62 (317.09) (367.46) (206.46) 105.21
-ukr_cyrl 76.60 116.07 64.77 (129.73) (145.72) 72.19
-umb_latn 182.34 211.34 (199.72) (209.91) (221.59) 174.09
-urd_arab 83.96 117.74 90.05 79.04 85.58 98.80
-uzn_latn 71.09 107.26 112.09 (243.88) (217.18) 96.67
-vec_latn 114.88 147.99 (161.71) (155.44) (160.51) 108.87
-vie_latn 77.66 120.08 68.06 64.89 61.00 74.36
-war_latn 118.02 153.85 (161.42) (203.73) (174.18) 132.17
-wol_latn 141.12 158.76 202.72 (225.23) 167.95 161.96
-xho_latn 93.68 121.39 144.06 (216.55) 155.76 122.42
-ydd_hebr 109.90 144.63 (260.53) (286.49) (210.28) 128.80
-yor_latn 123.24 167.30 174.33 (246.05) 154.45 148.67
-yue_hant 90.03 121.49 (70.31) (86.41) 61.81 69.76
-zho_hans 78.92 121.82 66.34 65.42 59.08 70.09
-zho_hant 93.56 125.41 (72.53) (89.36) 63.06 (75.40)
-zsm_latn 73.25 100.45 67.03 (83.84) (76.22) 72.90
-zul_latn 87.90 118.40 135.47 (218.33) 186.47 115.47
-Table 6: Goldfish languages with corresponding dataset sizes.
+Table 5: Goldfish languages with corresponding dataset sizes. Horizontal lines separate languages with at least
+1000MB, 100MB, 10MB, and 5MB of available data.
 Language   Language Script Byte Scaled Tokens Dataset Proportions